A Complete Bibliography of ACM Transactions on Internet Technology

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

24 August 2024 Version 1.76

Title word cross-reference

3 [LYW^{+}21, YLC^{+}22, ZXP^{+}22]. < [BMS02]. > [BMS02]. _{2} [VSKEOZM22]. _{2R} [SABG17]. δ [BCCA^{+}21]. K [CYD^{+}20, BGK14]. N [HZ11, WZKP19].

-barrier [CYD^{+}20]. -Risk [BCCA^{+}21]. -Tier [WZKP19].

19 [CLL23, NZQX22, SPE^{+}22, SDB21, SZT22, YV22].

24-hour [GS07a].

4.0 [CLW^{+}22].

5G [LWFD21, LQW21, LLSW22, SPCC23]. 5G-Aided [LLSW22]. 5G/6G [LWFD21].

6G [LWFD21].

802.15.4 [CMML22].

Abductive [GAL18]. Ability [MHCF22]. Abnormal [DCD^{+}21]. Abuse [TBG^{+}18]. Accelerated [EDC20]. ACCENT [PP11]. Acceptability [VDV18]. Acceptance [SPM^{+}13]. Access [Ano15, ADGM23, DSVA19, DLZ^{+}16, GSZ^{+}23, PV17, SK17, TSY^{+}21, DSNK08, KS07]. accesses [DK04]. Accessibility [PMFS17]. accessible [SMFR08]. Account [CCC^{+}23, CLJ^{+}21, WL23]. Accountability [BBP18]. Accountable [BCFB18, GAL18].

Accounts [CLJ^{+}21]. Accumulator [RQL^{+}21]. Accuracy [YXP^{+}18]. Achieving [GWF^{+}21, SS11, YBW19]. ACM [MFR^{+}21]. Acoustic [WWZ^{+}23]. across [BZVS18]. Active [YZL^{+}24]. Activities [MMP^{+}14]. Activity [CLM19]. Actuator [SS20, WVHTK21]. Ad [APAC18, SLG^{+}22, ZHDD07]. Ad-Based [APAC18]. ad-hoc [ZHDD07]. Adapt [RPR22]. Adaptation [DSVA19, SS20, HNF^{+}05, SMFR08]. Adapting [WL07]. Adaptive [ATD22, ADAP19, EHY19, HCW^{+}21, MP14, OGP^{+}18, YLC^{+}22, ZWC^{+}17, CH05, HJ03, KS03, LWM^{+}21, MPS04, RZJ20, ZHH04]. Addiction [FLR23]. address [HZCS10, HBGF02]. administering [HBGF02]. Administrative [Sin13a]. admission [CH05]. Adoption [GdOW14, Web17]. ADTO [YZL^{+}24]. Advanced [SST^{+}16]. Advances [CSS17, DNJ19, FLLM22, SLPZ23]. Adversarial [JPSS17, QLJ^{+}19, ZZW^{+}22]. Advertisement [CDM^{+}14]. Advertisement-Financed [CDM^{+}14]. advertisements [AM03]. Advertiser [Glu10]. Affect [CDPR17, MS17]. Affecting [PVL^{+}17]. Affective [AVB17, FPR16]. Affinities [GdOW14, RS09]. Affinity [Ji02]. Affinity-based [Ji02]. against [AKA^{+}23, BRK04, HJ08, MMK^{+}22, YW10]. Age [ALS23]. Agent [CDPR17, LB04, STK17, YCC17, AJ03, AGPS05, HS19, HJ03, MEAK^{+}21]. Agent-Based [CDPR17]. Agents [AVB17, KMB^{+}22, CPV03, JMSP06, BGL04]. Aggregate [BLD^{+}15]. Aggregation [ABC^{+}17, XCL07]. Agile [JSAA22]. Agnostic [Nov19]. AI [CGG^{+}22, CSW^{+}22, CTS^{+}24, GDLM22, GJAT^{+}21, GAL^{+}22, HSRK23, HXB^{+}22, LQSW21, LQVK21, OKM21, RKY^{+}22, TNJJ22, WRWM21, ZBF^{+}19]. AI-Based [OKM21]. AI-Empowered [WRWM21, LQSW21]. AI-Enabled [GAL^{+}22, LQVK21]. AI-IoT [CGG^{+}22, GDLM22]. Aided [CLS^{+}22, LLSW22, ZXP^{+}22]. Air [GJAT^{+}21]. Airborne [SRK22]. Airport [GAT^{+}21]. Alert [CTS^{+}24, CDJ^{+}22]. Alexa [LHAT22, MHCF22]. Algorithm [ABCL17, BBS21, DWF24, HML^{+}21, LMS^{+}21, LPX^{+}21, SZT22, SGOS19, TF21, WCC20, ZDCB18, MBBW07]. Algorithms [HKV14, MQUXK22, MMJ21, SK17, BRRT05, KRRT06, MPS04, MS05]. aliasing [GM04]. Aligning [EM19]. Allocation [ADAP19, JSAA22, KA20, MRS^{+}22b, MMI23]. allowing [FLL06]. Alzheimer [HCW^{+}21]. Am [MHCF22]. Amazon [MHCF22, WLL^{+}13]. Ambiguous [LJLN16]. Analyse [MBE22]. Analysing [SCPB22]. Analysis [BLD^{+}15, BG21, CLZ^{+}20, CCJ^{+}14, FPA^{+}23, Gel09, GRR20, GVM^{+}23, HZ11, JYW^{+}19, JLC20, LMSS23, LS21, MMJ21, NGER20, PSZ24, SABG17, TGBG20, WS17, WMW^{+}22, YV22, YDZ^{+}21, BRRT05, EV07, GNK11, GBAR08, Liu12, MBBW07, OHKS04, SHH^{+}06]. Analytical [RPA^{+}17, SR13]. Analytics [LSK^{+}17a, LQVK21, MA23, PGP^{+}21, SH22]. Analyze [YV22]. Anarchy [DABP14]. Android [BAM^{+}22, UNBAT22]. Anger [DP17]. angles [PRD09]. Anomalous [ZHL^{+}16]. Anomaly [MSW^{+}16, ZOC11]. Anonymity [AJSS13]. anonymization [QLJ^{+}19]. Anonymized [MMV11]. Anonymous [CLJ^{+}21, PVL^{+}17, PO19]. Answering [GR04, LSLY19, LMSTM14, ZSL^{+}17]. answers [ALG04]. Ant [WSLT21]. Anti [CLS^{+}22, WSM21]. Anti-Eavesdropping [CLS^{+}22]. Anti-spam [WSM21]. Apples [TBG^{+}18]. Appliances [SH22]. Application [BBS21, BTC^{+}23, CLM19, HCW^{+}21, MRB19, MED19, Mor17, OGP^{+}18, SKA^{+}23, VDV18, XWML19, ATB^{+}11, CH05, KMW09, OSSV05, SHH^{+}06, USR09, VPR07, CDM10]. Application-Driven [XWML19]. application-level [CH05]. Applications [AO22, CLW^{+}22, CXG21, CGT^{+}21, CMML22, CGL^{+}16, FCS^{+}18, GSS^{+}14, KBBI15, LLC^{+}23, LGKL20, NZ22, PDS20, PCV^{+}21, PCBG19, RAR22, SAB^{+}18, SDB21, SS11, SSC23, PBL^{+}22, SCPB22, TMK^{+}12, WZKP19, WG23, WK18, BLSW04, BCF^{+}07, CDMF07, CGMH^{+}06, GLJ^{+}12, JN08, MBC^{+}05, Var03]. Approach [ADAP19, BBH^{+}14, EM19, GWF^{+}21, KYY17, LMZ13, LYM^{+}18, MGHB16, Nov19, OWK^{+}19, PGP^{+}21, RPA^{+}17, RZP^{+}22, RTcR19, RCP^{+}15, SR13, SPAT21, SPKTG22, VAS24, VBD^{+}22, XWML19, XLL20, ZGF^{+}23, CDIW05, FS04, GM04, MAM03, MGB^{+}07, Rin09, TGRBD07, TJLC08, YASU01, ZHDD07, ZH09]. Approximate [HC14, TGRBD07]. Approximation [HKV14, PWSG22]. Apps [JCH^{+}18, MHCF22]. Araneus [MAM03]. Architectural [PJZ18]. Architecture [AVB17, FYW^{+}22, FXYX23, FYZ19, KLMH03, LMS^{+}21, LGKL20, MEAK^{+}21, MMP^{+}14, PRKD20, PCBG19, PPDG19, BCP^{+}04, FT02, Jor09, LHTL06, WRC01]. Architectures [IRJ^{+}21, SSKW20, SLPZ23, KS07]. archiving [MPC06]. area [AOVP08, BVT06]. Argument [LSK^{+}17a]. Argumentation [ABC^{+}17, CT17, GLT17, KYY17, LT16, WS17]. Argumentation-Enabled [ABC^{+}17]. Argumentative [LPB^{+}17]. arguments [BC01]. Art [LT16]. Articles [GRR20]. Artificial [IRJ^{+}21, LPX^{+}21, PSL^{+}20]. Aspect [HJWW20]. assess [ZH09]. Assessment [ABC^{+}17, CRP17, LJG18, RDC16, ZKC^{+}22, Dal11]. Assignment [CLFX24, HBGF02]. Assistant [ASÖY23, UY22]. Assisted [CDC14, KSL^{+}21, LHL^{+}22, ZMGW22]. Asynchronous [WZKP19]. Attachment [JS13]. Attack [AKA^{+}23, LZK^{+}22, MMK^{+}22, BRK04, MBBW07]. Attack-Resistant [LZK^{+}22]. Attacking [SO17, TSM^{+}23]. Attacks [AE24, DWGC23, YCM^{+}13, DCAT12, HJ08, HGW07, YW10]. Attention [ZMT^{+}23, BGL04, TJGY22]. Attentive [HLG^{+}21, HMLH21]. Attribute [BJ15, DSVA19, PO19]. Attribute-based [DSVA19, PO19]. Attributed [YMY^{+}23]. Attributes [BWL16]. Auction [CKKK14, NT21, Guo02]. Auctions [HKV14, RML12, AJ03, DRJ^{+}07, Gel09, HJPB06]. Audience [DTE17]. Audit [BCFB18]. Auditing [TPQC22]. Augmentation [YXL^{+}21]. Augmented [MBS19, PDS20]. Augmenting [DWGC23]. Australia [ZW17]. Authenticated [BCO13, ZXH16]. Authentication [ATS^{+}21, ADA^{+}22, CIY^{+}21, JKI^{+}21, LXZ^{+}22, MRS^{+}22a, SGC16, XZG^{+}22, YLM^{+}23, DCAT12]. Authenticity [RKY^{+}22]. Authority [XJ20]. Authorization [MRS^{+}22a]. Autism [CXH^{+}21]. Automated [DCL^{+}22, JMSP06, ZKC^{+}22]. Automatic [KBNV18, LSLY19, ZGB18]. Automatically [EM19]. Autonomic [MED19]. Autonomous [KMB^{+}22, HJ03]. Autonomously [RPR22]. Auxiliary [VJL^{+}14]. Avionic [SPKTG22]. Avoid [MRY^{+}23]. Aware [ASÖY23, GSS^{+}14, GLWH17, JPCL22, LGC20, MRB19, NYB^{+}19, OKR^{+}14, RIB18, WLW^{+}23, YZY^{+}14, YBW19, AR12, BCMS06, BCCA^{+}21, CDMF07, FLD12, HAST21, HMLH21, JN08, MYS^{+}12, SD12, SLG^{+}22, SCPB22, TJLC08]. Awareness [ZWW^{+}23].

Backbone [WNN^{+}22, BSS02]. Backdoor [WWJ^{+}22]. BACKM [WG23]. BACKM-EHA [WG23]. Bad [FLR23, TBG^{+}18]. balanced [GLJ^{+}12]. Balancing [CLM^{+}11, DOG^{+}22, DABP14, WY01]. Bandwidth [GD17, SKA^{+}23, TPK10]. banner [AM03]. barrier [CYD^{+}20]. BASECASS [HCBRM23]. Based [ASBH^{+}16, ASO^{+}22, APAC18, AHM^{+}15, ATS^{+}21, ABCL17, Ano15, BBC14, BJ15, BBH^{+}14, CHC^{+}21, CLW^{+}22, CLS^{+}22, CDPR17, CLM19, CO16, DFLT22, DLZ^{+}16, FYW^{+}22, FFKG19, GSZ^{+}23, Glu10, HML^{+}21, JPSS17, JKI^{+}21, KBNV18, KKMK16, KZLG21, LMZ13, LPX^{+}21, LYW23, LXZ^{+}22, LGGB^{+}21, LYW^{+}21, LMSTM14, LP21, MRS^{+}22a, DMGR^{+}17, MED19, NBM19, NPP^{+}15, OKM21, PAS13, QZDG22, RWXC20, RQL^{+}21, RCP^{+}15, RZAD17, SAB^{+}18, SF21, SGOS19, SH22, STJ^{+}21, SCZ^{+}21, TPQC22, WQC^{+}19, WMG^{+}21, WNN^{+}22, WLW^{+}23, XZG^{+}22, XvHWW18, XM17, XSW^{+}22, YPFY21, YCH^{+}22, YYM^{+}19, YLC^{+}22, ZXYL16, ZSY^{+}17, ZTL^{+}21, ZGF^{+}23, ZWW^{+}23, ZLS^{+}22, ZB20, AE24, AAJ21, AGPS05, AKR01, ADA^{+}22, BTGM22, BLMP20, BLMP22, BGL04, BCP08, BC23, CE21, CDIW05, CIY^{+}21, CH05, CXG21, CMML22, DSVA19, DNJ19, EV07, FYZ^{+}21, FMC19, GNW^{+}20, GHD21, GCP^{+}20, GH06, GLF02, HZHC12, HXZ^{+}20, HJWW20, Ji02]. based [JN08, KFB^{+}14, KGKK21, KKK21, KK21, KA20, KGAR22, KG10, LLNF12, LNTL23, LHZ^{+}21, LZJ^{+}21, LSZ^{+}21, MEAK^{+}21, MSG^{+}21, MDDB19, MMI23, MQB22, MBS19, MGB^{+}07, NT21, NGER20, PRKD20, PSK10, PCV^{+}21, PO19, RKY^{+}22, RS09, SSA^{+}21, SLG^{+}22, SCLB24, SHH^{+}06, TPK10, UNBAT22, WCX^{+}23, WSLT21, WMW^{+}22, WYC^{+}23, XCRY22, YZL^{+}24, YASU01, Zdu08, ZJQ^{+}21, RPA^{+}17, SS11, AJP07, BRK04, PP11, WG23]. Bases [LSLY19, ZSY^{+}17]. Batch [ZJL^{+}15]. Battlefield [SSA^{+}21]. Bayes [MBS19]. Bayes-based [MBS19]. Bayesian [AZKG21]. BCI [KKK21]. BCI-based [KKK21]. BDI [AVB17]. BDS [WCX^{+}23].

Beauty [YCC17]. Bee [LPX^{+}21]. Behavior [ASO^{+}22, BLD^{+}15, IDS19, LGGB^{+}21, PSZ24, SHH^{+}06, vdADO^{+}08]. Behavior-Based [ASO^{+}22, SHH^{+}06]. behavioral [MS05]. Behaviors [GD17]. Behind [LFL17, CIY^{+}21]. benchmarking [LYW^{+}05]. BERT [ZMT^{+}23]. Between [ZLHD15, YC18, YBMV22]. Betweenness [WQC^{+}19]. Beyond [GLFV^{+}21, PSL^{+}20]. Bi [FYZ19]. Bi-Lanczos [FYZ19]. bid [DRJ^{+}07]. bidding [AJ03, HJPB06]. Bids [Glu10]. Big [LS21, LCS21, MAK^{+}22, PSA^{+}20, RPA^{+}17, SPG22]. Big-Data [PSA^{+}20]. bigwig [BMS02]. Bilinear [RQL^{+}21]. binary [GH06]. Biomedical [CE21, MAK^{+}22]. Bistatic [CYD^{+}20]. Bitcoin [TNJJ22, MCS18]. BLE [ZKC^{+}22]. BLE-enabled [ZKC^{+}22]. Blind [DXP^{+}23]. Block [JYW^{+}19, JDZ^{+}21, RMMH22]. Blockchain [AE24, AAJ21, AKA^{+}23, BLTH22, CLJ^{+}21, CGT^{+}21, CXW^{+}21, DCZ^{+}21, FYZ^{+}21, GSZ^{+}23, GWF^{+}21, LNTL23, LHZ^{+}21, LGGB^{+}21, LSZ^{+}21, MFR^{+}21, NT21, DFL^{+}23, SCZ^{+}21, SXZ^{+}21, TSY^{+}21, WMG^{+}21, WCX^{+}23, WG23, WYC^{+}23, XZY^{+}21, XSSD23, YPFY21]. Blockchain-Based [GSZ^{+}23, WMG^{+}21, AE24, LNTL23, LHZ^{+}21, LSZ^{+}21, NT21, WCX^{+}23]. Blockchain-empowered [TSY^{+}21]. Blockchain-Enabled [DCZ^{+}21, AKA^{+}23, FYZ^{+}21, WG23]. Blocks [FYT17]. blog [LYF^{+}09]. Bloom [GNW^{+}20]. Blowfish [CAN^{+}21]. BogusBiter [YW10]. Bootstrapping [MQB22]. Bot [ZTH^{+}23]. Bottom [AHM^{+}15]. Bottom-Up [AHM^{+}15]. Bound [DZHV16, ABMW05]. Box [PMN23]. BPMS [PPDG19]. BPMS-RA [PPDG19]. Brain [HML^{+}21, KGAR22]. Brains [YCC17]. brave [BC01]. Breach [GAC18]. Breaking [FLR23, HCBRM23]. Breast [MHA^{+}21]. Bridge [YBMV22]. Broker [XIS22, SMFR08]. Brokering [AV16]. Browser [XM17]. browsers [HJ08]. Browsing [LYM^{+}18, HNF^{+}05]. Bundled [GdOW14]. Bundling [YWML19]. Bursting [GSS^{+}14]. Bus [ALA^{+}19]. Business [ACDLM19, FYT17, GNR19, PPDG19, YBW19]. Buyer [HNGN23]. buying [HJPB06]. Byzantine [XZY^{+}21].

C [Van08]. Cache [RMP10, JAT^{+}06, YADI02]. caching [CLN05, LSCZ05, Wil02]. Calibrating [YXP^{+}18]. California [CBM23]. Cameras [DCL^{+}22]. Can [ABR17, CPV03, SHB06, MBS19]. CAN-TM [MBS19]. Cancer [KSL^{+}21, MHA^{+}21]. Cannot [SdMA^{+}14]. Canonical [Mor17]. Capabilities [GL14, BDT04]. Caps [DJ15]. CaptchaStar [HCBRM23]. Card [GAC18]. Cards [GAC18]. Care [OALA17, RPA^{+}17, MGB^{+}21]. CARES [JPCL22]. Carrying [PV17]. Cascading [FPA^{+}23]. Case [EHY19, FAGB14, GAT^{+}21, DD07]. Categories [FSC15]. categorization [LXW^{+}12]. Category [WMWM20, JKR07]. centered [BHPY21, SDB21, ZMT^{+}23]. Centrality [DMGR^{+}17, WQC^{+}19]. Centres [CTS^{+}24]. Centric [CAN^{+}21, DLZ^{+}16, KKMK16, LGKL20, MRS^{+}22b, TEMH19, TMK^{+}12, FFKG19, KA20, MPR^{+}23, PC22]. Certificate [PCP^{+}20]. certified [Ung05]. Chain [MBS19, RMMH22, DFL^{+}23, XZY^{+}21, YPFY21, HGW07, SCZ^{+}21]. Chaining [WCC20]. Challenges [AHJ^{+}20, AGKW14, BSBP16, DFLT22, KMB^{+}22, LWM^{+}21, NZ22, SLG^{+}22, SLPZ23]. change [CGM03, Liu12]. change-impact [Liu12]. Channel [CSS20, MMJ21]. Channels [NT21, WMW^{+}22]. Chaos [LC16]. Character [MBP^{+}17]. Characteristics [LCKN05]. Characterization [BYCE07, BLD^{+}15, DPCM16].

Characterizing [AKR01, ACG^{+}11, GD17, GS05, SK13, CFTV03]. charging [TPK10]. Chasing [RCP^{+}15]. Chat [LXC^{+}13]. Cheating [BKS^{+}14]. checking [NCEF02, vdADO^{+}08]. Children [LHAT22]. China [ZW17]. Chinese [LWH^{+}21]. Choreographies [SBC20]. Cipher [JDZ^{+}21]. Circuit [LXZ^{+}22]. Cities [AZKG21, CGG^{+}22, DKP17, GDLM22, KZLG21, LHZ^{+}21, LQSW21, SPE^{+}22, WRWM21]. Citizen [LFL17]. City [CXG21, SdMA^{+}14, PBL^{+}22, SPCC23, PMFS17, WLW^{+}23]. Classification [CT17, JG10, KSL^{+}21, KGAR22, MSG^{+}21, QZDG22, UNBAT22, ZWC^{+}22, ZH09]. Classifier [BC23]. CLEVER [KRRT06]. Click [Glu10]. Click-Through [Glu10]. Client [RMP10]. Clients [SK13]. CloseUp [VAKK19]. Closing [ZLHD15]. Closure [MCS18]. Cloud [AHJ^{+}20, AO22, BMG^{+}19, BLMP20, BLMP22, CIY^{+}21, CMTD16, FYZ19, FCS^{+}18, GHD21, GD17, GSS^{+}14, JSAA22, KGKK21, KSL^{+}21, LPR19, MMK^{+}22, MQUXK22, MDDB19, MBS19, PJZ18, RMP10, RWXC20, SAB^{+}18, SPG22, TEMH19, TPQC22, TGBG20, UNBAT22, VASD19, XFL^{+}23, XSW^{+}22, YJL^{+}22, YSZ^{+}22, YBZ14, ZZF^{+}23, ZB20, DKP12, HAST21, PP11]. Cloud-Assisted [KSL^{+}21]. Cloud-Based [SAB^{+}18, BLMP20, BLMP22, CIY^{+}21, KGKK21]. Cloud-edge-based [GHD21]. Cloud-enabled [AHJ^{+}20]. Cloud-native [ZZF^{+}23]. cloudlet [MAB19]. CloudMF [FCS^{+}18]. Clouds [CGS23, FGS20, GS17, LC16, OGP^{+}18, SCL^{+}19, ZDCB18]. Cluster [CMML22]. Cluster-tree-based [CMML22]. clustered [WY01]. Clustering [JCH^{+}18, VAS24, ZJQ^{+}21, LYF^{+}09, PRD09]. Clusters [CLFX24, FXYX23, Ji02]. CNN [CLFX24, LYW^{+}21]. Co [YMY^{+}23, VSKEOZM22]. Co-guided [YMY^{+}23]. code [ZZF^{+}23]. Coding [YLL^{+}17]. Cognitive [AKOB^{+}21, CLF^{+}19, CXH^{+}21, Liu20, LWFD21, LQW21, MED19, PP11, ZTH^{+}23]. collaboration [SBG07]. Collaborative [BCFB18, CO16, DNJ19, FFKG19, HSLH17, OHKS04, PMN23, PO19, SS11, SPCC23, YSZ^{+}22, YSW^{+}17, ZLL^{+}20, LLSL08]. Collection [LZBN17, SPCC23]. Collective [ABC^{+}17]. Collusion [YJL^{+}22]. Collusion-free [YJL^{+}22]. Colony [LPX^{+}21, WSLT21]. Commerce [BWL16, DCZ^{+}21, GWF^{+}21, SXZ^{+}21, Var03, VPR07, XLL20, ZH09, Ung05]. Commitment [BBC14]. Commitment-Based [BBC14]. Commons [KAS14]. Communication [BPSD17, Liu20, LZK^{+}22, MRS^{+}22a, ZZW^{+}22]. Communication-Efficient [LZK^{+}22]. Communications [FMC19, MPR^{+}23, PACH20]. Communities [DKP17, NPP^{+}15, RZAD17, YMY^{+}23, ZWC^{+}17, ZSL^{+}17]. Community [BCN17, KLS^{+}17, VAKK19, GS05]. Community-Driven [VAKK19]. Comorbidity [MED19]. Comparative [NGER20, OKM21]. Comparing [GNK11]. Comparison [MS17]. Compatible [LDG^{+}23]. compete [BGL04]. Competition [CB15]. Competitive [KAS14, BSS02]. Complementary [SGOS19]. Complex [OKR^{+}14, YLM^{+}23, ZTH^{+}23, CTZZ06]. component [JN08]. component-based [JN08]. components [CGMH^{+}06, GBAR08, Van08]. compose [MGB^{+}07]. Composite [MQB22]. Composition [AV16, LJS^{+}14, LMZ13, PGT^{+}18, YBZ14, BCP08]. composition-oriented [BCP08]. Compositions [BBH^{+}14]. Compressed [PCP^{+}20, ABMP07]. Compression [STJ^{+}21, ZZW^{+}22, PP11]. Compression-Based [STJ^{+}21]. compressor [MPC06]. Computation [ADAP19, DCZ^{+}21, LHL^{+}22, LYM^{+}18, MAB19, DFL^{+}23]. Computation-Intensive [LYM^{+}18]. Computational [BBP18, BCO13, GAL^{+}22, SSC23]. Computer [SK17]. Computing [AAJ21, AZKG21, ATS^{+}21, BAM^{+}22, CGG^{+}22, CLF^{+}19, CYG^{+}21, CAV14, CSS17, FYW^{+}22, FGS20, GDLM22, HAST21, KBBI15, LOD19, LMS^{+}21, Liu20, LCS21, LLL22, LLSW22, MMK^{+}22, MRB19, MAB19, MDDB19, PML^{+}19, SF21, SPAT21, SSA^{+}21, SZT22, SKH22, PBL^{+}22, SPCC23, THS06, VAS24, WCC20, WWJ^{+}22, WTS^{+}21, XZG^{+}22, XZJO22, XFL^{+}23, XSW^{+}22, YSZ^{+}22, YZL^{+}24, ZLS^{+}22, ZB20, ZMGW22, BCMS06, DKP12, ML08, PP11, Van08]. Computing-based [XZG^{+}22]. Concept [GK23, LLNF12]. concept-based [LLNF12]. Concepts [BSBP16, LJLN16, SLG^{+}22]. Conceptual [SPM^{+}13, ZHH04]. concerns [DR05]. Conco [ZTH^{+}23]. Conco-ERNIE [ZTH^{+}23]. Concurrency [ACDLM19]. Concurrent [XZY^{+}21]. Condemning [DP17]. Conditional [FYZ^{+}21, XZG^{+}22]. Conduct [RCP^{+}15]. confidence [KG10]. Confidential [CGS23]. Confidentiality [MAK^{+}22]. configuration [ATB^{+}11]. Configurator [MD22]. Configure [RPR22]. configuring [HBGF02]. Conflicts [KBNV18, LMZ13]. Conformance [vdADO^{+}08]. Congestion [DFLT22, SK24]. Connected [BCN17, BCCA^{+}21, DKP17, HXB^{+}22, LQVK21, MJ22, RKY^{+}22, SATPR22, SPE^{+}22, VBD^{+}22, ZWC^{+}17]. Connecting [BI17]. conscious [ABMP07]. Consensus [ABCL17, DRC^{+}23, RZJ20, SXZ^{+}21, XSSD23]. Considering [BWL16]. Consistency [SS11, KMW09, NCEF02, YADI02]. Consistent [DWF24]. Consolidation [DvRDHB22]. Constrained [GZL^{+}21, Nov19, YLL^{+}17]. Constraint [RPR22]. Constructing [GPM^{+}18, JYW^{+}19, LJLN16].

Construction [ADGM23]. consumer [BGL04]. Consumption [MRY^{+}23, VSKEOZM22]. Container [BLMP20, ZB20]. Containerized [ZGD23]. contemporary [BF06]. Contempt [DP17]. Content [AHM14, AAF18, CDM^{+}14, FPR16, GLWH17, LHAT22, LJLN16, LXC^{+}13, CDIW05, Coo03, HNF^{+}05, JKR07]. Context [AR12, BCCA^{+}21, JPCL22, LMZ13, LZJ^{+}21, MYS^{+}12, PSA21, SNBC12, TEMH19, BCMS06, CDMF07, FLD12, HZHC12, Hoc02, MGB^{+}07, SD12]. Context-Aware [JPCL22, AR12, BCCA^{+}21, MYS^{+}12, BCMS06, CDMF07, FLD12, SD12]. Context-Based [LMZ13, LZJ^{+}21, MGB^{+}07]. Context-Driven [TEMH19]. Context-sensitive [SNBC12]. ContextAiDe [PCBG19]. Contexts [CJW^{+}23]. Contextual [SO17, YSW^{+}17]. Contextualization [SS11]. Continuity [FYT17]. Continuum [BMG^{+}19]. Contract [KK21]. Control [AHJ^{+}20, APAC18, AKOB^{+}21, ACG^{+}11, ADAP19, DWGC23, DSVA19, DFLT22, DLZ^{+}16, EDC20, GSZ^{+}23, KZLG21, LGGB^{+}21, PV17, SK17, TSY^{+}21, ZTL^{+}21, BDT04, CH05, KS07]. Control-path [DWGC23]. Control-Theoretic [ADAP19]. Controlled [PLZW18]. Controller [SK24]. Controlling [CMTD16, KMW09, MD22, PACH20]. Controls [SDB21]. Convolutional [FYZ^{+}21, MHA^{+}21, WCZ^{+}21, XCRY22]. cookies [DCAT12, Kri01]. Cooperation [XZG^{+}22]. Cooperative [CYWW22, IDS19, XFL^{+}23]. coordinated [LSCZ05]. Coordination [PHR^{+}21]. Core [KRRT06]. Coronavirus [GJAT^{+}21, JGH^{+}22]. Correlated [GdOW14]. Correlating [GD17]. Correlation [GJAT^{+}21, WEJ14].

Corrigenda [Vas05]. corrupted [CS09]. Cost [GSS^{+}14, HAST21, ISG^{+}22, Web17, ZB20, AAA^{+}20]. Cost-Aware [GSS^{+}14]. Cost-Efficient [ZB20, HAST21]. Costs [BTH^{+}17]. countermeasures [FLD12]. Coupled [GZL^{+}21]. Coupled-Layer [GZL^{+}21]. Coverage [CYD^{+}20]. COVID [CLL23, NZQX22, SPE^{+}22, SDB21, SZT22, YV22]. COVID-19 [CLL23, NZQX22, SPE^{+}22, SDB21, SZT22, YV22]. CPS [YXP^{+}18]. Cracking [CSS20]. crawlers [MPS04]. creating [CDIW05]. Credential [PO19]. Credit [DGWW15, GAC18]. Crime [HLLS21]. Crisis [NYB^{+}19]. Crisis-Relevant [NYB^{+}19]. criteria [DOG^{+}22]. Critical [CRP17, OKM21]. Cross [CL24, GSZ^{+}23, XM17, ZXH16]. Cross-Browser [XM17]. Cross-Domain [CL24]. Cross-Layer [ZXH16]. Cross-Organizational [GSZ^{+}23]. Crow [MSG^{+}21]. Crowd [ASO^{+}22, GHK17, LZBN17, NZ22, PCBG19, RDC16, CH05, MSG^{+}21]. Crowd-sensing [NZ22, PCBG19]. Crowd-Sourced [LZBN17]. Crowdsensing [JPCL22, LOD19, LZW^{+}22]. CrowdService [PGT^{+}18]. Crowdsourced [BB23, DZHV16, JCH^{+}18]. Crowdsourcing [CWLZ19, NBM19, PGT^{+}18, PMFS17]. Cryptocurrencies [LBC^{+}24]. Cryptocurrency [LMSS23]. Cryptographic [MJ22]. cryptography [PP11]. CSR [GPM^{+}18]. Curated [ZSY^{+}17]. Curation [AHM14]. Currency [MCS18]. Current [BSBP16, CPV03]. Customer [BWL16]. Customers [NGER20]. customized [THS06]. Customizing [SKA^{+}23]. Cyber [ALS23, CDJ^{+}22, CGT^{+}21, DXP^{+}23, FYZ19, FYZ^{+}21, FPA^{+}23, GAT^{+}21, HAD22, ISG^{+}22, JDZ^{+}21, KGKK21, LJS^{+}14, LMS^{+}21, LSZ^{+}21, NLLC21, PBJP21, SLPZ23, VAK17, WCY^{+}23, YXL^{+}21]. Cyber-alert [CDJ^{+}22]. Cyber-Espionage [LJS^{+}14]. Cyber-Manufacturing [DXP^{+}23, FPA^{+}23, SLPZ23, WCY^{+}23]. Cyber-Physical [CGT^{+}21, GAT^{+}21, ISG^{+}22, KGKK21, NLLC21, PBJP21, VAK17, LSZ^{+}21, YXL^{+}21]. Cyber-Physical-Social [FYZ19, FYZ^{+}21]. Cyberdeception [GCK^{+}22]. Cybersecurity [AO22, LNTL23, WMW^{+}22]. cycling [CMML22].

D [LYW^{+}21, YLC^{+}22, ZXP^{+}22]. DaaS [WHM^{+}22]. DADC [CMML22]. DAN [HMLH21]. DAN-SNR [HMLH21]. DANCE [ZZW^{+}22]. Dandelion [WDK^{+}24]. Darknet [GVM^{+}23]. Darknets [CCJ^{+}14]. DarkVec [GVM^{+}23]. Data [ASBH^{+}16, ADGM23, ASW^{+}22, BCFB18, BBS21, CPV^{+}16, CCM17, CLW^{+}22, CDJ^{+}22, DJ15, DZHV16, DLZ^{+}16, EM19, FFKG19, GSZ^{+}23, GAC18, GWF^{+}21, HSLH17, KIG^{+}19, KBBI15, LMZ13, LBC^{+}24, LHZ^{+}21, LZBN17, LGGB^{+}21, LGKL20, LQVK21, LS21, LCS21, LP21, MSW^{+}16, MGHB16, MEAK^{+}21, MBE22, MMV11, MAK^{+}22, NDO^{+}17, NZ22, PV17, PSA^{+}20, PVL^{+}17, PGP^{+}21, RKY^{+}22, RPA^{+}17, RQL^{+}21, RTcR19, SF21, SS11, SKH22, SWAHP21, SPCC23, SPG22, TEMH19, TPQC22, WCX^{+}23, WLW^{+}23, WVHTK21, YLL^{+}17, YXL^{+}21, ZGB18, ZXYL16, ZTL^{+}21, BCMS06, CS09, FFP09, MAM03, PPV05, XCL07, PBJP21]. Data-Centric [DLZ^{+}16, LGKL20, TEMH19, FFKG19]. Data-Driven [ASW^{+}22, CDJ^{+}22]. Data-Hiding [RKY^{+}22]. data-intensive [MAM03]. Data-throttling [RTcR19]. database [ABMP07, Ji02, LYW^{+}05, ZXS08]. Databases [GPM^{+}18, YASU01]. DataOps [GAT^{+}21]. Datasets [CAN^{+}21, PLZW18, WQC^{+}19]. DDoS [AE24, DWGC23, HGW07]. De-anonymization [QLJ^{+}19]. Deal [DWGC23]. Debates [LPB^{+}17]. Debating [LSK^{+}17a]. DECENT [MD22]. Decentralized [ABCL17, KBBI15, MD22, PSP22, PAS13, WEJ14, YPFY21, CGT^{+}21]. Decision [CRP17, SAB^{+}18, YBW19]. Decision-Aware [YBW19]. Decisions [ASÖY23]. decoder [XCRY22]. Deconvolution [DXP^{+}23]. Decryption [PCV^{+}21]. Deduplication [SKH22]. Deep [CE21, CLS^{+}22, CLL23, FYZ^{+}21, HSRK23, HLG^{+}21, HMLH21, KSL^{+}21, LOD19, LXZ^{+}22, MQUXK22, MSG^{+}21, MMJ21, MAK^{+}22, PGP^{+}21, RTR^{+}22, RWXC20, SPE^{+}22, SZT22, UNBAT22, VBD^{+}22, WNN^{+}22, XSW^{+}22, YDZ^{+}21, ZLS^{+}22]. Deep-Confidentiality [MAK^{+}22]. Defacements [BDM10]. Defeating [HGW07]. Defect [GK23, WCY^{+}23]. Defending [BRK04]. Defense [GCK^{+}22, LMS^{+}21, EL09]. Defined [WQC^{+}19, YLZ^{+}21]. Degree [SGOS19]. Delay [DZHV16]. Delegatable [PO19]. deletion [FLL06]. deliberation [VDV18]. Deliberative [LPB^{+}17]. Delivery [BCGN16, TMK^{+}12, WMW^{+}22, HNF^{+}05]. Delivery-Centric [TMK^{+}12]. Demand [KAS14]. Demand-Invariant [KAS14]. Democracy [LPB^{+}17]. Dense [GAL^{+}22]. Density [RMP10]. Dependencies [CSMM17]. dependent [MS05, WL07]. Depletion [AKA^{+}23]. Deployment [BLMP20, TGBG20, WK18, MBC^{+}05]. Deployments [EDC20, VSID16]. depth [JMSP06]. Derivation [CWW^{+}21]. Description [NGER20]. Descriptions [NGER20]. Descriptor [LZJ^{+}21]. Design [AOVP08, DOG^{+}22, DJ15, FXYX23, KKK21, MAM03, OWK^{+}19, PCP^{+}20, SK17, SS06, ZXH16, BC01, BCF^{+}07, DRJ^{+}07, FT02, MBC^{+}05, Zdu08, HZCS10]. Designing [CBM23]. Detect [CLL23, MMP^{+}14, ZTH^{+}23]. Detecting [BC23, CDM10, GK23, PSA^{+}20, PDAMGULMV20, RM17, YZL^{+}24, YLZ^{+}21].

Detecting-based [YZL^{+}24]. Detection [ABR17, AAF18, ACDLM19, BDM10, CBM23, CPL^{+}21, CS09, FPR16, LMSS23, LXC^{+}13, MSW^{+}16, MEAK^{+}21, MHA^{+}21, OKM21, PSZ24, SAJL16, SR13, SZT22, VBD^{+}22, WARCD17, WCY^{+}23, WDK^{+}24, XM17, YLL^{+}17, YYM^{+}19, ZLZ^{+}23, ZOC11, ZHL^{+}16, ZSL^{+}17, ZMT^{+}23, CDM10, GNK11]. detectiOns [CMTT24]. determine [GMM09]. Developing [AJ03, CBM23, AGPS05]. Development [BTC^{+}23, CDC14, SH22, ZZF^{+}23, BCF^{+}07, CDMF07, MAM03, OSSV05]. Device [ABCL17, JS13, LGGB^{+}21, RAR22]. Devices [AKA^{+}23, CLM19, FGS20, FMC19, HSRK23, JLC20, STB^{+}19, SZT22, SST^{+}16, TSY^{+}21, YBMV22, ZKC^{+}22, DMT07]. DevOps [SCL^{+}19, XvHWW18]. Diagnosis [LPX^{+}21, NZQX22, SPE^{+}22, ZJQ^{+}21]. Dialogical [LSK^{+}17a]. Dialogue [LWH^{+}21]. Differences [XM17, LYW^{+}05]. Differential [LP21]. Diffusion [ZHL^{+}16]. Digital [ALS23, CWC14, PAS13, RCM^{+}22, STJ^{+}21, WCY^{+}23, YZY^{+}14]. Dimensional [KLS^{+}17, RIB18, YSW^{+}17]. Dimensionality [CSMM17]. Direct [JHC^{+}22]. directed [KLMH03]. Directions [SLPZ23]. disaster [PRKD20]. Disclosure [PVL^{+}17, HTG06]. Discount [XLL20]. Discourse [WS17]. Discourse-Level [WS17]. Discovery [BJ15, DCL^{+}22, GLWH17, KLS^{+}17, PHC^{+}21, AOVP08, BCP08, GLF02, SBG07]. Discrete [LPX^{+}21, SZT22, DRJ^{+}07]. Discrimination [CB15]. Disease [GJAT^{+}21, JGH^{+}22, LLL22, MSG^{+}21, SRK22, XZJO22]. DisguisedNets [CGS23]. Disgust [DP17]. Disorder [VBD^{+}22]. Disputes [KYY17]. Disruptive [ABR17]. Dissecting [CCJ^{+}14]. Dissemination [CLW^{+}22]. Distance [YC18, LLSM08, TJLC08]. Distinguish [MS17]. Distributed [AHM14, AO22, ATB^{+}11, BAM^{+}22, CLFX24, FLLM22, GS17, MMK^{+}22, MA23, PCV^{+}21, RPR22, SCL^{+}19, TGBG20, WK18, ZZW^{+}22, AJP07, GBAR08, JN08, KMW09, LLSL08]. Distribution [AAF18, PT09, BVT06]. Diverse [LZJ^{+}21, PC22]. Diversity [HZ11]. Divisions [YCH^{+}22]. DM2 [MAB19]. DM2-ECOP [MAB19]. DNN [CYWW22, FXYX23, TF21]. DNS [DPD22, RMP10, SK13]. do [CPV03]. document [KRML09]. documentation [GMM09]. Documents [Mor17, STJ^{+}21, CTZZ06, MPC06, YASU01]. Does [TSM^{+}23]. Doge [LMSS23]. DoH [TSM^{+}23]. Domain [Ano15, CL24, LHTL06, LSK^{+}17b, PACH20, Thi05, ZLZ^{+}23, YCM^{+}13]. Domain-Specific [LSK^{+}17b, Thi05]. domains [BYCE07]. Dominance [BBH^{+}14]. DONAS [Ano15]. Double [NT21]. Downtimes [GD17]. Drift [GK23]. Driven [ASW^{+}22, DCZ^{+}21, FCS^{+}18, GNR19, TEMH19, VAKK19, XWML19, YBZ14, BCF^{+}07, CDMF07, CLN05, CDJ^{+}22, KGKK21, MBC^{+}05, Rin09, SF21, TJGY22, WHM^{+}22, XIS22]. Driver [RTR^{+}22]. Drone [SABL24]. Drones [SPCC23, ZXP^{+}22]. Dual [GNW^{+}20, HCW^{+}21, HLLS21, YLL^{+}17]. Dual-layer [GNW^{+}20]. Dual-robust [HLLS21]. Dual-Structured [HCW^{+}21]. Dump [LMSS23]. Duplicate [ZSL^{+}17]. During [MBE22]. DuroNet [HLLS21]. Duty [CMML22]. Duty-cycling [CMML22]. DVE [CLN05]. DWT [KGAR22]. DWT-based [KGAR22]. DxHash [DWF24]. Dyadic [RSS17]. Dynamic [CLF^{+}19, CJW^{+}23, GNW^{+}20, GHD21, LMS^{+}21, MD22, PSP22, RTcR19, ZOC11, ZXYL16, CDIW05, HBGF02]. Dynamics [ABDL14, FAGB14, PWSG22].

E-Commerce [BWL16, DCZ^{+}21, SXZ^{+}21, XLL20, GWF^{+}21, VPR07, ZH09, Ung05]. E-deliberation [VDV18]. E-health [PO19].

E-healthcare [WG23]. easIE [GPM^{+}18]. Easy [GPM^{+}18]. Easy-to-Use [GPM^{+}18]. Eavesdropping [CLS^{+}22]. ECC [MMJ21]. ECH [TSM^{+}23]. Ecommerce [GHD21, MFR^{+}21]. Economic [CWC14, YBZ14]. Economics [BCG^{+}18, CDM^{+}14, XWML19]. Economy [APAC18, BKK03]. ECOP [MAB19]. Ecosystem [YBMV22]. Ecosystems [BG21]. Edge [AZKG21, ACG^{+}11, ATD22, BMG^{+}19, BLMP20, CGG^{+}22, CLF^{+}19, CYG^{+}21, CYWW22, CSW^{+}22, CLFX24, FYW^{+}22, FGS20, FYZ19, GDLM22, GJAT^{+}21, HSRK23, HXB^{+}22, JPCL22, KA20, LDG^{+}23, LZK^{+}22, LPR19, LQVK21, LLL22, MAB19, MDDB19, MD22, STB^{+}19, SF21, SZT22, SKH22, SLG^{+}22, SPCC23, VAS24, WCC20, WWJ^{+}22, XZJO22, XFL^{+}23, YZL^{+}24, ZZF^{+}23, ZGD23, ZLS^{+}22, ZMGW22, GHD21, RKY^{+}22]. Edge-AI [GJAT^{+}21, HXB^{+}22, RKY^{+}22]. Edge-centric [KA20]. Edge-Fog-Cloud [FYZ19]. Edge/Fog [XZJO22]. EdgeCI [CLFX24]. Editor [SSC23]. Editorial [CCM17, MQUXK22, FFGM04a, FFGM04b, GS07b, ML08]. Editors [AGKW14, BCG^{+}18, CDPR17, CGL^{+}14, GNR19, KCR^{+}17, LPR19, TSS19]. education [LLSM08, TJLC08]. EECDN [CYWW22]. Effect [DJ15]. Effective [HNF^{+}05, TF21, WCX^{+}23, MPC06]. Effects [CWLZ19, YWML19, BSS02, Wil02]. Efficiency [BL17]. Efficient [AM03, CYG^{+}21, EDC20, GDLM22, GEFT14, LHL^{+}22, LZK^{+}22, MAB19, MJ22, OGP^{+}18, PK20, PCV^{+}21, PHC^{+}21, SPAT21, SCW17, SKH22, SL22, TJGY22, WCC20, WGW^{+}24, WTS^{+}21, YLC^{+}22, ZXYL16, ZB20, CGG^{+}22, CYWW22, HAST21, JSAA22]. efficiently [CDIW05]. egress [GNK11]. EHA [WG23]. eHealth [PHC^{+}21]. Elastically [DWGC23]. Elasticity [CMTD16, GS17, MD22]. Election [MBE22]. Electric [ASW^{+}22]. electronic [CPV03, MS05]. Elements [FLR23]. Eliciting [GHK17]. Email [SHH^{+}06]. embedded [Thi05]. Embeddings [GVM^{+}23, WMWM20]. Emerging [BCN17, LT16, SRK22, XvHWW18]. Emo2Vec [WMWM20]. Emotion [WMWM20, YYM^{+}19]. Emotional [GRR20, LWH^{+}21, WMWM20]. Emotions [DP17, MS17]. Empathy [OALA17]. Empirical [DvRDHB22, XM17, ZH09]. Empowered [WRWM21, LQSW21, TSY^{+}21]. Enabled [ABC^{+}17, DCZ^{+}21, GAL^{+}22, MAK^{+}22, SGC16, SSA^{+}21, ZWC^{+}22, AHJ^{+}20, AKA^{+}23, FYZ^{+}21, LQVK21, MBC^{+}05, SK24, SS06, WG23, ZKC^{+}22, MA23]. Enabling [BLMP20, BLTH22, KBBI15, MDDB19, RHT20, SDB21, GBAR08]. Encoder [XCRY22]. Encoder-decoder [XCRY22]. Encoding [SCLB24]. Encoding-based [SCLB24]. Encrypted [GWF^{+}21, ZXYL16]. Encryption [RMMH22, STJ^{+}21, TSM^{+}23]. End [BB23, PCBG19, SPKTG22, BC01, CFTV03]. End-to-End [PCBG19, SPKTG22, BB23, BC01, CFTV03]. Energy [AKA^{+}23, ASW^{+}22, BLTH22, CGG^{+}22, CYWW22, GDLM22, JSAA22, KGKK21, LHL^{+}22, MRS^{+}22b, SH22, VSKEOZM22, WMG^{+}21, YLC^{+}22]. Energy-Centric [MRS^{+}22b]. Energy-Efficient [LHL^{+}22, YLC^{+}22, CGG^{+}22, CYWW22, JSAA22]. Engagement [LSK^{+}17a, MBE22]. Engender [YCC17]. Engine [JPSS17, VAKK19, NDL07]. Engineering [MDDB19, YADI02, AR12]. engines [JMSP06, LM04]. English [DRJ^{+}07, HJPB06, LLC^{+}23]. Enhance [SPKTG22, WHM^{+}22]. Enhanced [BCFB18, DTE17, HSLH17, HLLS21, HZB19, KK21]. Enhancement [BCN17, CXH^{+}21]. Enhancing [AO22, MQUXK22, ZLS^{+}22]. Enriched [KLS^{+}17, AKS07]. enroute [LSCZ05]. Ensemble [BC23, CYG^{+}21, KA20]. Enterprise [GSS^{+}14]. Entity [PC22, KMW09]. Entity-centric [PC22]. Entropy [ZJQ^{+}21, ZGF^{+}23]. Entropy-based [ZJQ^{+}21]. Environment [CIY^{+}21, MAB19, PO19, VSKEOZM22, WWZ^{+}23, Var03]. Environments [BCCA^{+}21, CCD^{+}22, GHD21, LPR19, MRB19, MMI23, PAS13, RPR22, SL22, VBD^{+}22, WSLT21, XSW^{+}22, MYS^{+}12, SBG07]. Epidemiological [MGHB16]. Epilepsy [ZJQ^{+}21]. EPRT [PHC^{+}21]. Equal [HZB19]. Equality [Mor17]. Equipping [DMT07, GL14]. ERNIE [ZTH^{+}23]. Error [SABL24]. eRulemaking [LPB^{+}17]. esDNN [XSW^{+}22]. Espionage [LJS^{+}14]. Establishment [BCO13]. Estimating [CGM03]. Estimation [EDC20, JPCL22, MMR16, RMP10]. Estimators [ZOC11]. Ethereum [CLZ^{+}20, CPL^{+}21, PSZ24]. EtherShield [PSZ24]. Ethics [BBP18, VDV18]. Evaluating [BSS02, MBP^{+}17, MPS04]. Evaluation [HZ11, JWW15, YPFY21]. Event [ABR17, ACDLM19, AGKW14, MP14, OKR^{+}14, WARCD17, WEJ14, YLL^{+}17]. Events [HC14, PSL^{+}20]. Everybody [HZB19]. Everything [BCN17]. Evidence [LBC^{+}24]. Evolution [CL24, GLQ11, MMV11, SSKW20, FLL06]. Evolutionary [RWXC20]. Evolved [PDF^{+}23]. Evolving [WFZ^{+}20]. examination [Hoc02, JMSP06]. Exchange [CYG^{+}21, LZW^{+}22, MCS18, ZXH16, LB04]. Execution [OGP^{+}18]. Exfiltration [MEAK^{+}21]. Existing [LDG^{+}23]. Experience [GAL^{+}22, HS19, PDS20, WHM^{+}22]. Experiences [CCN^{+}21, LHTL06]. Experimental [ABC^{+}17, JLC20, GNK11]. Experiments [NDO^{+}17, BRRT05]. expertise [BF06]. experts [BF06].

EXplanations [CMTT24]. Exploiting [AAF18, BCN17, LC12, PK20, SO17, TK11]. Exploring [ALS23, WLL^{+}13]. Exposure [RML12]. Extending [DKP17]. External [LSLY19]. Externalities [GdOW14, Web17]. Extracting [EM19, HNGN23, HHS^{+}22]. Extraction [BWL16, BC23, GPM^{+}18, WL07]. Extractor [MSG^{+}21].

Fabric [JKI^{+}21]. Facade [ADGM23].

Facebook [OALA17]. Facial [GZL^{+}21, XCRY22]. Facilitating [Web17, WYC^{+}23]. Factorisation [De 19]. Factors [LFL17, PVL^{+}17]. factory [GS07a]. Failing [HZB19]. Failures [FPA^{+}23]. fairness [PT09]. Fake [BC23, CLL23]. fall [KSA^{+}10]. False [GRR20]. far [DLLM07]. farm [WY01]. Fast [JDZ^{+}21, KRML09, WGW^{+}24]. Fatigue [CTS^{+}24]. Fault [AHM^{+}15, SCPB22, WEJ14, XFL^{+}23, XZY^{+}21]. Fault-aware [SCPB22]. Fault-Tolerant [WEJ14, XFL^{+}23]. Feasibility [RDC16]. Feature [BC23, KSL^{+}21, LPX^{+}21, MSG^{+}21, YLL^{+}17, Dal11]. Features [JHC^{+}22, LSK^{+}17b, SZT22, WL07]. Federal [MBE22]. Federated [CE21, LZK^{+}22, PSA21, SPG22, WWJ^{+}22, WGW^{+}24, ZLZ^{+}23]. Federation [ALA^{+}19]. federations [Zdu08]. Fees [TNJJ22]. Fighting [GM04]. files [ZHH04]. Filter [GNW^{+}20, Wil02]. Filter-based [GNW^{+}20]. Filtering [HSLH17, PMN23, YSZ^{+}22, ZLL^{+}20, JKR07, KRML09]. Filters [HZB19]. Finance [PWGQ22]. Financed [CDM^{+}14]. financial [LB04]. find [ALG04]. Finding [PSL^{+}20, ZGF^{+}23]. Fine [APAC18, BG21, BDT04, CJW^{+}23, PV17, YZY^{+}14, YYM^{+}19]. Fine-Grained [APAC18, BG21, CJW^{+}23, PV17, YZY^{+}14, BDT04, YYM^{+}19]. Fingerprint [WZB^{+}21]. FinPrivacy [WZB^{+}21]. Firewall [Liu12]. Five [AHJ^{+}20]. fixing [HGW07]. Flash [CH05]. Fleets [ASW^{+}22]. Flexibility [BLTH22]. Flexible [SPJ09, SPG22, YSZ^{+}22]. Flickr [YLL^{+}17]. Flow [GAT^{+}21, MEAK^{+}21, MMV11, WCZ^{+}21, WLW^{+}23]. Flow-based [MEAK^{+}21]. Flows [NDO^{+}17, PSP22]. Fog [AKOB^{+}21, CCN^{+}21, CLM19, EDC20, FGS20, FYZ19, FMC19, HAST21, LOD19, LPR19, MRB19, MDDB19, PML^{+}19, PBL^{+}22, VASD19, XZG^{+}22, XZJO22]. Fog-Based [CLM19, FMC19]. Fog-cloud [HAST21]. Footprint [VSKEOZM22]. Footprints [YZY^{+}14]. Force [ZTL^{+}21]. Forecasting [DCD^{+}21, Glu10, JGH^{+}22, PGP^{+}21]. Forgiveness [BL17]. Form [Mor17]. Formal [AVB17, MLMK05]. Formation [DGWW15, RSS17, YC18]. Formats [HHS^{+}22]. Foundations [CAV14, KBBI15]. Fourier [PWSG22]. fragment [CDIW05]. fragment-based [CDIW05]. Framework [AE24, BB23, BTGM22, BDM10, CDJ^{+}22, CMTD16, ISG^{+}22, JG10, KGKK21, KKMK16, KSL^{+}21, KGAR22, LDG^{+}23, LYW23, MKJB21, MGB^{+}21, MAK^{+}22, MA23, RWXC20, RZAD17, SST^{+}16, SCZ^{+}21, TSY^{+}21, VSID16, WCZ^{+}21, WSM21, WHM^{+}22, YDZ^{+}21, ZKC^{+}22, ZSY^{+}17, ZZF^{+}23, GBAR08, LLNF12, TPK10, Van08]. free [BVT06, YJL^{+}22]. Frequency [GLFV^{+}21, CGM03]. Frisber [RCP^{+}15]. Function [WCC20]. functions [ABMW05]. Fusion [ABCL17, KGAR22, WWZ^{+}23]. Future [SLPZ23, SD12]. Fuzzy [BBH^{+}14, JCH^{+}18, JGH^{+}22, YLC^{+}22, ZH09].

Gait [YLM^{+}23]. Game [PHR^{+}21, YJL^{+}22, YC18, YLC^{+}22, LZW^{+}22]. Game-Based [YLC^{+}22]. Game-Theoretic [PHR^{+}21]. Games [BKS^{+}14, DABP14, WYC^{+}23]. GAN [FYZ^{+}21]. Gap [ZLHD15]. Gaps [SPM^{+}13]. Gas [MRY^{+}23, MRY^{+}23]. Gateway [PCV^{+}21]. Gateway-based [PCV^{+}21]. Gathering [ACG^{+}11, JMSP06].

GDWCN [BBS21]. GDWCN-PSO [BBS21]. Generalized [CKKK14, SO17]. Generating [AKS07, MRY^{+}23]. Generation [CGT^{+}21, LWH^{+}21, NGER20, AAA^{+}20, BCP^{+}04, NCEF02]. Generative [WWJ^{+}22, ZZW^{+}22]. Genetic [SK17, SGOS19]. Genres [RM17]. Geo [GS17, MA23]. Geo-Distributed [MA23]. Geo-Elasticity [GS17]. Geographically [GS17]. Geolocation [DPD22]. GEP [DCD^{+}21]. German [MBE22]. globally [GBAR08]. GLV [MMJ21]. goods [HJPB06]. Google [WLL^{+}13]. Governance [GAT^{+}21, KMB^{+}22]. Graded [BBH^{+}14]. Grained [APAC18, BG21, CJW^{+}23, PV17, YZY^{+}14, BDT04, YYM^{+}19]. Graph [ADGM23, BLD^{+}15, CLZ^{+}20, CSS20, CAN^{+}21, CXG21, PWSG22, PLZW18, SR13, ZMT^{+}23, DLLM07, WCZ^{+}21]. Graph-based [CXG21]. Graphical [ADA^{+}22, PPV05]. Green [ADA^{+}22, LZW^{+}22, LLSW22, MRS^{+}22b, SH22, TSY^{+}21]. GREENHOME [VSKEOZM22]. Grid [DLZ^{+}16, LZJ^{+}21, DKP12]. Group [LMSTM14, WJL^{+}22, ZXH16]. Group-Level [WJL^{+}22]. Grouping [SGOS19]. Guarantee [CLJ^{+}21, SKA^{+}23, ZLZ^{+}23]. Guarantees [CKKK14, BLSW04]. Guest [CCM17, FFGM04a, FFGM04b, GS07b, ML08, MQUXK22, SSC23, AGKW14, BCG^{+}18, CDPR17, CGL^{+}14, GNR19, KCR^{+}17, LPR19, TSS19]. guided [YMY^{+}23].

Hadoop [RPA^{+}17]. Hadoop-Based [RPA^{+}17]. Handling [GK23]. hard [ABMW05, LSZ^{+}21]. Hardware [EDC20, MJ22]. Hardware-Accelerated [EDC20]. Harvest [TBG^{+}18]. Hashing [DWF24, LSZ^{+}21]. Hate [PSA^{+}20]. Healing [SBC20]. Health [CSW^{+}22, SPE^{+}22, ZKC^{+}22, PO19].

Healthcare [AKOB^{+}21, SWAHP21, ZTH^{+}23, WG23]. help [SHB06]. Helpfulness [DMGR^{+}17]. Helpfulness-Based [DMGR^{+}17]. Heterogeneous [ALA^{+}19, ADGM23, DCL^{+}22, XCRY22, YLM^{+}23, YSNL16, ZB20, ZDCB18, ZLL^{+}20, AJ03, FFP09]. heuristic [HJPB06]. HICS [RPA^{+}17]. Hiding [RKY^{+}22]. Hierarchical [DSVA19]. hierarchies [Wil02, ZHH04]. High [FYZ19, MRY^{+}23, ZXP^{+}22]. High-order [FYZ19]. High-Quality [ZXP^{+}22]. Highly [LDG^{+}23, WWZ^{+}23]. hijacking [DCAT12]. Histopathology [KSL^{+}21]. Hoc [SLG^{+}22, ZHDD07]. Home [SH22]. Homes [KLS^{+}17, SCLB24]. Honest [BTGM22]. Honey [JPSS17]. Honey-Based [JPSS17]. Hopping [CSS20]. Horizontally [SCPB22]. Hospitals [HML^{+}21]. hosting [USR09]. hosts [CPV03]. Hotspot [NBM19, SPCC23]. hour [GS07a]. Household [VSKEOZM22]. HTTP [Kri01]. Human [BHPY21, CPV^{+}16, CTS^{+}24, CLM19, HS19, LYW^{+}21]. Human-agent [HS19]. Human-AI [CTS^{+}24]. Human-centered [BHPY21]. Human-Robot [LYW^{+}21]. Humidity [RZP^{+}22]. Hybrid [AJSS13, LPX^{+}21, MPR^{+}23, NLLC21, OWK^{+}19, SKH22, YDZ^{+}21]. Hydraulic [WVHTK21]. Hyper [LFL17]. Hyper-Local [LFL17]. Hyperledger [JKI^{+}21]. hyperlink [FS04]. hypermedia [ZHDD07]. Hyperparameter [TF21].

i-DarkVec [GVM^{+}23]. i-Jacob [LYM^{+}18]. IaaS [LC16, ZLHD15]. IBRDM [KGAR22]. ICE [ASW^{+}22]. ICMN [SATPR22]. ICN [FYW^{+}22]. ICT [SRK22]. Identification [CCC^{+}23, NYB^{+}19, RTR^{+}22, SCLB24, WZB^{+}21, HJ08]. identified [QLJ^{+}19]. Identify [MHCF22, Coo03]. Identifying [LHAT22, LFL17]. Identity [TPQC22, XZG^{+}22]. Identity-Based [TPQC22]. IDEs [GL14]. IDN [LHTL06]. IEEE [CMML22]. Image [CGS23, GZL^{+}21, MKJB21, ZJL^{+}15, XZZ08]. Images [HLG^{+}21, KSL^{+}21, MHA^{+}21, YDZ^{+}21]. Imbalance [RTcR19]. Impact [AJP07, GLQ11, WZKP19, YV22, Liu12]. Implementation [KKK21, PCP^{+}20, AOVP08, HZCS10, SS06]. implementations [LYW^{+}05]. Implementing [MBP^{+}17]. Implications [Jor09]. Implicit [NDO^{+}17, YLM^{+}23]. Improve [AAF18, CT17]. Improved [CIY^{+}21, DCD^{+}21]. Improvement [YBW19]. Improving [BL17, CXW^{+}21, GAL^{+}22, OWK^{+}19, XZZ08, YXP^{+}18, YCH^{+}22, ZSY^{+}17, ZLL^{+}20]. in-depth [JMSP06]. Incentive [AAJ21, CWLZ19, DCZ^{+}21, LZW^{+}22, NBM19, HGW07]. Incentive-Based [NBM19, AAJ21]. Incentive-Driven [DCZ^{+}21]. Incentives [CGL^{+}14, SXZ^{+}21]. Inclusion [TNJJ22]. Incompatible [SL22]. Incorporating [BL17]. Incremental [GVM^{+}23, WJL^{+}22]. independent [YV22]. Index [WLB22, ZXYL16]. Index-Based [ZXYL16]. Indexing [CSMM17]. India [DD07]. indicators [HJ08]. indices [LM04]. Individual [DRC^{+}23]. Indoor [KIG^{+}19, WWZ^{+}23]. Industrial [CSS20, DXP^{+}23, FXYX23, JLC20, LZK^{+}22, LLSW22, RAR22, SS20, ZTL^{+}21]. Industry [CLW^{+}22]. Inertial [JHC^{+}22]. Infectious [LLL22, XZJO22]. InFeMo [SPG22]. Inference [CYWW22, CLFX24, FXYX23, HSRK23, MMK^{+}22, NLLC21, SL22, KG10]. Inferring [BPSD17, GH06]. Influence [CDM^{+}14, IDS19, LGC20, WFZ^{+}20, ZLL^{+}20]. Influencers [RM17]. Influences [HS19]. Information [ABCL17, CSW^{+}22, FSC15, GRR20, GPM^{+}18, LFL17, MPR^{+}23, NZQX22, NYB^{+}19, SO17, TK11, WMW^{+}22, YSW^{+}17, YYM^{+}19, ZLL^{+}20, BKK03, HTG06, JMSP06, Rin09, WL07].

Information-centric [MPR^{+}23]. Infrastructure [BBC14, AGPS05]. Infrastructures [CGT^{+}21, CRP17, OKM21, ZB20, DKP12]. ingress [GNK11]. Initial [PAS13]. Innovation [CB15]. Inpainting [MKJB21]. Inputs [MRY^{+}23]. Inquiries [PDF^{+}23]. Insider [LJS^{+}14]. Inspection [CHC^{+}21]. Instagram [WL23]. Instance [CXG21, MS05]. Instrumentation [GEFT14]. Integrated [CGG^{+}22, FYZ19, GDLM22]. Integrating [LSLY19, DFL^{+}23, VJL^{+}14, YSW^{+}17]. Integration [LPR19, XZY^{+}21, CS09, ZXS08]. Integrity [RQL^{+}21, JKS^{+}10]. Intelligence [AHJ^{+}20, AE24, ACG^{+}11, ABC^{+}17, IRJ^{+}21, LLL22, XZJO22]. Intelligent [AKOB^{+}21, CGG^{+}22, GDLM22, HML^{+}21, KKK21, KK21, KA20, KZLG21, KGAR22, LWM^{+}21, RPA^{+}17, SAJL16, WCY^{+}23, YLM^{+}23, AM07, CS07]. Intensive [LYM^{+}18, ETRDRO^{+}19, MAM03]. Intent [WHM^{+}22, ZTH^{+}23]. Intent-driven [WHM^{+}22]. Intentional [FYT17]. Inter [ZLZ^{+}23]. Inter-domain [ZLZ^{+}23]. interacting [JMSP06]. Interaction [CDPR17, LYW^{+}21, MGB^{+}21, NPP^{+}15, SL22, YXP^{+}18]. Interaction-Based [NPP^{+}15]. Interactions [YCC17]. interactive [KMW09]. interdomain [GNK11]. Interest [GCP^{+}20, GLWH17, YSNL16, HMLH21]. Interest-Aware [GLWH17]. interesting [Coo03]. Interfaces [OWK^{+}19, ZSY^{+}17, PPV05, SNBC12]. Intermediate [Glu10]. Intermittently [SATPR22]. International [FYT17]. Internationalized [LHTL06]. Internet [MFR^{+}21, AM03, AJP07, ADA^{+}22, BLSW04, BHPY21, BCMS06, BCN17, BCGN16, BSS02, BI17, BC01, BTC^{+}23, BRK04, CLW^{+}22, CFTV03, CZPS22, DWGC23, DD07, DNJ19, FFGM04a, FFGM04b, GCK^{+}22, GS07b, GS07a, GBAR08, HC14, HAD22, IRJ^{+}21, JKR07, Jor09, LLSM08, LLSL08, LNTL23, Liu20, LZK^{+}22, LS21, LP21, LLSW22, MGHB16, MRS^{+}22b, MRS^{+}22a, MJ22, MMV11, PC22, PT09, PML^{+}19, RMMH22, SSA^{+}21, SD12, SCZ^{+}21, TSY^{+}21, TSS19, TPQC22, TGBG20, USR09, Var03, WQC^{+}19, WNN^{+}22, Web17, WRWM21, XvHWW18, XWML19, YZL^{+}24, YSNL16, YCH^{+}22, YLZ^{+}21, ZTL^{+}21]. Internet-based [AJP07, BRK04, XvHWW18, DNJ19]. Internet-of-things [GCK^{+}22]. Internet-of-Vehicles [TPQC22]. Internet-scale [PT09]. Internetware [LYM^{+}18, XvHWW18, OGP^{+}18]. Internetware-Oriented [LYM^{+}18]. Interpret [LPB^{+}17]. Interpretation [LMZ13]. interval [PSZ24]. Intra [XSSD23]. Intra-Shard [XSSD23]. Introduction [AM07, AGKW14, BHPY21, BCG^{+}18, CGT^{+}21, CAV14, CSS17, CZPS22, CDPR17, CGL^{+}14, DNJ19, FLLM22, GDLM22, GNR19, HAD22, KCR^{+}17, LLSM08, LPR19, LWFD21, MQUXK22, MBB07, NBFZ15, PBJP21, SD12, SSC23, SWAHP21, SSKW20, TSS19, XZJO22, XvHWW18, ZBF^{+}19]. Intrusion [OKM21, SAJL16, WDK^{+}24]. Intrusions [AAF18]. Invariant [KAS14, WL07]. Inverted [ZXYL16]. Investigating [GJAT^{+}21, SPM^{+}13]. Investigation [TNJJ22, ZH09]. Invocations [WZKP19]. IoMT [WG23]. IoMT-based [WG23]. IoT [BCGN16, AE24, AKA^{+}23, AAA^{+}20, BB23, BLMP22, CE21, CIY^{+}21, CGG^{+}22, CXG21, CMML22, FXYX23, FGS20, FFKG19, FMC19, GDLM22, JLC20, JSAA22, KLS^{+}17, KKMK16, KIG^{+}19, LYW23, LLC^{+}23, LGGB^{+}21, LGKL20, LQSW21, LQVK21, MED19, MAK^{+}22, Nov19, PACH20, PCV^{+}21, PGP^{+}21, QZDG22, RKY^{+}22, RPA^{+}17, RAR22, RPR22, STB^{+}19, SF21, SGC16, SBC20, SSC23, SH22, SST^{+}16, SSKW20, SL22, TEMH19, UNBAT22, VASD19, VSID16, WCX^{+}23, WDK^{+}24, YBMV22]. IoT-Based [FFKG19, KKMK16, MED19, SH22, CE21, UNBAT22]. IoT-Edge [LQVK21]. IoT-Enabled [MAK^{+}22, SGC16]. IoT-Enriched [KLS^{+}17]. IoT-oriented [JSAA22]. IoTs [SAJL16]. IoTvar [BTC^{+}23]. IoV [JHC^{+}22, ZWC^{+}22]. IoVs [XZG^{+}22]. IP [DPD22, EL09, Nov19]. IP-Agnostic [Nov19]. IPFS [LYW23]. IPv6 [ATS^{+}21, ZW17]. IRGA [YLM^{+}23]. Irony [FPR16]. isotonic [JKR07]. ISP [DJ15, JS13]. Issue [BBP18, BCG^{+}18, CGT^{+}21, CAV14, CSS17, CZPS22, CGL^{+}14, GNR19, KBBI15, LPR19, MBE22, MFR^{+}21, SSC23, SSKW20, TSS19, XvHWW18, LLSM08, MBB07, SD12]. Issuers [GAC18]. Issues [LLL22, CPV03]. Item [GLFV^{+}21]. Item-specific [GLFV^{+}21]. Iterative [NT21].

Jacob [LYM^{+}18]. Jamming [CLS^{+}22]. Jamming-Aided [CLS^{+}22]. JavaScript [FLR23]. JCloudScale [ZLHD15]. Job [KGKK21]. Johnny [KSA^{+}10]. Joint [FXYX23, HAST21, STJ^{+}21]. Juno [TMK^{+}12].

Kautz [GLJ^{+}12]. Kernel [GZL^{+}21]. Key [BCO13, NYB^{+}19, ZXH16, DMT07]. Key-Establishment [BCO13]. Keypoints [XCRY22]. Keypoints-based [XCRY22]. Keyword [CWW^{+}21, LC12, ZXYL16]. Knowledge [ADGM23, ETRDRO^{+}19, GNR19, GHK17, LSLY19, QLJ^{+}19, ZSY^{+}17, GS07a, WL07]. Knowledge-Driven [GNR19]. Knowledge-intensive [ETRDRO^{+}19]. KRDB [GR04]. Kubernetes [ZB20]. Kubernetes-Based [ZB20].

L2DART [DFL^{+}23]. Labeling [NZQX22]. LAKE [BCO13]. Lanczos [FYZ19].

Landing [SABL24]. Language [CT17, LYW^{+}21, NLLC21, PDAMGULMV20, XIS22, YV22, HP03, MLMK05, Thi05]. Language-independent [YV22]. languages [MLMK05]. Large [BDM10, DRC^{+}23, GNW^{+}20, PK20, TSM21, VSID16, ZHL^{+}16, AKR01, JKS^{+}10]. Large-Scale [BDM10, DRC^{+}23, PK20, TSM21, VSID16, GNW^{+}20, JKS^{+}10]. Latency [EDC20, MRB19, MEAK^{+}21, SKA^{+}23]. Latency-Aware [MRB19]. Laughing [MBP^{+}17]. Layer [GZL^{+}21, MQB22, ZXH16, GNW^{+}20]. Layer-based [MQB22]. Layouts [JYW^{+}19]. Leak [ZLZ^{+}23]. Leakage [STK17]. Learning [ASO^{+}22, ALG04, AZKG21, CE21, CYG^{+}21, CLS^{+}22, CLL23, CL24, CXG21, DXP^{+}23, DRC^{+}23, HCW^{+}21, HSRK23, HXZ^{+}20, HLLS21, HLG^{+}21, KZLG21, LWH^{+}21, LXZ^{+}22, LSK^{+}17b, LZK^{+}22, LLSW22, MMK^{+}22, MQUXK22, MRS^{+}22b, MSG^{+}21, MMJ21, MS05, PSA21, PGP^{+}21, RTR^{+}22, RWXC20, RZP^{+}22, SABL24, SSA^{+}21, SPE^{+}22, SZT22, UNBAT22, VBD^{+}22, WMWM20, WNN^{+}22, XLL20, YDZ^{+}21, ZSY^{+}17, ZLS^{+}22, DSNK08, FFGM04a, FFGM04b, LLSL08, SMFR08]. Learning-Based [WNN^{+}22, ZSY^{+}17, HXZ^{+}20, SSA^{+}21]. Learning-powered [LLSW22]. Least [TSM21]. Level [WS17, WJL^{+}22, CH05, LSLY19, ZMT^{+}23]. levels [DRJ^{+}07]. leverage [GS07a]. Leveraging [YSNL16, YXL^{+}21]. Light [SK24]. Lightweight [CIY^{+}21, JDZ^{+}21]. like [JDZ^{+}21]. Limb [KKK21]. Line [JHC^{+}22]. Linguistic [DRC^{+}23, OALA17]. Linguistics [SSC23]. Link [BRRT05, EV07, FLL06, NCEF02, ZHDD07, ZHH04]. LinkSelector [FS04]. Literature [GLF02, PSA21]. Literature-based [GLF02]. Live [PSL^{+}20, VAKK19, TJLC08].

Living [HXB^{+}22, LQVK21, RKY^{+}22, SPE^{+}22, VBD^{+}22]. Load [CLM^{+}11, DCD^{+}21, DABP14, WY01]. Load-Balancing [DABP14]. Local [ACDLM19, CSMM17, LFL17]. Locality [GZL^{+}21, TJLC08]. locality-aware [TJLC08]. Locality-Constrained [GZL^{+}21]. Localization [YLC^{+}22]. Location [Var03, YSW^{+}17, BCMS06]. location- [BCMS06]. locator [BF06]. log [ZHH04]. Logic [GAL18]. Logs [ACDLM19]. Lossless [RKY^{+}22]. Low [AKA^{+}23, AAA^{+}20, BCO13, FMC19, SAJL16, ZZF^{+}23, MEAK^{+}21]. Low-code [ZZF^{+}23]. Low-cost [AAA^{+}20]. Low-Power [AKA^{+}23, SAJL16, FMC19]. LQR [SK24]. LSTM [HML^{+}21]. LTE [SGC16].

Machine [ASO^{+}22, JSAA22, MRS^{+}22b, MMJ21, RZP^{+}22, SSA^{+}21, FFGM04a, FFGM04b]. main [Ji02]. maintain [KMW09]. Maintainable [LJLN16]. Maintaining [LC12]. makes [LYW^{+}05]. Making [ASÖY23, CLJ^{+}21, Nov19, SAB^{+}18]. Malicious [CCC^{+}23, PSZ24, YLZ^{+}21]. Malware [QZDG22]. Manageability [MED19]. Management [AHM^{+}15, ATD22, BB23, BCCA^{+}21, CDJ^{+}22, EHY19, FFKG19, FCS^{+}18, GNR19, JG10, JS13, KBBI15, MRB19, MGB^{+}21, MED19, PPDG19, RAR22, RZAD17, DFL^{+}23, SCPB22, SPG22, TSS19, TK11, WMG^{+}21, WLB22, WHM^{+}22, ATB^{+}11, Ji02, JN08, JAT^{+}06, KS07, Var03]. Managing [NDO^{+}17]. Mandarin [LLC^{+}23]. MANDOLA [PSA^{+}20]. MANET [SPAT21]. Manipulation [LBC^{+}24]. Manipulations [LMSS23]. Manufacturing [ALS23, DXP^{+}23, FPA^{+}23, SLPZ23, WCY^{+}23]. Map [RQL^{+}21]. Mapping [ZXS08]. mappings [CS09]. MapReduce [KGKK21]. Maps [LZJ^{+}21].

Market [BGL04, KAS14, LBC^{+}24, MMI23, PWGQ22, TPK10]. Market-based [BGL04, TPK10]. Marketplace [BL17, CPV^{+}16]. Markets [BLTH22, GHK17, UNBAT22, YWML19]. Markov [DK04]. MARSA [CPV^{+}16]. mash [GMM09]. mash-ups [GMM09]. Mashup [CDC14, RDC16]. Match [WYC^{+}23]. Match-based [WYC^{+}23]. Matching [HC14, LYW^{+}21, TJGY22, ZWW^{+}23]. Materialized [LC12]. Matrix [De 19]. Matter [HHF^{+}21]. Maximization [LGC20, WFZ^{+}20]. Maximize [MGHB16]. Maximizing [HSRK23]. MCEP [OKR^{+}14]. Me [OALA17]. Measure [DABP14]. Measurement [CSW^{+}22, CCJ^{+}14, RZAD17, WLB22]. Measurements [DTE17, GD17, HTG06]. Measures [DMGR^{+}17, PRD09]. Measuring [BZVS18, CFTV03, ETRDRO^{+}19, TBG^{+}18, VDV18]. MEC [CLS^{+}22, LHL^{+}22, ZWC^{+}22]. MEC-Based [CLS^{+}22]. MEC-Enabled [ZWC^{+}22]. Mechanism [AAJ21, ATS^{+}21, BL17, CLF^{+}19, CWLZ19, CAN^{+}21, LZW^{+}22, RQL^{+}21, WZB^{+}21]. Mechanisms [BLMP20]. Media [CCD^{+}22, CDPR17, FAGB14, GRR20, GLT17, HLG^{+}21, LBC^{+}24, MBE22, MS17, WARCD17, YZY^{+}14, Dal11, LCKN05]. mediation [MGB^{+}07]. Mediator [KK21]. Mediator-based [KK21]. Medical [BBS21, LP21, PHC^{+}21, PSA21, SWAHP21, WSLT21, YDZ^{+}21, ZJL^{+}15]. medoid [ZJQ^{+}21]. Meets [WLL^{+}13]. Memory [DWF24, LSZ^{+}21, ABMW05, Ji02]. memory-bound [ABMW05]. Memory-hard [LSZ^{+}21]. Memory-saving [DWF24]. Mental [CSW^{+}22]. mergers [BSS02]. Merging [LZJ^{+}21]. Mesh [SLG^{+}22]. Mesh-based [SLG^{+}22]. Message [MJ22, ZWC^{+}17]. Messages [HHS^{+}22]. Metaheuristics [JDZ^{+}21].

Metering [VSKEOZM22]. Method [ADGM23, GK23, LXZ^{+}22, RMMH22]. Methodology [HCBRM23, SF21, SATPR22]. Methods [MS17, NGER20, LSCZ05]. Metric [XM17]. Metrics [GAC18]. Metro [CWC14, TF21]. Microcomputations [KFB^{+}14]. Micropayments [KFB^{+}14]. Middleware [BTC^{+}23, MDDB19, TMK^{+}12, BCMS06, Zdu08]. Migrating [SAB^{+}18]. Migration [BLMP20, CLF^{+}19, RWXC20]. Mimic [LMS^{+}21]. Mimicry [OALA17]. Miners [TNJJ22]. Minersoft [DKP12]. Minimal [LDG^{+}23, WVHTK21]. Minimize [RTcR19]. Minimum [GLFV^{+}21]. Mining [GLFV^{+}21, LT16, NDL07, RDC16, SF21, WTS^{+}21, YZY^{+}14, ZGB18, EV03, FS04, WL07, ZHH04]. Misogyny [PDAMGULMV20]. Mist [SSA^{+}21, VASD19]. Mitigate [CTS^{+}24]. Mitigating [HSLH17, WZKP19]. mitigation [CH05]. Mixing [LLC^{+}23]. Mobile [ASO^{+}22, AZKG21, ATS^{+}21, AJSS13, BMG^{+}19, BAM^{+}22, BZVS18, DZHV16, GHD21, LOD19, LYM^{+}18, LZBN17, LZW^{+}22, MAB19, NZ22, PACH20, PGT^{+}18, PVL^{+}17, PCBG19, SATPR22, SDB21, VAS24, WCC20, XFL^{+}23, ZWC^{+}17, ZLS^{+}22, ZDCB18, ZMGW22, ZJL^{+}15, BCMS06, CPV03, SMFR08, Var03, PDS20]. Mobile-Edge-Cloud [BMG^{+}19]. Mobility [OKR^{+}14]. Mobility-Aware [OKR^{+}14]. mode [STB^{+}19]. Model [ASO^{+}22, AKOB^{+}21, AO22, BMG^{+}19, BCF^{+}07, CDMF07, CBM23, CGS23, CLFX24, CDC14, CO16, CGL^{+}16, FCS^{+}18, MBC^{+}05, MBS19, RKY^{+}22, RTR^{+}22, RDC16, SPM^{+}13, WS17, WSLT21, YC18, YBZ14, ZTH^{+}23, FLL06, GMM09, ZXS08]. Model-Based [CO16]. Model-Driven [FCS^{+}18, YBZ14, BCF^{+}07, CDMF07, MBC^{+}05]. Modeling [AVB17, PAS13, VJL^{+}14, YXP^{+}18, SHH^{+}06]. Modelling [ISG^{+}22, SCPB22, SWD15].

Models [AR12, CLL23, KA20, SABL24, WWJ^{+}22, DK04, KG10, MBBW07]. Moderately [ABMW05]. Modern [BG21, FT02]. Module [MRB19]. Monitoring [CE21, CSW^{+}22, LZBN17, PK20, PSA^{+}20, PSL^{+}20, WVHTK21, ZKC^{+}22, AJP07]. Monocular [JHC^{+}22]. Moral [DP17, VDV18]. Motion [CLN05]. Motivators [HTG06]. mouse [CLN05]. mouse-driven [CLN05]. Movie [WL23]. Moving [GCK^{+}22]. MPARS [PDS20]. MRI [KGAR22]. Multi [AE24, BJ15, BCCA^{+}21, BC23, CCD^{+}22, DOG^{+}22, FCS^{+}18, HCW^{+}21, HSRK23, JGH^{+}22, KLS^{+}17, LSLY19, LZW^{+}22, MMK^{+}22, MEAK^{+}21, MAB19, RMMH22, RIB18, SCLB24, STK17, SCL^{+}19, WMWM20, WCZ^{+}21, WLW^{+}23, WK18, WSLT21, YSW^{+}17, ZJQ^{+}21, ZWW^{+}23, AGPS05]. Multi-Agent [STK17, MEAK^{+}21, AGPS05]. Multi-Attribute [BJ15]. Multi-Cloud [FCS^{+}18]. Multi-cloudlet [MAB19]. Multi-criteria [DOG^{+}22]. Multi-Dimensional [KLS^{+}17, RIB18, YSW^{+}17]. Multi-Emotion [WMWM20]. Multi-graph [WCZ^{+}21]. Multi-level [LSLY19]. Multi-medoid [ZJQ^{+}21]. Multi-Objective [WK18, BCCA^{+}21, SCL^{+}19]. Multi-Party [WLW^{+}23]. Multi-resident [SCLB24]. Multi-service [LZW^{+}22]. Multi-Task [HCW^{+}21, JGH^{+}22, MMK^{+}22]. Multi-Tenancy [HSRK23]. Multi-Threshold [WSLT21]. Multi-Tier [RMMH22]. Multi-turn [ZWW^{+}23]. Multi-type [BC23]. Multi-user [CCD^{+}22, MAB19]. Multi-vector [AE24]. Multi-view [ZJQ^{+}21]. Multiagent [CZPS22]. Multicast [SLG^{+}22]. Multicloud [AV16]. Multidevice [DPCM16]. multidimensional [PRD09].

Multifaceted [VJL^{+}14]. Multilateral [JKI^{+}21]. Multilayer [QZDG22]. Multilevel [PLZW18]. Multimedia [AdM^{+}13, ADA^{+}22, SSC23, SMFR08, LLC^{+}23]. Multimedia-based [ADA^{+}22]. Multimedia-IoT [LLC^{+}23]. Multimodal [HML^{+}21, HLG^{+}21, YLL^{+}17]. Multiobjective [AV16, BBS21]. Multiparty [MPR^{+}23, NT21]. Multiple [CXG21, PHR^{+}21, RM17, WLL^{+}13, XZG^{+}22, AJ03, HJPB06]. Multivariate [XSW^{+}22]. multiversion [CTZZ06]. Mutual [LXZ^{+}22]. MWPoW [XSSD23]. Myths [LFL17].

Naïve [MBS19]. Nakamoto [RZJ20]. Name [Ano15, TSM^{+}23, YCM^{+}13, HBGF02, LHTL06]. NAT [Nov19]. national [BYCE07, GS05]. native [ZZF^{+}23]. Natural [CT17, NLLC21, XIS22]. Navigable [YC18]. Navigation [GCP^{+}20, KIG^{+}19, PHR^{+}21, CLN05, ZHH04]. navigational [EV07]. Nearcast [TJLC08]. Need [PMFS17]. Needs [XWML19]. Negative [CSW^{+}22]. Negotiating [CGL^{+}16]. negotiations [MS05]. Net [CB15, Jor09]. Network [AHS14, ALA^{+}19, ACG^{+}11, BGK14, BLMP20, BLMP22, BKS^{+}14, BG21, CCC^{+}23, CWLZ19, CPL^{+}21, CHC^{+}21, CSMM17, DCL^{+}22, DFLT22, FYT17, GdOW14, HLLS21, HLG^{+}21, HMLH21, LJLN16, LDG^{+}23, NLLC21, PWSG22, PRKD20, PWGQ22, QLJ^{+}19, SGC16, SATPR22, SLBD20, SCW17, TJGY22, WCZ^{+}21, WNN^{+}22, WLB22, WTS^{+}21, WDK^{+}24, XCRY22, XSW^{+}22, YV22, YWML19, ZHL^{+}16, GLJ^{+}12, HZCS10, BVT06]. Networked [LJG18, PWSG22, Gel09]. Networking [MPR^{+}23, PSP22, SSKW20, YPFY21]. Networks [ATS^{+}21, ABCL17, AAA^{+}20, ABDL14, Ano15, AJSS13, BCFB18, BPSD17, CYD^{+}20, CYWW22, CSS20, CRP17, CO16, CGL^{+}14, DGWW15, FLLM22, GNW^{+}20, GAL^{+}22, GLWH17, JPCL22, JWW15, KKY18, KYY17, LWFD21, LQW21, MEAK^{+}21, MHA^{+}21, MMV11, DMGR^{+}17, MD22, NBFZ15, PK20, RCP^{+}15, SK17, SKA^{+}23, SS20, SPKTG22, SLG^{+}22, WNN^{+}22, WJL^{+}22, WLB22, VAK17, WFZ^{+}20, YC18, YMY^{+}23, YLC^{+}22, ZWC^{+}17, ZZW^{+}22, ZGF^{+}23, ZMT^{+}23, ZLL^{+}20, ZJL^{+}15, DSNK08, GH06, KG10, LSCZ05, PT09]. Networks-The [YC18]. Neural [MHA^{+}21, NLLC21, PWSG22, PWGQ22, WWJ^{+}22, XSW^{+}22]. Neutrality [CB15, CDM^{+}14, Jor09]. News [CLL23, GRR20]. Next [AAA^{+}20, CGT^{+}21, HMLH21, BCP^{+}04]. Next-generation [AAA^{+}20, BCP^{+}04]. NIST [SS06]. NLoS [WWZ^{+}23]. NLUBroker [XIS22]. Nobody [HZB19]. Nodes [ZWC^{+}22]. Nonneutral [AHS14]. Normative [KBNV18]. Novel [BBS21, CL24, CMML22, JYW^{+}19, KSL^{+}21, LSZ^{+}21, MKJB21, PPDG19, SPAT21, SPKTG22, WLB22, WG23, WSM21, WYC^{+}23]. Novelty [HZ11].

Obfuscation [ABCL17].

Obfuscation-Based [ABCL17]. object [Zdu08]. Objective [WK18, BCCA^{+}21, SCL^{+}19]. objects [SMFR08]. Obscene [LXC^{+}13]. Observation [WQC^{+}19]. observations [CH05]. ODIN [ABCL17]. Odometry [JHC^{+}22]. Off [AHS14, DFL^{+}23]. Off-Chain [DFL^{+}23]. Off-Network [AHS14]. Offensive [RCP^{+}15]. Offering [PDF^{+}23]. Offloading [ADAP19, DCZ^{+}21, GAL^{+}22, LHL^{+}22, MRS^{+}22b, MAB19, YZL^{+}24, ZWC^{+}22, ZDCB18]. offs [AOVP08]. offshore [AJP07]. offshored [DD07]. On-Device [RAR22]. One [DCAT12]. One-time [DCAT12]. Online [ASBH^{+}16, ALA^{+}19, BGK14, BPSD17, BL17, BKS^{+}14, CCM17, HTG06, JWW15, KKY18, KYY17, LPB^{+}17, LXC^{+}13, NPP^{+}15, PSA^{+}20, RIB18, RM17, RZAD17, SCL^{+}19, VAKK19, WJL^{+}22, WYC^{+}23, YWML19, ZDCB18, Guo02, JKS^{+}10, LYF^{+}09]. Ontology [LMSTM14, Rin09]. Ontology-Based [LMSTM14]. ontology-driven [Rin09]. Oops [STB^{+}19]. Open [MMI23, WDK^{+}24, BCP^{+}04]. OpenStack [BLMP22, MDDB19]. OpenStack-based [MDDB19]. Operating [LWM^{+}21]. Operation [STB^{+}19]. Operation-mode [STB^{+}19]. Operational [AE24]. Operations [CTS^{+}24, PRKD20]. Operator [GEFT14]. Opportunistic [BI17, XFL^{+}23, ZWC^{+}17]. Opportunities [DFLT22, KMB^{+}22, LWM^{+}21]. Optimal [CYD^{+}20, DRJ^{+}07, LSCZ05, MRS^{+}22b, Guo02]. Optimally [SBC20]. Optimisation [SCL^{+}19]. Optimization [AV16, ASW^{+}22, DFLT22, LHL^{+}22, LLSW22, SZT22, TF21, WK18]. Optimization-Based [DFLT22]. Optimize [SK24, XLL20]. Optimized [RTR^{+}22]. Optimizing [LM04, LYM^{+}18, PGT^{+}18, STB^{+}19, TNJJ22, TK11]. Options [RML12]. Orchestration [ZB20]. Order [MP14, FYZ19]. Organizational [GSZ^{+}23]. Oriented [LYM^{+}18, BCP08, JSAA22, LXW^{+}12, Van08, Zdu08, ML08]. Other [DP17]. Other-Condemning [DP17]. OTI [AE24]. OTI-IoT [AE24]. Out-of-Gas [MRY^{+}23]. Out-of-Order [MP14]. Outcomes [KAS14]. Outdoor [PDS20]. Outlook [Liu20]. Outreach [DKP17]. Outsourcing [CGS23, GS07b, XCL07]. overbooking [USR09]. Overexposure [LGC20]. Overexposure-Aware [LGC20]. overhead [JAT^{+}06]. overload [SHB06]. OWL [ZXS08].

P [Ano15, CLM^{+}11]. P-DONAS [Ano15]. P-Ring [CLM^{+}11]. P2P [Ano15, BJ15, CLM^{+}11, TJLC08]. P2P-Based [Ano15, BJ15]. PaaS [ZLHD15]. Packet [SPAT21]. PADUA [MMP^{+}14]. Page [XM17, DK04, THS06]. PageCluster [ZHH04]. PageRank [BGS05, Bri06]. Pages [DCL^{+}22, CDM10, LXW^{+}12]. PANOLA [UY22]. Parallel [MMP^{+}14]. Parameter [SS20]. Paris [CWC14]. Parked [ZMGW22]. Parked-vehicle-assisted [ZMGW22]. Parking [PGP^{+}21]. Parkinson [LPX^{+}21, MSG^{+}21]. Participation [LFL17, VDV18]. Particle [SZT22]. Partitioning [CLFX24, FXYX23]. Party [MHCF22, WLW^{+}23, BZVS18, XJ20]. Passenger [GAT^{+}21]. Passengers [TF21]. Passive [CYD^{+}20]. Password [LSZ^{+}21, ZXH16]. Password-Authenticated [ZXH16]. Past [HS19]. Path [SLBD20, DWGC23, YASU01, GL14]. path-based [YASU01]. Pattern [MED19, TNJJ22, Zdu08]. Pattern-based [Zdu08]. Patterns [BPSD17, CDC14, LC16, RDC16, WTS^{+}21, Coo03, EV07, KRML09]. Pay [XWML19]. Payloads [HHS^{+}22]. PCAM [CDJ^{+}22]. PDG [UNBAT22]. PDG-based [UNBAT22]. Pedestrian [XCRY22]. Peeking [RMP10]. Peer [BGK14, GLWH17, RS09, ZHDD07]. Peer-to-Peer [BGK14, GLWH17, RS09, ZHDD07]. Peering [CGL^{+}16]. Peers [SGOS19]. Perceived [PDS20, Dal11]. Perception [CXH^{+}21, QZDG22]. Performance [CCJ^{+}14, ETRDRO^{+}19, JAT^{+}06, LC16, PMN23, RZJ20, CFTV03, HZCS10, KLMH03]. Personal [ASÖY23, CLM19, JKI^{+}21, PVL^{+}17, UY22]. personalization [AKS07, AM07, EV03, EV07, NDL07]. Personalized [ASÖY23, CJW^{+}23, CO16, DRC^{+}23, HJWW20, AGPS05, LYF^{+}09]. Personalizing [BGK14, DSNK08, LLNF12]. Perspective [BKS^{+}14, CSW^{+}22, GHD21, SDB21, GR04].

Perspectives [SPM^{+}13]. Pervasive [PDS20, YPFY21]. phish [KSA^{+}10]. Phishing [CPL^{+}21, CMTT24, CDM10, HJ08, YW10]. Physical [CGT^{+}21, FYZ19, GAT^{+}21, ISG^{+}22, KGKK21, NLLC21, PBJP21, VAK17, BRK04, FYZ^{+}21, LSZ^{+}21, YXL^{+}21]. Placement [CYD^{+}20, VAS24, WCC20]. Planning [AZKG21, LLG22, STK17]. Platform [PSA^{+}20, RMMH22, TMK^{+}12, Hoc02, USR09]. Platforms [CCC^{+}23, PBL^{+}22]. plugged [PP11]. plush [ATB^{+}11]. POI [CJW^{+}23]. Point [HMLH21, JHC^{+}22]. Point-of-interest [HMLH21]. Points [GCP^{+}20]. Points-of-Interest [GCP^{+}20]. Poisoning [YCM^{+}13]. Polarized [YMY^{+}23]. Policies [ZGB18, Ung05]. Policy [BTH^{+}17, DSVA19, MAB19, PV17, Hoc02, Liu12]. Policy-Carrying [PV17]. Policymaking [GAC18]. Polishing [ZTL^{+}21]. politics [Kri01]. Pollution [GJAT^{+}21]. Popular [BWL16]. Popularity [EDC20, FAGB14, WJL^{+}22]. portals [FS04]. Portfolio [JKI^{+}21]. Portlet [DR05]. Positional [SCLB24]. Positioning [WWZ^{+}23]. PoSSUM [PC22]. Post [PRKD20, YCH^{+}22]. Post-disaster [PRKD20]. Post-quantum [YCH^{+}22]. Potential [ALS23]. Power [AKA^{+}23, BZVS18, MMJ21, SAJL16, WMG^{+}21, FMC19]. powered [LLSW22]. PPRP [LLG22]. Practical [FYZ19, RCP^{+}15, SABG17, VDV18, WQC^{+}19, XZY^{+}21]. Practices [JG10]. Pre [MHA^{+}21]. Pre-Trained [MHA^{+}21]. Precise [SABL24]. Predict [ABR17, DMGR^{+}17, TF21]. Predictability [LC16]. predicting [DK04]. Prediction [ASW^{+}22, CLW^{+}22, De 19, GK23, GHD21, HLLS21, HZB19, PMN23, WCZ^{+}21, WNN^{+}22, WJL^{+}22, WLW^{+}23, XCRY22, XSW^{+}22, YXL^{+}21, CLN05]. Predictive [DFLT22, PGP^{+}21, SH22]. Preference [YZY^{+}14, Hoc02, NDL07]. Preference-Aware [YZY^{+}14]. Preferences [BBH^{+}14, LMSTM14, PDS20]. Prefetching [KIG^{+}19, CLN05, LM04]. Premium [CGL^{+}16]. Presence [FYT17]. Preservation [EHY19]. Preserving [ABCL17, CSMM17, KKY18, LLG22, MMK^{+}22, MAK^{+}22, PLZW18, PHC^{+}21, UY22, XCL07, YSZ^{+}22, CE21, CCD^{+}22, FYZ19, PSK10, SLBD20, WZB^{+}21, XZG^{+}22, YDZ^{+}21]. Preserving-Privacy [LLG22]. Presses [WVHTK21]. Prestige [KSAB^{+}21]. Preventing [DCAT12]. Prevention [LLL22, SRK22]. Price [CKKK14, DABP14, HZB19, KAS14]. Priced [RML12]. Prices [CGL^{+}16]. Pricing [AHS14, CGL^{+}14, MMI23, XWML19, CWC14]. Pricing-based [MMI23]. Primitives [JDZ^{+}21]. Principled [FT02]. Principles [ABC^{+}17, PJZ18]. Privacy [ABCL17, ASÖY23, BHPY21, BCG^{+}18, BCCA^{+}21, CE21, CCD^{+}22, CIY^{+}21, CAN^{+}21, DTE17, FYZ19, KKY18, KK21, KS03, KYY17, LYW23, LLG22, LGGB^{+}21, LP21, MMK^{+}22, MGB^{+}21, MAK^{+}22, NZQX22, PLZW18, PSK10, PHC^{+}21, PDF^{+}23, SLBD20, SDB21, SWAHP21, STK17, TSM^{+}23, UY22, WZB^{+}21, WLW^{+}23, XZG^{+}22, YSZ^{+}22, YDZ^{+}21, ZGB18, ZLZ^{+}23, ZJQ^{+}21, Hoc02, Kri01, XCL07, MGB^{+}21]. Privacy-Aware [WLW^{+}23]. Privacy-Enhanced [DTE17]. Privacy-Preserving [ABCL17, MMK^{+}22, MAK^{+}22, PLZW18, PHC^{+}21, YSZ^{+}22, CE21, CCD^{+}22, FYZ19, PSK10, SLBD20, WZB^{+}21, XZG^{+}22, YDZ^{+}21]. PrivacyCheck [ZGB18]. Private [KAS14, ZXYL16]. Privileged [NZQX22]. Proactive [GCK^{+}22]. Probabilistic [CDJ^{+}22, KG10]. Probing [RMP10]. Problem [RML12, ZLS^{+}22]. Problems [CT17, SK17]. Process [ACDLM19, DRC^{+}23, GNR19, PPDG19, YBW19, GMM09]. Processes [ETRDRO^{+}19, SABG17, YBW19]. Processing [BGK14, LCS21, MS17, MP14, OKR^{+}14, PSA^{+}20, ZJL^{+}15, HP03]. Product [BWL16, HNGN23, NGER20, WLL^{+}13, WVHTK21]. profiles [AKS07, LLNF12]. profiling [USR09]. Profitability [YWML19]. Programmable [HHF^{+}21, HZCS10]. Programming [BBC14, GAL18, ZSL^{+}17]. Progressive [CSMM17, ZJL^{+}15]. project [BMS02]. PROLISEAN [HHF^{+}21]. Proof [KSAB^{+}21]. Proof-of-Prestige [KSAB^{+}21]. Properties [MMV11]. Property [EHY19]. Protect [TSM^{+}23]. Protecting [LYW23]. Protection [KK21, NZQX22, ZJQ^{+}21, YW10]. Protocol [HHF^{+}21, NT21, PCP^{+}20, SGC16, SL22, XSSD23, Hoc02]. Protocols [GAL18, SLG^{+}22]. PROV [Mor17, SABG17]. Provenance [BTGM22, BTH^{+}17, CCM17, GEFT14, NDO^{+}17, RIB18, SABG17, GMM09]. Provenance-Aware [RIB18]. Provide [FGS20]. providers [BSS02]. Providing [AJSS13, GS17, ZGD23, ZMGW22, LHTL06]. Provisioning [MA23, TEMH19, VPR07, VSID16, SPJ09]. proximity [PRD09]. Proxy [ATS^{+}21, BI17, PK20, RMMH22, YCM^{+}13, LHTL06]. Pruning [PWGQ22]. pseudonymity [KS03]. Pseudoperiodic [MSW^{+}16]. PSO [BBS21, JSAA22]. Public [LC16, TPQC22, DMT07]. Publish [DLZ^{+}16, PC22]. Publish/Subscribe [DLZ^{+}16, PC22]. Publishing [PLZW18, WRC01]. PUF [LXZ^{+}22]. Pump [LMSS23]. Purchase [PDF^{+}23]. Pure [EM19].

QoE [XIS22]. QoE-driven [XIS22]. QoS [GHD21, HAST21, JN08, SLG^{+}22, YXL^{+}21]. QoS-aware [HAST21, JN08, SLG^{+}22]. Quality [ASBH^{+}16, BKK03, CHC^{+}21, DOG^{+}22, GAL^{+}22, LSK^{+}17b, OWK^{+}19, PDS20, RDC16, SPKTG22, WVHTK21, WHM^{+}22, YCM^{+}13, ZXP^{+}22, Dal11]. Quality-Based [ASBH^{+}16]. Quality-of-Service [LSK^{+}17b]. Quantify [BCN17]. Quantifying [FLR23, STK17]. Quantitative [CGL^{+}16]. Quantized [SK24]. quantum [YCH^{+}22]. Queries [BJ15, CLM^{+}11, KA20, LC12, CTZZ06, GR04, LXW^{+}12]. Query [LMSTM14, ABMP07, PPV05]. query-conscious [ABMP07]. Querying [ZSY^{+}17, FFP09]. Question [LSLY19, VASD19, ZSL^{+}17]. questions [ALG04]. Quota [ABDL14]. QURSED [PPV05].

RA [PPDG19]. Radar [CYD^{+}20]. Radiomics [KGAR22]. Radiomics- [KGAR22]. Raising [DR05]. Random [CXG21, CSMM17, YMY^{+}23]. Range [CLM^{+}11]. ranking [BRRT05, LYF^{+}09]. ranks [THS06]. Rates [Glu10]. Rating [CO16, RIB18, FLD12]. ratings [JKR07]. Re [QLJ^{+}19, RMMH22]. Re-Encryption [RMMH22]. Re-identified [QLJ^{+}19]. Reachable [Nov19]. Reaching [HSRK23]. reading [LYF^{+}09]. Real [BJ15, MMI23, MPR^{+}23, TEMH19, WARCD17, WSM21, YLM^{+}23]. Real-Time [TEMH19, WARCD17, MMI23, MPR^{+}23, WSM21, YLM^{+}23]. Real-World [BJ15]. Reality [PDS20, PSL^{+}20, ZXP^{+}22]. Realtime [CPV^{+}16, JPCL22, ZGD23]. Reasonable [JG10]. Reasoning [EHY19, GL14, JPSS17, RPR22]. Receiver [CYD^{+}20]. Reciprocation [RSS17]. Reciprocity [YC18]. Recognition [AGKW14, CLM19, DCD^{+}21]. Recommendation [CJW^{+}23, CL24, CDC14, CO16, HXZ^{+}20, HJWW20, HMLH21, HZ11, LSLY19, PMN23, PHC^{+}21, WL23, YSNL16, YSW^{+}17, BGL04, OHKS04].

Recommendations [NPP^{+}15]. Recommender [AdM^{+}13, MBBW07, RS09]. Recommenders [JWW15]. Reconciliation [ASBH^{+}16]. Reconciling [LMZ13]. Reconfiguration [GHD21, SK17]. Reconstruction [ZXP^{+}22]. Recovery [BLSW04]. Recruitment [ASO^{+}22]. Recurrent [PWGQ22]. Recursive [VAS24]. reduced [Dal11]. Reduction [BTH^{+}17, CSMM17, KZLG21]. Redundancies [NZ22]. Reference [PPDG19, RHT20]. Regression [GZL^{+}21, Glu10, MKJB21]. Regular [GD17]. regulate [Ung05]. Regulation [AHS14]. Rehabilitation [KKK21]. Reinforcement [CLS^{+}22, HSLH17, KZLG21, LWH^{+}21, LOD19, RWXC20, SABL24, XLL20]. Reinforcement-Enhanced [HSLH17]. Reissue [GAC18]. Relation [LJLN16]. relational [YASU01]. Relations [YSNL16]. Relationship [BBH^{+}14, SGOS19]. Relationship-Based [BBH^{+}14]. Relationships [KAS14, SWD15, GH06]. Releasing [CAN^{+}21]. Relevance [FSC15]. Relevant [NYB^{+}19]. Reliable [MBS19, ZMGW22]. remailer [GM04]. Remote [ZXP^{+}22, KMW09, Zdu08]. Replica [SCPB22]. Replica- [SCPB22]. Replication [ZWC^{+}17]. Reporting [BTGM22]. Reports [JCH^{+}18]. repository [SS06]. Representation [HLG^{+}21]. Reputation [BTGM22, MMR16, DMGR^{+}17, MQB22, PAS13, RIB18, RCP^{+}15, SXZ^{+}21, XLL20]. Reputation-Based [PAS13, RCP^{+}15, BTGM22]. Requirements [KS07]. Research [SLPZ23]. Resident [SCLB24]. Resilience [BCN17]. Resilient [RPR22]. Resistant [LZK^{+}22]. Resolution [GZL^{+}21, KBNV18, LHTL06]. Resolutions [LZJ^{+}21]. Resolvers [SK13]. Resolving [KYY17]. Resource [AZKG21, ADAP19, BJ15, JSAA22, LWM^{+}21, LLSW22, MRS^{+}22b, MMI23, MA23, TK11, USR09, ZXS08, AOVP08, ZHDD07]. Resource-adaptive [LWM^{+}21]. Resources [AKOB^{+}21, BJ15, HAST21, ZB20]. RESP [VAS24]. Response [GAC18, LWH^{+}21, WZKP19, ZWW^{+}23]. Responsibility [KKY18]. restrictive [GM04]. result [LM04]. Rethinking [BC01]. Retraining [WGW^{+}24]. Retrieval [ZJL^{+}15, DKP12, MPC06, PSK10, Rin09, TGRBD07, YASU01]. Retrieving [FFP09]. Retweet [BLD^{+}15, YYM^{+}19]. Reusable [CDC14]. Revealed [SK13]. Revealing [SdMA^{+}14]. Revenue [CKKK14]. Reverse [DPD22]. Review [HJWW20, NGER20, PSA21, BF06]. Review-based [NGER20]. Reviewers [Sin17, Sin18]. Reviews [BWL16, BC23, HNGN23, LSK^{+}17b]. revisited [Bri06]. Revisiting [MCS18]. Reward [KSAB^{+}21]. RFID [LXZ^{+}22]. RFID-PUF [LXZ^{+}22]. Right [DABP14]. Rights [JS13]. Ring [CLM^{+}11]. Riot [ABR17]. Risk [BCCA^{+}21, CRP17, LJG18]. Risks [MCS18]. Risky [LHAT22]. RL [RWXC20]. RNS [MMJ21]. Robot [KKK21, LYW^{+}21, ZTL^{+}21]. Robotic [CCN^{+}21]. Robotics [CXH^{+}21, LWFD21, LQW21]. Robots [PHR^{+}21]. Robust [GZL^{+}21, RZJ20, HLLS21, WRC01]. robustness [MBBW07, OHKS04]. Role [FPR16, PDS20, SWD15, YYM^{+}19, DD07]. Rotating [CIY^{+}21]. Rotten [TBG^{+}18]. Route [LLG22, ZLZ^{+}23]. Routes [CSS20]. Routing [GNW^{+}20, SLG^{+}22, WQC^{+}19, ZLZ^{+}23, ZWC^{+}17, GNK11]. rSYBL [CMTD16]. RTChain [SXZ^{+}21]. Runtime [ATD22].

S [WCX^{+}23]. S-BDS [WCX^{+}23]. safe [Thi05]. Safety [CXW^{+}21, MJ22]. SafeVchat [LXC^{+}13]. Sale [YWML19].

SAM [ZWW^{+}23]. Sample [CYG^{+}21, WVHTK21]. Sampling [PWSG22]. Sanitization [WSLT21]. SANTM [TJGY22]. Sarcasm [ZMT^{+}23]. Satisfiability [ATD22]. saving [DWF24]. scalability [AKR01]. Scalable [MPR^{+}23, SCPB22, VSID16, KS07]. Scale [BDM10, DRC^{+}23, PK20, TSM21, VSID16, GNW^{+}20, JKS^{+}10, PT09]. Scams [CPL^{+}21]. sCARE [MMR16]. Scenarios [YLM^{+}23]. Scheduling [HAST21, KGKK21, LOD19, LMS^{+}21, PSP22, AM03, SHB06]. Schema [GLQ11, CS09, MLMK05]. Scheme [CIY^{+}21, CLJ^{+}21, CMML22, GSZ^{+}23, GNW^{+}20, KLS^{+}17, KA20, LLG22, LHL^{+}22, LSZ^{+}21, MRS^{+}22b, PCV^{+}21, PHC^{+}21, PO19, RMMH22, SL22, WCX^{+}23, XZG^{+}22, YSZ^{+}22, YZL^{+}24]. Science [PBJP21]. Scientific [NDO^{+}17]. Score [IDS19]. Screw [CHC^{+}21]. scripting [Thi05]. SDN [DWGC23, MA23, SK24]. SDN-enabled [MA23, SK24]. Seamless [FYT17]. Search [CDM^{+}14, Glu10, GWF^{+}21, JDZ^{+}21, MSG^{+}21, VAKK19, YZY^{+}14, YMY^{+}23, ZXYL16, CS07, JMSP06, LM04, LLNF12, MYS^{+}12, NDL07, XZZ08]. Searching [ACGM^{+}01, BF06]. Second [CKKK14]. Secondary [HKV14]. Section [BHPY21, DNJ19, FLLM22, GDLM22, HXB^{+}22, HAD22, LWFD21, MQUXK22, NBFZ15, PBJP21, SWAHP21, SLPZ23, WRWM21, XZJO22, ZBF^{+}19]. Secure [ATS^{+}21, BCGN16, BAM^{+}22, CCD^{+}22, CGS23, DLZ^{+}16, FMC19, GWF^{+}21, KSL^{+}21, LJS^{+}14, LDG^{+}23, MRS^{+}22a, Nov19, SKH22, WNN^{+}22, YLZ^{+}21, CPV03, GNK11, SBG07]. Secured [UNBAT22]. Securing [AKA^{+}23, MPR^{+}23]. Security [AKOB^{+}21, AAA^{+}20, BHPY21, BBS21, BCG^{+}18, CTS^{+}24, CRP17, GAC18, GBAR08, HJ08, HHF^{+}21, HAD22, ISG^{+}22, IRJ^{+}21, JLC20, JDZ^{+}21, LYW23, LXZ^{+}22, LQSW21, LP21, LLL22, MQUXK22, QZDG22, SST^{+}16, SWAHP21, STJ^{+}21, WG23, YCH^{+}22, ZKC^{+}22, ZLS^{+}22, BDT04, CPV03, KS07]. Security-Problem-Based [ZLS^{+}22]. See [SdMA^{+}14]. Segmentation [HML^{+}21]. segmented [LM04]. Selecting [JWW15]. Selection [DOG^{+}22, LPX^{+}21, MBS19, STB^{+}19, ZWC^{+}22, ZWW^{+}23, FS04]. Selective [DK04]. Self [DXP^{+}23, DKM^{+}02, RZJ20, SBC20, SS20, TJGY22, HBGF02]. Self-Adaptation [SS20]. Self-adaptive [RZJ20]. self-administering [HBGF02]. Self-attention-driven [TJGY22]. self-configuring [HBGF02]. Self-Healing [SBC20]. Self-similarity [DKM^{+}02]. Self-supervised [DXP^{+}23]. sellers [Guo02]. Semantic [HC14, JKS^{+}10, LJLN16, LYW^{+}21, RAR22, YBMV22, ZWW^{+}23, BCF^{+}07, GR04, JAT^{+}06, MBB07, MGB^{+}07, Rin09, SNBC12, TGRBD07, OSSV05]. semantically [AKS07]. Semantics [BCP08, DRC^{+}23, VJL^{+}14]. Semantics-based [BCP08]. Semi [HXZ^{+}20, JHC^{+}22]. Semi-Direct [JHC^{+}22]. Semi-supervised [HXZ^{+}20]. SemIoTic [YBMV22]. semistructured [PPV05]. Sensemaking [LSK^{+}17a]. Sensing [CPV^{+}16, LHZ^{+}21, PK20, PMFS17, RZP^{+}22, NZ22, PCBG19]. Sensing-as-a-Service [LHZ^{+}21]. sensitive [PSP22, SNBC12]. Sensor [CYD^{+}20, CYWW22, PK20, RQL^{+}21, SS20, SPKTG22, WLW^{+}23, WVHTK21, YLC^{+}22, MYS^{+}12]. Sensors [BI17, LZBN17, PSL^{+}20]. Sentence [LYW^{+}21]. Sentiment [HZB19, HJWW20, MSK17, YV22]. separation [JKR07]. separations [GS07a]. Sequence [CJW^{+}23]. Sequences [CSS20, KGAR22]. sequencing [KRML09]. Sequential [RML12]. Sequentially [CAN^{+}21]. Serendipity [GCP^{+}20]. Serendipity-based [GCP^{+}20]. Series [ZTL^{+}21, YDZ^{+}21]. Server [BCO13, TK11, TSM^{+}23, VAS24, KLMH03, LHTL06, Thi05]. server-directed [KLMH03]. Server-Side [BCO13, Thi05]. Serverless [WYC^{+}23]. Servers [XZG^{+}22, LB04, SHB06, VPR07]. Service [AO22, AHM^{+}15, AV16, BBH^{+}14, BCGN16, CLF^{+}19, DOG^{+}22, DJ15, FYW^{+}22, GHD21, HHS^{+}22, KKMK16, LHZ^{+}21, LSK^{+}17b, MBS19, OWK^{+}19, PGT^{+}18, PHC^{+}21, SPKTG22, TSS19, TK11, UNBAT22, WCC20, XWML19, YBZ14, YWML19, YXL^{+}21, ZMGW22, BCF^{+}07, BKK03, CFTV03, HZHC12, JN08, LZW^{+}22, MBC^{+}05, NCEF02, PRD09, SPJ09, TGRBD07, Van08, Zdu08, vdADO^{+}08, ML08, YCM^{+}13]. Service-Based [AHM^{+}15]. service-enabled [MBC^{+}05]. service-oriented [Van08, Zdu08, ML08]. Services [ALA^{+}19, BB23, CWC14, CZPS22, CMTD16, DOG^{+}22, DLZ^{+}16, GdOW14, JPCL22, KFB^{+}14, LMZ13, LXC^{+}13, LGKL20, MMR16, MQUXK22, NBM19, RWXC20, SSC23, TEMH19, Web17, XIS22, ZGD23, AR12, AJP07, BCMS06, BCP^{+}04, BCP08, DD07, FLD12, LHTL06, MBB07, MGB^{+}07, PP11, SBG07, SD12, SNBC12, XCL07, ZHDD07]. Serving [FYW^{+}22]. SESAME [YZY^{+}14]. session [DCAT12]. Set [SO17, WDK^{+}24]. Set-Generalized [SO17]. sets [Dal11]. Shard [XSSD23]. Sharding [XSSD23]. SHARE [JPSS17]. Shared [AO22, WSLT21, USR09]. Sharing [AO22, BCFB18, GSZ^{+}23, LHZ^{+}21, SCW17, ZHDD07]. shopping [AKR01]. Short [BLTH22, CWW^{+}21, CLW^{+}22, DCD^{+}21, SCW17]. Short-Term [BLTH22, CLW^{+}22, DCD^{+}21]. Short-Video [SCW17]. Should [GAC18]. Show [OALA17]. Siamese [NLLC21]. Side [BCO13, MMJ21, Thi05]. Side-Channel [MMJ21]. Sign [SPM^{+}13]. Sign-On [SPM^{+}13]. Signal [KZLG21, RZP^{+}22]. Signature [Mor17]. Signatures [YCH^{+}22, DMT07]. Signed [CO16, YMY^{+}23]. similar [CDM10]. Similarity [HSLH17, XM17, DKM^{+}02, PSK10]. similarity-based [PSK10]. Simulation [SF21]. Simulation-driven [SF21]. simulations [JKS^{+}10]. Simulator [PSP22]. Single [SPM^{+}13, Gel09, MS05]. single-instance [MS05]. Site [BDM10, EV07, WL07, ZHH04]. site-dependent [WL07]. site-invariant [WL07]. Sites [BWL16, MAM03, ZH09]. Situated [GHK17]. SkillBot [LHAT22]. Skills [LHAT22]. Sky [HSRK23]. Skyline [WTS^{+}21]. SLA [KGKK21]. SLA-driven [KGKK21]. Slot [CHC^{+}21]. Small [WCY^{+}23, YC18]. Small-World [YC18]. Smart [AZKG21, ABCL17, BCGN16, CCD^{+}22, CGG^{+}22, CXG21, CLM19, DKP17, DLZ^{+}16, GDLM22, HML^{+}21, KLS^{+}17, KK21, KZLG21, LHZ^{+}21, LPR19, LQSW21, MED19, PGP^{+}21, RTR^{+}22, SK24, SPE^{+}22, SH22, PBL^{+}22, SWAHP21, SPCC23, SCLB24, TSY^{+}21, VBD^{+}22, WRWM21, YBMV22, ZTH^{+}23, DMT07, HZHC12, NCEF02, PMFS17, WLW^{+}23]. Smartphone [PRKD20, WWZ^{+}23]. Smartphone-based [PRKD20]. SMig [RWXC20]. SMig-RL [RWXC20]. snippets [XZZ08]. SNR [HMLH21]. SOAs [KIG^{+}19]. Social [ALA^{+}19, BCFB18, BGK14, BPSD17, BKS^{+}14, CCC^{+}23, CAV14, CSS17, CDPR17, CO16, FYZ19, FYZ^{+}21, FAGB14, GRR20, GLWH17, GLT17, HLG^{+}21, HMLH21, JWW15, KKY18, KYY17, KBBI15, LBC^{+}24, MBE22, MS17, NBFZ15, PSL^{+}20, QLJ^{+}19, RCP^{+}15, RZAD17, SCW17, SZT22, SGOS19, SWD15, VJL^{+}14, WARCD17, WJL^{+}22, VAK17, YPFY21, YZY^{+}14, YLL^{+}17, ZLL^{+}20, FLD12, GH06, Hoc02, KG10]. Social-aware [HMLH21]. Social-Chain [YPFY21]. Socio [BBC14]. Socio-Technical [BBC14]. Software [BG21, DKP12, GK23, LWM^{+}21, PJZ18, SCL^{+}19, WQC^{+}19, XvHWW18, YLZ^{+}21, BVT06]. Software-defined [YLZ^{+}21]. Soil [RZP^{+}22]. SoIoT [KKMK16]. Solution [WG23]. Solutions [BSBP16, NZ22, CPV03]. Solve [LLL22, RML12]. Solving [SK17]. Source [NYB^{+}19, ZGF^{+}23]. Source-Aware [NYB^{+}19]. Sourced [LZBN17]. Sources [ADGM23, FSC15, WLL^{+}13, ZHL^{+}16, FFP09]. Sourcing [ASO^{+}22, AJP07]. SouthamptonTAC [HJ03]. space [ZXS08]. Spaces [YBMV22]. spam [GM04, WSM21]. Spanish [PDAMGULMV20]. Sparse [HXZ^{+}20, PWSG22]. Sparsity [HSLH17]. Spatial [AAF18, HLLS21, GS07a]. Spatial-temporal [HLLS21]. Spatially [TGBG20]. Spatio [AZKG21]. Spatio-temporal [AZKG21]. Speaking [MHCF22]. Special [BHPY21, BBP18, BCG^{+}18, CGT^{+}21, CAV14, CSS17, CZPS22, CGL^{+}14, DNJ19, FLLM22, GDLM22, GNR19, HXB^{+}22, HAD22, KBBI15, LPR19, LWFD21, MQUXK22, MFR^{+}21, PBJP21, SSC23, SWAHP21, SSKW20, SLPZ23, TSS19, WRWM21, XZJO22, XvHWW18, ZBF^{+}19, LLSM08, MBB07, SD12]. Specific [LSK^{+}17b, GLFV^{+}21, Thi05]. Specifying [CMTD16]. Spectrum [DXP^{+}23, HKV14]. Speculation [OGP^{+}18]. Speculative [MP14]. Speech [PSA^{+}20]. Spinel [BI17]. Sponsored [Glu10]. spontaneous [RS09]. spoofing [EL09, HJ08]. Spread [GJAT^{+}21]. Spy [NDL07]. Squares [TSM21]. SSL [HXZ^{+}20, PP11]. SSL-SVD [HXZ^{+}20]. SSL/TLS [PP11]. SSL/TLS-based [PP11]. Stable [WWZ^{+}23]. Stack [RMMH22]. Stackelberg [JPSS17, LZW^{+}22]. Stackelberg-game [LZW^{+}22]. Stage [LHL^{+}22]. Stance [MSK17, ZMT^{+}23]. Stance-centered [ZMT^{+}23]. Stance-level [ZMT^{+}23]. Standards [Kri01]. Stanford [CGMH^{+}06]. State [KZLG21, LT16, NT21, EL09, KMW09]. stateless [DCAT12]. statically [HP03]. Station [TF21]. Statistical [LSK^{+}17b, WLB22]. Status [PCP^{+}20]. stochastic [FLL06]. Stock [HZB19]. Storage [LYW23, Liu20, TPQC22, WCX^{+}23, YASU01]. stored [LCKN05]. Strategic [DGWW15, PHR^{+}21]. Strategies [BCFB18, YCM^{+}13]. Strategy [YWML19, ZB20, Guo02, HJPB06]. Stream [GEFT14]. Streaming [CCD^{+}22, MA23, Dal11, LCKN05, TJLC08]. Streams [MSW^{+}16, MP14]. Street [LMSS23]. Strength [RZP^{+}22]. Strong [XSSD23]. Structural [ZGF^{+}23]. Structure [LPB^{+}17, YLL^{+}17, Coo03]. Structured [CXG21, EM19, GHK17, HCW^{+}21]. structures [GLJ^{+}12]. Study [FAGB14, HCW^{+}21, LC16, OKM21, RDC16, DD07]. Style [OALA17]. subjectively [Coo03]. Subscribe [DLZ^{+}16, PC22]. Subsidization [Web17]. Summarization [NYB^{+}19, PC22, ZGB18]. Summary [CWW^{+}21]. Super [GZL^{+}21]. Super-Resolution [GZL^{+}21]. Supervised [CLJ^{+}21, MSW^{+}16, DXP^{+}23, HXZ^{+}20]. Supply [SCZ^{+}21, XZY^{+}21]. Supply-chain [SCZ^{+}21]. Support [APAC18, DRC^{+}23, JSAA22, ZGD23, SMFR08]. Supporting [CTZZ06, CS07, OSSV05, TMK^{+}12, UY22, ZHDD07]. Supportive [KBNV18]. supports [LLSL08]. Surface [WCY^{+}23]. Survey [PML^{+}19, PBL^{+}22, CPV03]. Survival [MGHB16, YCM^{+}13]. Sustainability [LFL17]. Sustainable [IRJ^{+}21]. SVD [HXZ^{+}20]. SVM [NZQX22]. SVMs [TSM21]. Swarm [JDZ^{+}21, SZT22]. Swarm-like [JDZ^{+}21]. Switches [YLZ^{+}21]. syndication [DR05]. Synergic [SPE^{+}22]. System [ALS23, AdM^{+}13, Ano15, CGG^{+}22, CHC^{+}21, GDLM22, HAST21, JKI^{+}21, JGH^{+}22, KKK21, KSAB^{+}21, LXC^{+}13, LHZ^{+}21, LS21, MED19, OKR^{+}14, PMN23, PC22, PHC^{+}21, RPA^{+}17, RIB18, DFL^{+}23, SCZ^{+}21, SPG22, SXZ^{+}21, WWZ^{+}23, WSLT21, XIS22, YLM^{+}23, ZTL^{+}21, AKR01, HBGF02, KRRT06, LYF^{+}09, PPV05, RS09].

System-based [WSLT21]. Systematic [LJG18, PSA21]. Systems [AKOB^{+}21, AHM^{+}15, ATD22, BBC14, CWLZ19, CZPS22, CDPR17, CGL^{+}14, CLM^{+}11, DSVA19, DLZ^{+}16, FFKG19, FYZ19, FYZ^{+}21, FLLM22, FPA^{+}23, GAT^{+}21, ISG^{+}22, KGKK21, LJG18, LWM^{+}21, LFL17, LSZ^{+}21, MQUXK22, NLLC21, NBFZ15, PDS20, PPDG19, RIB18, TGBG20, WCY^{+}23, XvHWW18, XLL20, YXL^{+}21, ZOC11, ZZF^{+}23, AGPS05, AJP07, BF06, CS09, KS03, LB04, MBBW07, VPR07, WRC01, CGT^{+}21, PBJP21]. Systolic [YCH^{+}22].

tactic [MS05]. Tactile [CCN^{+}21, CHC^{+}21, YLZ^{+}21]. Tag [LSLY19]. Tagging [BGK14]. Tail [WZKP19]. Taiwanese [LLC^{+}23]. Taming [BTH^{+}17, BTC^{+}23]. Tangible [MGB^{+}21]. Target [GCK^{+}22]. Task [GAL^{+}22, HCW^{+}21, HAST21, JGH^{+}22, MMK^{+}22, YZL^{+}24, ZWC^{+}22, ZDCB18]. Tasks [KSAB^{+}21]. Taxonomy [ADA^{+}22, MLMK05, LXW^{+}12]. taxonomy-oriented [LXW^{+}12]. TBchain [LYW23]. TCPS [PSP22]. Teaching [KSA^{+}10]. Team [LJS^{+}14]. Teaming [CTS^{+}24]. Teamwork [HS19]. Technical [BBC14]. Technique [STJ^{+}21]. Techniques [AGKW14, OKM21, AM07]. Technologies [BCN17, DNJ19, PDAMGULMV20, Web17, WYC^{+}23, XvHWW18, LLSM08]. Technology [KBNV18, LSK^{+}17a, LLSL08, Liu20, LP21, SCZ^{+}21, GS07a, GBAR08]. telecommunication [BCP^{+}04]. Television [DTE17]. Temperature [WLB22]. temporal [AZKG21, GS07a, HLLS21]. Tenancy [HSRK23]. Tensor [FYZ^{+}21]. Tensor-based [FYZ^{+}21]. Term [BLTH22, CLW^{+}22, DCD^{+}21]. Test [JCH^{+}18]. Testbed [SST^{+}16]. Tethering [PRKD20]. Text [TJGY22, WMW^{+}22, PSK10]. Text-based [WMW^{+}22]. Texts [CWW^{+}21]. Textual [BC23]. Textual-based [BC23]. their [SK13]. Theme [NBFZ15]. Theoretic [ADAP19, PHR^{+}21, YC18]. Theory [GLJ^{+}12, RZAD17, YJL^{+}22, BRRT05, MLMK05]. Theory-Based [RZAD17]. There [ZW17]. Things [BCGN16, Nov19, YSNL16, HZHC12, ADA^{+}22, BHPY21, BI17, BTC^{+}23, CZPS22, HC14, IRJ^{+}21, LNTL23, LZK^{+}22, LS21, LLSW22, MGHB16, MRS^{+}22b, PC22, PML^{+}19, RMMH22, SSA^{+}21, SCZ^{+}21, TSY^{+}21, TSS19, TGBG20, WRWM21, YZL^{+}24, YSNL16, ZTL^{+}21, GCK^{+}22, MFR^{+}21]. Third [BZVS18, MHCF22, XJ20]. Third-Party [MHCF22, BZVS18, XJ20]. Thistle [CBM23]. Threat [AE24, FFKG19]. Threats [LJS^{+}14]. Three [LYW23]. Three-tier [LYW23]. Threshold [WSLT21]. throttling [RTcR19]. Throughput [DWGC23, HSRK23, RZJ20]. Thwart [LJS^{+}14]. Ticket [ATS^{+}21]. Ticket-Based [ATS^{+}21]. Tier [DJ15, RMMH22, WZKP19, LYW23, VPR07]. Time [CYG^{+}21, PSZ24, PSP22, TEMH19, WARCD17, WZKP19, YDZ^{+}21, ZTL^{+}21, DCAT12, MMI23, MS05, MPR^{+}23, WSM21, YLM^{+}23]. time-dependent [MS05]. Time-Efficient [CYG^{+}21]. Time-interval [PSZ24]. Time-sensitive [PSP22]. Time-series [YDZ^{+}21]. Tip [HNGN23]. Tips [HNGN23]. TLS-based [PP11]. TM [MBS19]. TOIT [Sin13a, Sin13b, Sin17, Sin18]. Token [MRS^{+}22a]. Token-Based [MRS^{+}22a]. tokens [DCAT12]. Tolerance [FPA^{+}23, XZY^{+}21]. Tolerant [WEJ14, XFL^{+}23]. Top [BGK14, HZ11]. Top- [BGK14, HZ11]. Topic [SR13, VJL^{+}14, LYF^{+}09]. Topical [MPS04]. Topics [WMW^{+}22]. Topologies [WK18]. Tor [DFLT22]. Tourist [WCZ^{+}21]. Tracker [BZVS18]. Tracking [APAC18]. trade [AOVP08, LB04]. trade-offs [AOVP08]. Tradeoff [YC18]. Tradeoffs [TGBG20, XLL20]. Trading [WMG^{+}21, HJ03]. Traffic [CLW^{+}22, GVM^{+}23, JG10, KZLG21, MMV11, SK24, WARCD17, WNN^{+}22, WLW^{+}23, XCRY22]. Trained [MHA^{+}21]. Training [CGS23]. Trait [OALA17]. Trajectory [XCRY22]. Transaction [CPL^{+}21, LBC^{+}24, SXZ^{+}21, TNJJ22]. Transactions [MFR^{+}21, PAS13, SO17, CPV03, Ung05]. transcoding [KLMH03]. Transfer [DZHV16, LLSW22]. Transform [PWSG22]. transformations [AR12]. Transit [ASW^{+}22]. translator [HZCS10]. Transmission [SPAT21]. Transmit [PACH20]. Transmitting [SATPR22]. Transparency [GAC18]. Transparent [XJ20, YW10]. Transportation [CGG^{+}22, GDLM22, RTR^{+}22]. Traversal [Nov19]. tree [CMML22, GLJ^{+}12, LSCZ05]. Trend [JGH^{+}22]. Trending [WMW^{+}22]. Trends [LT16, SRK22]. Tripartite [SATPR22]. Trust [BB23, BHPY21, De 19, GSZ^{+}23, HS19, HXZ^{+}20, HZB19, IDS19, JPCL22, JWW15, LNTL23, DMGR^{+}17, MBS19, NBFZ15, PHC^{+}21, PAS13, RSS17, RZAD17, DFL^{+}23, SWD15, WCX^{+}23, WLW^{+}23, YPFY21, YZL^{+}24, YCC17, ZBF^{+}19, GH06, KG10]. Trusting [FSC15]. Trustworthy [BTH^{+}17, PMFS17, XJ20, MBBW07]. TSCH [CSS20]. TSK [JGH^{+}22]. Tumor [HML^{+}21, KGAR22]. Tuneman [SKA^{+}23]. turn [ZWW^{+}23]. Tweet [NYB^{+}19]. Tweets [MS17, MSK17, PDAMGULMV20]. twig [KRML09]. Twin [TSM21, WCY^{+}23]. Twins [RCM^{+}22]. Twitter [ABR17, BLD^{+}15, FPR16, HZB19, VJL^{+}14]. Two [AO22, LHL^{+}22, PDF^{+}23]. Two-Stage [LHL^{+}22]. Two-way [AO22, PDF^{+}23]. type [BC23, Thi05]. type-safe [Thi05]. typed [HP03].

U.S. [Hoc02]. UAV [LHL^{+}22].

UAV-Assisted [LHL^{+}22]. UAVs [FGS20]. Ubiquitous [YBW19, MYS^{+}12]. UK [CB15]. Ultra [GAL^{+}22, MEAK^{+}21]. Ultra-low [MEAK^{+}21]. Ultrasound [MHA^{+}21]. Uncertain [BSBP16, MSW^{+}16, MMR16]. Uncertainty [ASÖY23, GAC18, YXP^{+}18]. Uncertainty-Aware [ASÖY23]. Understanding [ABDL14, CLZ^{+}20, HS19, MHCF22, XIS22, PVL^{+}17]. Underwater [YLC^{+}22]. UNET [HML^{+}21]. Unexplained [MMP^{+}14]. Unified [ADGM23, BMG^{+}19]. UNION [XFL^{+}23]. Universal [ALA^{+}19, WS17]. Unlearning [WGW^{+}24]. Unpaired [DXP^{+}23]. Unreasonable [JG10]. Unstructured [MAK^{+}22, SABG17]. Unsupervised [BWL16, CWW^{+}21]. untrusted [CPV03]. Unverifiable [KSAB^{+}21]. Update [SCL^{+}19]. Updates [Sin13a, SL22, KMW09]. updating [MPC06]. Upgrades [LDG^{+}23]. upon [DJ15]. Upper [KKK21]. ups [GMM09]. Urban [HLLS21, LZBN17, LGKL20, PMFS17, SH22]. Ursa [RZJ20]. Usage [SH22, TK11, Coo03]. Use [ASW^{+}22, GPM^{+}18, Coo03]. Useful [KSAB^{+}21]. User [AO22, AJSS13, ADA^{+}22, BLD^{+}15, CAN^{+}21, Dal11, HZB19, HJWW20, JS13, KBNV18, KKMK16, LSK^{+}17b, MHCF22, PDS20, SDB21, SPM^{+}13, TSM^{+}23, WHM^{+}22, YZY^{+}14, YCC17, ZTH^{+}23, AKS07, CCD^{+}22, KS03, LLNF12, MAB19, SNBC12, NDL07]. user-adaptive [KS03]. User-Agent [YCC17]. User-centered [SDB21]. User-Centric [CAN^{+}21, KKMK16]. User-perceived [Dal11]. Users [DJ15, DPCM16, QLJ^{+}19, UY22]. Using [AAJ21, ABR17, CT17, CLL23, CLM19, CLM^{+}11, CXW^{+}21, DCD^{+}21, HCBRM23, HSLH17, HZB19, IRJ^{+}21, JHC^{+}22, KGAR22, KG10, LPB^{+}17, LGGB^{+}21, MKJB21, MGHB16, MRS^{+}22b, MHA^{+}21, MMJ21, MBE22, DMGR^{+}17, NT21, NZQX22, PRKD20, PDAMGULMV20, RTR^{+}22, RML12, SABL24, SZT22, SCZ^{+}21, Ung05, WVHTK21, YDZ^{+}21, ZGB18, ZOC11, ZJQ^{+}21, ZH09, Dal11, GR04, JKR07, JGH^{+}22, JSAA22, MS05, MLMK05, NDL07, PRD09, SGOS19, TNJJ22, UNBAT22, XCRY22, XZZ08, YASU01, GS07a]. Utility [GLFV^{+}21, PLZW18, SAB^{+}18]. Utility-Based [SAB^{+}18]. Utility-Controlled [PLZW18]. UTS [BCN17].

V [MRY^{+}23]. V-Gas [MRY^{+}23]. Vaccine [CXW^{+}21]. vague [FFP09]. Validation [Mor17, SLBD20]. Values [KBNV18]. VANETs [LLG22, YSZ^{+}22]. variability [DR05]. Variable [Glu10]. Variation [LC16]. Varied [GLFV^{+}21]. Varying [HHS^{+}22]. Vector [JSAA22, AE24]. vehicle [ZMGW22]. Vehicles [ASW^{+}22, CLW^{+}22, HAD22, MRS^{+}22a, MJ22, TPQC22, WNN^{+}22, YCH^{+}22]. Vehicular [JPCL22]. Verifiability [RHT20]. Verification [LDG^{+}23, MJ22, RQL^{+}21, YJL^{+}22, YXP^{+}18, AR12]. Vertical [WGW^{+}24]. via [AKOB^{+}21, CH05, CLZ^{+}20, De 19, EDC20, GEFT14, GJAT^{+}21, JDZ^{+}21, KBBI15, LWH^{+}21, LZBN17, PV17, RMP10, WMWM20, WGW^{+}24, Web17, WL07, YV22, YMY^{+}23, YBW19]. Viability [CWC14]. Video [LXC^{+}13, SCW17]. View [DvRDHB22, YJL^{+}22, YCM^{+}13, ZJQ^{+}21]. Views [LC12, GR04]. Virtual [CCN^{+}21, FYT17, MBP^{+}17, ZXP^{+}22]. Virtualization [BLMP22]. Vision [Sin13b]. Visual [EM19, JHC^{+}22, XM17, CMTT24]. Visual-Inertial [JHC^{+}22]. Visualization [PSA^{+}20, WLL^{+}13, ATB^{+}11]. visually [CDM10]. Voice [VBD^{+}22]. VoiceTalk [LLC^{+}23]. Volatile [ATD22]. Volunteer [AAJ21, ATS^{+}21, BAM^{+}22, HAST21, LMS^{+}21, LCS21, WTS^{+}21]. Volunteered [SPAT21]. VORTEX [CMTT24]. voting [NDL07]. vs [BC01]. Vulnerabilities [FLD12, JLC20]. Vulnerability [MRY^{+}23].

Waiting [CCN^{+}21]. Walk [YMY^{+}23]. Wall [LMSS23]. wars [GM04]. Watermarking [STJ^{+}21]. way [AO22, PDF^{+}23]. WBANs [CLS^{+}22]. Weak [ZOC11]. Wearable [CE21, CXH^{+}21, SST^{+}16, ZKC^{+}22]. Weaving [CDC14]. web [AKR01, Coo03, DKM^{+}02, EV03, LLNF12, MPS04, MAM03, Wil02, WY01, YADI02, AHM14, APAC18, ALG04, AKS07, AM07, ACGM^{+}01, AGPS05, ADGM23, BYCE07, BDM10, BF06, BBH^{+}14, BSBP16, BWL16, BZVS18, BCF^{+}07, BCP08, CDMF07, CDIW05, CDM10, CS07, DOG^{+}22, DCL^{+}22, DK04, DvRDHB22, DLLM07, EV03, EV07, EM19, FS04, FLL06, FT02, FLR23, GPM^{+}18, GR04, GH06, GS05, GLF02, HNF^{+}05, HZHC12, JMSP06, KMB^{+}22, KFB^{+}14, KG10, LM04, LJLN16, LCKN05, LSCZ05, LMZ13, LHTL06, LYM^{+}18, MYS^{+}12, MMR16, MBC^{+}05, MBB07, MGB^{+}07, OSSV05, OWK^{+}19, PRD09, RCM^{+}22, Rin09, RHT20, SHB06, SBG07, SS11, SD12, SPJ09, SPM^{+}13, SS06, Thi05, Van08, WLL^{+}13, WL07, XM17, XZZ08, ZSY^{+}17, ZHDD07, ZH09, ZHH04]. web-based [AKR01, SS11, AGPS05, GH06, KFB^{+}14, KG10]. Web-enabled [SS06]. WebBase [CGMH^{+}06]. Webchain [RHT20]. Webpage [JYW^{+}19]. Weed [CBM23]. Weighted [JGH^{+}22]. Weights [PWGQ22]. Wheeled [PHR^{+}21]. White [PMN23]. Who [MHCF22]. Wide [GLF02, RHT20, AOVP08, BVT06]. wide-area [AOVP08, BVT06]. WiFi [PRKD20]. Wireless [ABDL14, CYWW22, DPCM16, SS20, SPKTG22, ZJL^{+}15, Var03]. within [GD17, Hoc02, KMW09]. Word2Vec [QZDG22]. Work [KSAB^{+}21]. Workflow [GHD21, RTcR19]. workflows [SPJ09]. Workload [BCO13, CLFX24, FXYX23, MDDB19, XSW^{+}22]. workplaces [GBAR08]. World [BJ15, YV22, YC18, BC01, BRK04, GLF02, RHT20].

XDuce [HP03]. Xenophobia [PDAMGULMV20]. xlinkit [NCEF02]. XML [ABMP07, CTZZ06, CS09, FFP09, GLQ11, GL14, HP03, KRML09, LC12, LYW^{+}05, MPC06, MLMK05, YASU01]. XML-Path [GL14]. XQueC [ABMP07]. XRel [YASU01].

yellow [LXW^{+}12]. YouTube [FAGB14, SCW17].

Zero [GSZ^{+}23, LNTL23, WCX^{+}23, WLW^{+}23]. Zero-Trust [WCX^{+}23].

References

Anagnostopoulos:2020:LCS

[AAA^{+}20] Nikolaos Athanasios Anagnostopoulos, Saad Ahmad, Tolga Arul, Daniel Steinmetzer, Matthias Hollick, and Stefan Katzenbeisser. Low-cost security for next-generation IoT networks. ACM Transactions on Internet Technology (TOIT), 20(3):30:1–30:31, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3406280.

Angiulli:2018:ECS

[AAF18] Fabrizio Angiulli, Luciano Argento, and Angelo Furfaro. Exploiting content spatial distribution to improve detection of intrusions. ACM Transactions on Internet Technology (TOIT), 18(2):25:1–25:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

AlRidhawi:2021:IBM

[AAJ21] Ismaeel Al Ridhawi, Moayad Aloqaily, and Yaser Jararweh. An incentive-based mechanism for volunteer computing using blockchain. ACM Transactions on Internet Technology (TOIT), 21(4):87:1–87:22, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3419104.

Awad:2017:EAA

[ABC^{+}17] Edmond Awad, Jean-François Bonnefon, Martin Caminada, Thomas W. Malone, and Iyad Rahwan. Experimental assessment of aggregation principles in argumentationenabled collective intelligence. ACM Transactions on Internet Technology (TOIT), 17(3): 29:1–29:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ambrosin:2017:OBB

[ABCL17] Moreno Ambrosin, Paolo Braca, Mauro Conti, and Riccardo Lazzeretti. ODIN: Obfuscation-based privacy-preserving consensus algorithm for decentralized information fusion in smart device networks. ACM Transactions on Internet Technology (TOIT), 18(1): 6:1–6:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Andrews:2014:UQD

[ABDL14] Matthew Andrews, Glenn Bruns, Mustafa Dogru, and Hyoseop Lee. Understanding quota dynamics in wireless networks. ACM Transactions on Internet Technology (TOIT), 14(2– 3):14:1–14:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Arion:2007:XQC

[ABMP07] Andrei Arion, Angela Bonifati, Ioana Manolescu, and Andrea Pugliese. XQueC: a query-conscious compressed XML database. ACM Transactions on Internet Technology (TOIT), 7(2):10:1–10:??, May 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Abadi:2005:MHM

[ABMW05] Martin Abadi, Mike Burrows, Mark Manasse, and Ted Wobber. Moderately hard, memorybound functions. ACM Transactions on Internet Technology (TOIT), 5(2): 299–327, May 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Alsaedi:2017:CWP

[ABR17] Nasser Alsaedi, Pete Burnap, and Omer Rana. Can we predict a riot? Disruptive event detection using Twitter. ACM Transactions on Internet Technology (TOIT), 17 (2):18:1–18:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Armas-Cervantes:2019:LCD

[ACDLM19] Abel Armas-Cervantes, Marlon Dumas, Marcello La Rosa, and Abderrahmane Maaradji. Local concurrency detection in business process event logs. ACM Transactions on Internet Technology (TOIT), 19(1): 16:1–16:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Arlitt:2011:CIG

[ACG^{+}11] Martin Arlitt, Niklas Carlsson, Phillipa Gill, Aniket Mahanti, and Carey Williamson. Characterizing intelligence gathering and control on an edge network. ACM Transactions on Internet Technology (TOIT), 11 (1):2:1–2:??, July 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Arasu:2001:SW

[ACGM^{+}01] Arvind Arasu, Junghoo Cho, Hector Garcia-Molina, Andreas Paepcke, and Sriram Raghavan. Searching the Web. ACM Transactions on Internet Technology (TOIT), 1(1): 2–43, August 2001. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Awan:2022:TMB

[ADA^{+}22] Kamran Ahmad Awan, Ikram Ud Din, Abeer Almogren, Neeraj Kumar, and Ahmad Almogren. A taxonomy of multimedia-based graphical user authentication for green Internet of Things. ACM Transactions on Internet Technology (TOIT), 22(2): 37:1–37:28, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3433544.

Avgeris:2019:ARA

[ADAP19] Marios Avgeris, Dimitrios Dechouniotis, Nikolaos Athanasopoulos, and Symeon Papavassiliou. Adaptive resource allocation for computation offloading: a controltheoretic approach. ACM Transactions on Internet Technology (TOIT), 19(2):23:1–23:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

                                        URL https://dl.acm.
                                        org/ft_gateway.cfm?
                                        id=3284553.

Asprino:2023:KGC

[ADGM23] Luigi Asprino, Enrico Daga, Aldo Gangemi, and Paul Mulholland. Knowledge graph construction with a façade: a unified method to access heterogeneous data sources on the Web. ACM Transactions on Internet Technology (TOIT), 23(1):6:1–6:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3555312.

Albanese:2013:MRS

[AdM^{+}13] Massimiliano Albanese, Antonio d’Acierno, Vincenzo Moscato, Fabio Persia, and Antonio Picariello. A multimedia recommender system. ACM Transactions on Internet Technology (TOIT), 13(1):3:1–3:??, November 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Aguru:2024:OIB

[AE24] Aswani Aguru and Suresh Erukala. OTI-IoT: a blockchain-based operational threat intelligence framework for multi-vector DDoS attacks. ACM Transactions on Internet Technology (TOIT), 24(3): 15:1–15:??, 2024. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3664287.

Artikis:2014:ERC

[AGKW14] Alexander Artikis, Avigdor Gal, Vana Kalogeraki, and Matthias Weidlich. Event recognition challenges and techniques: Guest Editors’ introduction. ACM Transactions on Internet Technology (TOIT), 14 (1):1:1–1:??, July 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ardissono:2005:MAI

[AGPS05] Liliana Ardissono, Anna Goy, Giovanna Petrone, and Marino Segnan. A multi-agent infrastructure for developing personalized Web-based systems. ACM Transactions on Internet Technology (TOIT), 5(1):47– 69, February 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Abdelzaher:2020:FCC

[AHJ^{+}20] Tarek Abdelzaher, Yifan Hao, Kasthuri Jayarajah, Archan Misra, Per Skarin, Shuochao Yao, Dulanga Weerakoon, and Karl-Erik Årzén. Five challenges in cloud-enabled intelligence and control. ACM Transactions on Internet Technology (TOIT), 20 (1):3:1–3:19, March 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/abs/
                                        10.1145/3366021.

Abbassi:2014:DCC

[AHM14] Zeinab Abbassi, Nidhi Hegde, and Laurent Massoulié. Distributed content curation on the Web. ACM Transactions on Internet Technology (TOIT), 14(2–3): 9:1–9:??, October 2014.

CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Alhosban:2015:BFM

[AHM^{+}15] Amal Alhosban, Khayyam Hashmi, Zaki Malik, Brahim Medjahed, and Salima Benbernou. Bottomup fault management in service-based systems. ACM Transactions on Internet Technology (TOIT), 15(2):7:1–7:??, June 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Altman:2014:RNP

[AHS14] Eitan Altman, Manjesh Kumar Hanawal, and Rajesh Sundaresan. Regulation of off-network pricing in a nonneutral network. ACM Transactions on Internet Technology (TOIT), 14(2– 3):11:1–11:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Anthony:2003:DBA

[AJ03] Patricia Anthony and Nicholas R. Jennings. Developing a bidding agent for multiple heterogeneous auctions. ACM Transactions on Internet Technology (TOIT), 3(3): 185–217, August 2003. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Aron:2007:IIB

[AJP07] Ravi Aron, Siddarth Jayanty, and Praveen Pathak. Impact of Internet-based distributed monitoring systems on offshore sourcing of services. ACM Transactions on Internet Technology (TOIT), 7(3):16:1–16:??, August 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ardagna:2013:PUA

[AJSS13] Claudio A. Ardagna, Sushil Jajodia, Pierangela Samarati, and Angelos Stavrou. Providing users’ anonymity in mobile hybrid networks. ACM Transactions on Internet Technology (TOIT), 12 (3):7:1–7:??, May 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Alsirhani:2023:SLP

[AKA^{+}23] Amjad Alsirhani, Muhammad Ali Khan, Abdullah Alomari, Sauda Maryam, Aiman Younas, Muddesar Iqbal, Muhammad Hameed Siqqidi, and Amjad Ali. Securing low-power blockchainenabled IoT devices against energy depletion attack. ACM Transactions on Internet Technology (TOIT), 23(3): 43:1–43:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3511903.

Al-Khafajiy:2021:ICS

[AKOB^{+}21] Mohammed Al-Khafajiy, Safa Otoum, Thar Baker, Muhammad Asim, Zakaria Maamar, Moayad Aloqaily, Mark Taylor, and Martin Randles. Intelligent control and security of fog resources in healthcare systems via a cognitive fog model. ACM Transactions on Internet Technology (TOIT), 21(3): 54:1–54:23, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3382770.

Arlitt:2001:CSL

[AKR01] Martin Arlitt, Diwakar Krishnamurthy, and Jerry Rolia. Characterizing the scalability of a large web-based shopping system. ACM Transactions on Internet Technology (TOIT), 1(1):44–69, August 2001. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Anand:2007:GSE

[AKS07] Sarabjot Singh Anand, Patricia Kearney, and Mary Shapcott. Generating semantically enriched user profiles for Web personalization. ACM Transactions on Internet Technology (TOIT), 7(4):22:1–22:??, October 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Angarita:2019:USN

[ALA^{+}19] Rafael Angarita, Bruno Lefèvre, Shohreh Ahvar, Ehsan Ahvar, Nikolaos Georgantas, and Valérie Issarny. Universal social network bus: Toward the federation of heterogeneous online social network services. ACM Transactions on Internet Technology (TOIT), 19 (3):38:1–38:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Agichtein:2004:LFA

[ALG04] Eugene Agichtein, Steve Lawrence, and Luis Gravano. Learning to find answers to questions on the Web. ACM Transactions on Internet Technology (TOIT), 4(2):129– 162, May 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ahmed:2023:EPC

[ALS23] Usman Ahmed, Jerry Chun-Wei Lin, and Gautam Srivastava. Exploring the potential of cyber manufacturing system in the digital age. ACM Transactions on Internet Technology (TOIT), 23(4):54:1–54:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3596602.

Amiri:2003:ESI

[AM03] Ali Amiri and Syam Menon. Efficient scheduling of Internet banner advertisements. ACM Transactions on Internet Technology (TOIT), 3(4): 334–346, November 2003. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Anand:2007:IIT

[AM07] Sarabjot Singh Anand and Bamshad Mobasher. Introduction to intelligent techniques for Web personalization. ACM Transactions on Internet Technology (TOIT), 7(4):18:1–18:??, October 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Anonymous:2015:PDP

[Ano15] Anonymous. P-DONAS: a P2P-based domain name system in access networks. ACM Transactions on Internet Technology (TOIT), 15(3): 11:1–11:??, September 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Al-Otaibi:2022:STW

[AO22] Yasser D. Al-Otaibi. A shared two-way cybersecurity model for enhancing cloud service sharing for distributed user applications. ACM Transactions on Internet Technology (TOIT), 22(2): 47:1–47:17, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3430508.

Albrecht:2008:DIT

[AOVP08] Jeannie Albrecht, David Oppenheimer, Amin Vahdat, and David A. Patterson. Design and implementation trade-offs for wide-area resource discovery. ACM Transactions on Internet Technology (TOIT), 8(4): 18:1–18:??, September 2008. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Achara:2018:FGC

[APAC18] Jagdish Prasad Achara, Javier Parra-Arnau, and Claude Castelluccia. Finegrained control over tracking to support the ad-based Web economy. ACM Transactions on Internet Technology (TOIT), 18(4): 51:1–51:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Abeywickrama:2012:CAS

[AR12] Dhaminda B. Abeywickrama and Sita Ramakrishnan. Contextaware services engineering: Models, transformations, and verification. ACM Transactions on Internet Technology (TOIT), 11(3):10:1– 10:??, January 2012. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Abboura:2016:QBO

[ASBH^{+}16] Asma Abboura, Soror Sahri, Latifa Baba-Hamed, Mourad Ouziri, and Salima Benbernou. Quality-based online data reconciliation. ACM Transactions on Internet Technology (TOIT), 16(1):3:1–3:??, February 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Abououf:2022:MLM

[ASO^{+}22] Menatalla Abououf, Shakti Singh, Hadi Otrok, Rabeb Mizouni, and Ernesto Damiani. Machine learning in mobile crowd sourcing: a behaviorbased recruitment model. ACM Transactions on Internet Technology (TOIT), 22(1):16:1–16:28, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3451163.

Ayci:2023:UAP

[ASÖY23] Gonul Ayci, Murat Sensoy, Arzucan Özgür, and Pinar Yolum. Uncertaintyaware personal assistant for making personalized privacy decisions. ACM Transactions on Internet Technology (TOIT), 23(1):13:1–13:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3561820.

Ayman:2022:DDP

[ASW^{+}22] Afiya Ayman, Amutheezan Sivagnanam, Michael Wilbur, Philip Pugliese, Abhishek Dubey, and Aron Laszka. Datadriven prediction and optimization of energy use for transit fleets of electric and ICE vehicles. ACM Transactions on Internet Technology (TOIT), 22(1):7:1–7:29, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3433992.

Albrecht:2011:DAC

[ATB^{+}11] Jeannie Albrecht, Christopher Tuttle, Ryan Braud, Darren Dao, Nikolay Topilski, Alex C. Snoeren, and Amin Vahdat. Distributed application configuration, management, and visualization with plush. ACM Transactions on Internet Technology (TOIT), 11(2):6:1–6:??, December 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Avasalcai:2022:AMV

[ATD22] Cosmin Avasalcai, Christos Tsigkanos, and Schahram Dustdar. Adaptive management of volatile edge systems at runtime with satisfiability. ACM Transactions on Internet Technology (TOIT), 22(1):26:1–26:21, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3470658.

Alizadeh:2021:STB

[ATS^{+}21] Mojtaba Alizadeh, Mohammad Hesam Tadayon, Kouichi Sakurai, Hiroaki Anada, and Alireza Jolfaei. A secure ticket-based authentication mechanism for proxy mobile IPv6 networks in volunteer computing. ACM Transactions on Internet Technology (TOIT), 21(4): 82:1–82:16, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3407189.

Amato:2016:MOB

[AV16] Alba Amato and Salvatore Venticinque. Multiobjective optimization for brokering of multicloud service composition. ACM Transactions on Internet Technology (TOIT), 16(2): 13:1–13:??, April 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Alfonso:2017:TFM

[AVB17] Bexy Alfonso, Emilio Vivancos, and Vicente Botti. Toward formal modeling of affective agents in a BDI architecture. ACM Transactions on Internet Technology (TOIT), 17(1):5:1–5:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ale:2021:STB

[AZKG21] Laha Ale, Ning Zhang, Scott A. King, and Jose Guardiola. Spatiotemporal Bayesian learning for mobile edge computing resource planning in smart cities. ACM Transactions on Internet Technology (TOIT), 21(3):72:1–72:21, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3448613.

Bibi:2022:SDM

[BAM^{+}22] Iram Bibi, Adnan Akhunzada, Jahanzaib Malik, Muhammad Khurram Khan, and Muhammad Dawood. Secure distributed mobile volunteer computing with Android. ACM Transactions on Internet Technology (TOIT), 22(1):2:1–2:21, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3428151.

Bahutair:2023:EET

[BB23]  Mohammed  Bahutair

and Athman Bouguettaya. An end-to-end trust management framework for crowdsourced IoT services. ACM Transactions on Internet Technology (TOIT), 23(3): 46:1–46:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3600232.

Baldoni:2014:CBI

[BBC14] Matteo Baldoni, Cristina Baroglio, and Federico Capuzzimati. A commitmentbased infrastructure for programming socio-technical systems. ACM Transactions on Internet Technology (TOIT), 14(4): 23:1–23:??, December 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Benouaret:2014:WSC

[BBH^{+}14] Karim Benouaret, Djamal Benslimane, Allel Hadjali, Mahmoud Barhamgi, Zakaria Maamar, and Quan Z. Sheng. Web service compositions with fuzzy preferences: a graded dominance relationship-based approach. ACM Transactions on Internet Technology (TOIT), 13(4): 12:1–12:??, July 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Baroglio:2018:SIC

[BBP18] Cristina Baroglio, Olivier Boissier, and Axel Polleres. Special issue: Computational ethics and accountability. ACM Transactions on Internet Technology (TOIT), 18(4): 40:1–40:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bharti:2021:NMG

[BBS21] Vandana Bharti, Bhaskar Biswas, and Kaushal Kumar Shukla. A novel multiobjective GDWCN-PSO algorithm and its application to medical data security. ACM Transactions on Internet Technology (TOIT), 21(2):46:1–46:28, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3397679.

Blumenthal:2001:RDI

[BC01] Marjory S. Blumenthal and David D. Clark. Rethinking the design of the Internet: the end-to-end arguments vs. the brave new world. ACM Transactions on Internet Technology (TOIT), 1(1):70– 109, August 2001. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Budhi:2023:MTC

[BC23] Gregorius Satia Budhi and Raymond Chiong. A multi-type classifier ensemble for detecting fake reviews through textualbased feature extraction. ACM Transactions on Internet Technology (TOIT), 23(1):16:1– 16:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3568676.

Bou-Chaaya:2021:RTC

[BCCA^{+}21] Karam Bou-Chaaya, Richard Chbeir, Mansour Naser Alraja, Philippe Arnould, Charith Perera, Mahmoud Barhamgi, and Djamal Benslimane. δrisk: Toward contextaware multi-objective privacy management in connected environments. ACM Transactions on Internet Technology (TOIT), 21(2):51:1–51:31, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3418499.

Brambilla:2007:MDD

[BCF^{+}07] Marco Brambilla, Stefano Ceri, Federico Michele Facca, Irene Celino, Dario Cerizza, and Emanuele Della Valle. Model-driven design and development of semantic Web service applications. ACM Transactions on Internet Technology (TOIT), 8(1):3:1– 3:??, November 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bahri:2018:EAS

[BCFB18] Leila Bahri, Barbara Carminati, Elena Ferrari, and Andrea Bianco. Enhanced audit strategies for collaborative and accountable data sharing in social networks. ACM Transactions on Internet Technology (TOIT), 18 (4):44:1–44:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bohme:2018:SIE

[BCG^{+}18] Rainer Böhme, Richard Clayton, Jens Grossklags, Katrina Ligett, Patrick Loiseau, and Galina Schwartz. Special issue on the economics of security and privacy: Guest Editors’ introduction. ACM Transactions on Internet Technology (TOIT), 18(4): 47:1–47:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bertino:2016:ITI

[BCGN16] Elisa Bertino, Kim-Kwang Raymond Choo, Dimitrios Georgakopolous, and Surya Nepal. Internet of Things (IoT): Smart and secure service delivery. ACM Transactions on Internet Technology (TOIT), 16 (4):22:1–22:??, December 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bellavista:2006:MCM

[BCMS06] Paolo Bellavista, Antonio Corradi, Rebecca Montanari, and Cesare Stefanelli. A mobile computing middleware for location- and contextaware Internet data services. ACM Transactions on Internet Technology (TOIT), 6(4):356– 380, November 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bellini:2017:QRE

[BCN17] Emanuele Bellini, Paolo Ceravolo, and Paolo Nesi. Quantify resilience enhancement of UTS through exploiting connected community and Internet of everything emerging technologies. ACM Transactions on Internet Technology (TOIT), 18(1):7:1–7:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bicakci:2013:LSS

[BCO13] Kemal Bicakci, Bruno Crispo, and Gabriele Oligeri. LAKE: a server-side authenticated key-establishment with low computational workload. ACM Transactions on Internet Technology (TOIT), 13(2):5:1–5:??, December 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bond:2004:OAN

[BCP^{+}04] Gregory W. Bond, Eric Cheung, K. Hal Purdy, Pamela Zave, and J. Christopher Ramming. An open architecture for nextgeneration telecommunication services. ACM Transactions on Internet Technology (TOIT), 4(1): 83–123, February 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Brogi:2008:SBC

[BCP08] Antonio Brogi, Sara Corfini, and Razvan Popescu. Semanticsbased composition-oriented discovery of Web services. ACM Transactions on Internet Technology (TOIT), 8(4):19:1–19:??, September 2008. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bartoli:2010:FLS

[BDM10] Alberto Bartoli, Giorgio Davanzo, and Eric Medvet. A framework for large-scale detection of Web site defacements. ACM Transactions on Internet Technology (TOIT), 10 (3):10:1–10:??, October 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Boneh:2004:FGC

[BDT04] Dan Boneh, Xuhua Ding, and Gene Tsudik. Finegrained control of security capabilities. ACM Transactions on Internet Technology (TOIT), 4(1): 60–82, February 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Becerra-Fernandez:2006:SEW

[BF06] Irma Becerra-Fernandez. Searching for experts on the Web: a review of contemporary expertise locator systems. ACM Transactions on Internet Technology (TOIT), 6(4):333– 355, November 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Boldi:2021:FGN

[BG21] Paolo Boldi and Georgios Gousios. Finegrained network analysis for modern software ecosystems. ACM Transactions on Internet Technology (TOIT), 21(1): 1:1–1:14, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3418209.

Bai:2014:PTK

[BGK14] Xiao Bai, Rachid Guerraoui, and Anne-Marie Kermarrec. Personalizing top-k processing online in a peer-to-peer social tagging network. ACM Transactions on Internet Technology (TOIT), 13 (4):11:1–11:??, July 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bohte:2004:MBR

[BGL04] Sander M. Bohte, Enrico Gerding, and Han La Poutré. Market-based recommendation: Agents that compete for consumer attention. ACM Transactions on Internet Technology (TOIT), 4(4):

420–448, November 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bianchini:2005:IP

[BGS05] Monica Bianchini, Marco Gori, and Franco Scarselli. Inside PageRank. ACM Transactions on Internet Technology (TOIT), 5(1): 92–128, February 2005. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Barhamgi:2021:ISS

[BHPY21]  Mahmoud  Barhamgi,

Michael N. Huhns, Charith Perera, and Pinar Yolum. Introduction to the special section on humancentered security, privacy, and trust in the Internet of Things. ACM Transactions on Internet Technology (TOIT), 21(1):16:1–16:3, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3445790.

Billet:2017:SOP

[BI17] Benjamin Billet and Valérie Issarny. Spinel: an opportunistic proxy for connecting sensors to the Internet of Things. ACM Transactions on Internet Technology (TOIT), 17(2):21:1–21:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bandara:2015:PBM

[BJ15] H. M. N. Dilum Bandara and Anura P. Jayasumana. P2P-based, multi-attribute resource discovery under realworld resources and queries. ACM Transactions on Internet Technology (TOIT), 15(1): 5:1–5:??, February 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Braumandl:2003:QSI

[BKK03] R. Braumandl, A. Kemper, and D. Kossmann. Quality of service in an information economy. ACM Transactions on Internet Technology (TOIT), 3(4):291– 333, November 2003. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Blackburn:2014:COG

[BKS^{+}14] Jeremy Blackburn, Nicolas Kourtellis, John Skvoretz, Matei Ripeanu, and Adriana Iamnitchi. Cheating in online games: a social network perspective. ACM Transactions on Internet Technology (TOIT), 13(3):9:1–9:??, May 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Binmad:2017:IEO

[BL17] Ruchdee Binmad and Mingchu Li. Improving the efficiency of an online marketplace by incorporating forgiveness mechanism. ACM Transactions on Internet Technology (TOIT), 17(1):9:1–9:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bild:2015:ACU

[BLD^{+}15] David R. Bild, Yue Liu, Robert P. Dick, Z. Morley Mao, and Dan S. Wallach. Aggregate characterization of user behavior in Twitter and analysis of the retweet graph. ACM Transactions on Internet Technology (TOIT), 15(1): 4:1–4:??, February 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Benomar:2020:CBE

[BLMP20] Zakaria Benomar, Francesco Longo, Giovanni Merlino, and Antonio Puliafito. Cloud-based enabling mechanisms for container deployment and migration at the network edge. ACM Transactions on Internet Technology (TOIT), 20(3):25:1–25:28, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3380955.

Benomar:2022:CBN

[BLMP22] Zakaria Benomar, Francesco Longo, Giovanni Merlino, and Antonio Puliafito. Cloud-based network virtualization in IoT with OpenStack. ACM Transactions on Internet Technology (TOIT), 22(1): 19:1–19:26, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3460818.

Barga:2004:RGI

[BLSW04] Roger Barga, David Lomet, German Shegalov, and Gerhard Weikum. Recovery guarantees for Internet applications. ACM Transactions on Internet Technology (TOIT), 4(3):289– 328, August 2004. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Boerger:2022:EST

[BLTH22] Michell Boerger, Philipp Lämmel, Nikolay Tcholtchev, and Manfred Hauswirth. Enabling short-term energy flexibility markets through blockchain. ACM Transactions on Internet Technology (TOIT), 22(4):108:1–108:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3542949.

Baresi:2019:UMM

[BMG^{+}19] L. Baresi, D. F. Mendonça, M. Garriga, S. Guinea, and G. Quattrocchi. A unified model for the mobile-edge-cloud continuum. ACM Transactions on Internet Technology (TOIT), 19(2): 29:1–29:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
ft_gateway.cfm?id=3226644.

Brabrand:2002:BP

[BMS02] Claus Brabrand, Anders Møller, and Michael I. Schwartzbach. The < bigwig> project. ACM Transactions on Internet Technology (TOIT), 2(2): 79–114, May 2002. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Balsa:2017:TIC

[BPSD17] Ero Balsa, Cristina Pérez-Solà, and Claudia Diaz. Towards inferring communication patterns in online social networks. ACM Transactions on Internet Technology (TOIT), 17(3): 32:1–32:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Brinkmeier:2006:PR

[Bri06] Michael Brinkmeier. PageRank revisited. ACM Transactions on Internet Technology (TOIT), 6(3): 282–301, August 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Byers:2004:DAI

[BRK04] Simon Byers, Aviel D.

Rubin, and David Kormann. Defending against an Internet-based attack on the physical world. ACM Transactions on Internet Technology (TOIT), 4(3):239– 254, August 2004. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Borodin:2005:LAR

[BRRT05] Allan Borodin, Gareth O.

Roberts, Jeffrey S. Rosenthal, and Panayiotis Tsaparas. Link analysis ranking: algorithms, theory, and experiments. ACM Transactions on Internet Technology (TOIT), 5(1):231– 297, February 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Benslimane:2016:UWC

[BSBP16] Djamal Benslimane, Quan Z. Sheng, Mahmoud Barhamgi, and Henri Prade. The uncertain Web: Concepts, challenges, and current solutions. ACM Transactions on Internet Technology (TOIT), 16(1): 1:1–1:??, February 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Besen:2002:ECE

[BSS02] Stanley M. Besen, Jeffrey S. Spigel, and Padmanabhan Srinagesh. Evaluating the competitive effects of mergers of Internet backbone providers. ACM Transactions on Internet Technology (TOIT), 2(3):187– 204, August 2002. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Borges:2023:TIT

[BTC^{+}23] Pedro Victor Borges, Chantal Taconet, Sophie Chabridon, Denis Conan, Everton Cavalcante, and Thais Batista. Taming Internet of Things application development with the IoTvar middleware. ACM Transactions on Internet Technology (TOIT), 23(2): 29:1–29:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3586010.

Barakat:2022:RBF

[BTGM22] Lina Barakat, Phillip Taylor, Nathan Griffiths, and Simon Miles. A reputation-based framework for honest provenance reporting. ACM Transactions on Internet Technology (TOIT), 22 (4):103:1–103:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3507908.

Bates:2017:TCT

[BTH^{+}17] Adam Bates, Dave (Jing)

Tian, Grant Hernandez, Thomas Moyer, Kevin R. B. Butler, and Trent Jaeger. Taming the costs of trustworthy provenance through policy reduction. ACM Transactions on Internet Technology (TOIT), 17(4):

34:1–34:??, September 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bakker:2006:WAD

[BVT06] Arno Bakker, Maarten Van Steen, and Andrew S. Tanenbaum. A wide-area Distribution Network for free software. ACM Transactions on Internet Technology (TOIT), 6(3):259– 281, August 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Bing:2016:UEP

[BWL16] Lidong Bing, Tak-Lam Wong, and Wai Lam. Unsupervised extraction of popular product attributes from e-commerce Web sites by considering customer reviews. ACM Transactions on Internet Technology (TOIT), 16(2):12:1–12:??, April 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Baeza-Yates:2007:CNW

[BYCE07]  Ricardo  Baeza-Yates,

Carlos Castillo, and Efthimis N. Efthimiadis. Characterization of national Web domains. ACM Transactions on Internet Technology (TOIT), 7(2):9:1– 9:??, May 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Binns:2018:MTP

[BZVS18] Reuben Binns, Jun Zhao, Max Van Kleek, and Nigel Shadbolt. Measuring third-party tracker power across Web and mobile. ACM Transactions on Internet Technology (TOIT), 18(4): 52:1–52:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chicha:2021:UCM

[CAN^{+}21] Elie Chicha, Bechara Al Bouna, Mohamed Nassar, Richard Chbeir, Ramzi A. Haraty, Mourad Oussalah, Djamal Benslimane, and Mansour Naser Alraja. A user-centric mechanism for sequentially releasing graph datasets under Blowfish privacy. ACM Transactions on Internet Technology (TOIT), 21(1):20:1–20:25, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3431501.

Chopra:2014:ISI

[CAV14] Amit K. Chopra, Raian Ali, and Maja Vukovic. Introduction to the special issue on foundations of social computing. ACM Transactions on Internet Technology (TOIT), 14(4): 22:1–22:??, December 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Cooper:2015:NND

[CB15] Alissa Cooper and Ian Brown. Net neutrality: Discrimination, competition, and innovation in the UK and US. ACM Transactions on Internet Technology (TOIT), 15(1):2:1–2:??, February 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chegini:2023:DDW

[CBM23] Hossein Chegini, Fernando Beltran, and Aniket Mahanti. Designing and developing a weed detection model for California Thistle. ACM Transactions on Internet Technology (TOIT), 23(3):48:1–48:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3544491.

Caruccio:2023:MAI

[CCC^{+}23] Loredana Caruccio, Gaetano Cimino, Stefano Cirillo, Domenico Desiato, Giuseppe Polese, and Genoveffa Tortora. Malicious account identification in social network platforms. ACM Transactions on Internet Technology (TOIT), 23(4):57:1–57:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3625097.

Carpentieri:2022:PPS

[CCD^{+}22] Bruno Carpentieri, Arcangelo Castiglione, Alfredo De Santis, Francesco Palmieri, and Raffaele Pizzolante. Privacypreserving secure media streaming for multiuser smart environments. ACM Transactions on Internet Technology (TOIT), 22(2):32:1–32:21, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3423047.

Chu:2014:DDM

[CCJ^{+}14] Xiaowen Chu, Xiaowei Chen, Adele Lu Jia, Johan A. Pouwelse, and Dick H. J. Epema. Dissecting Darknets: Measurement and performance analysis. ACM Transactions on Internet Technology (TOIT), 13 (3):7:1–7:??, May 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chapman:2017:GEP

[CCM17] Adriane Chapman, James Cheney, and Simon Miles. Guest editorial: The provenance of online data. ACM Transactions on Internet Technology (TOIT), 17(4): 33:1–33:??, September 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Cascone:2021:WTR

[CCN^{+}21] Lucia Cascone, Aniello Castiglione, Michele Nappi, Fabio Narducci, and Ignazio Passero. Waiting for tactile: Robotic and virtual experiences in the fog. ACM Transactions on Internet Technology (TOIT), 21(3): 79:1–79:19, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3421507.

Chowdhury:2014:RWR

[CDC14] Soudip Roy Chowdhury, Florian Daniel, and Fabio Casati. Recommendation and weaving of reusable mashup model patterns for assisted development. ACM Transactions on Internet Technology (TOIT), 14(2– 3):21:1–21:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Challenger:2005:FBA

[CDIW05] Jim Challenger, Paul Dantzig, Arun Iyengar, and Karen Witting. A fragment-based approach for efficiently creating dynamic Web content. ACM Transactions on Internet Technology (TOIT), 5(2):359– 389, May 2005. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chen:2022:PDD

[CDJ^{+}22] Haipeng Chen, Andrew Duncklee, Sushil Jajodia, Rui Liu, Sean Mcnamara, and V. S. Subrahmanian. PCAM: a data-driven probabilistic cyber-alert management framework. ACM Transactions on Internet Technology (TOIT), 22(3): 67:1–67:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3511101.

Chen:2010:DVS

[CDM10] Teh-Chung Chen, Scott Dick, and James Miller. Detecting visually similar Web pages: Application to phishing detection. ACM Transactions on Internet Technology (TOIT), 10(2):5:1–5:??, May 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Coucheney:2014:ISN

[CDM^{+}14] Pierre Coucheney, Giuseppe D’acquisto, Patrick Maillé, Maurizio Naldi, and Bruno Tuffin. Influence of search neutrality on the economics of advertisement-financed content. ACM Transactions on Internet Technology (TOIT), 14(2– 3):10:1–10:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ceri:2007:MDD

[CDMF07] Stefano Ceri, Florian Daniel, Maristella Matera, and Federico M. Facca. Model-driven development of contextaware Web applications. ACM Transactions on Internet Technology (TOIT), 7(1):2:1– 2:??, February 2007. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Clavel:2017:AIA

[CDPR17] Chloé Clavel, Rossana Damiano, Viviana Patti, and Paolo Rosso. Affect and interaction in agent-based systems and social media: Guest Editors’ introduction. ACM Transactions on Internet Technology (TOIT), 17 (1):1:1–1:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Can:2021:PPF

[CE21] Yekta Said Can and Cem Ersoy. Privacy-preserving federated deep learning for wearable IoTbased biomedical monitoring. ACM Transactions on Internet Technology (TOIT), 21(1): 21:1–21:17, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3428152.

Cherkasova:2003:MCE

[CFTV03] Ludmila Cherkasova, Yun Fu, Wenting Tang, and Amin Vahdat. Measuring and characterizing end-to-end Internet service performance. ACM Transactions on Internet Technology (TOIT), 3(4): 347–391, November 2003. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chavhan:2022:ECA

[CGG^{+}22] Suresh Chavhan, Deepak Gupta, Sarada Prasad Gochhayat, Chandana B. N., Ashish Khanna, K. Shankar, and Joel J. P. C. Rodrigues. Edge computing AI-IoT integrated energy-efficient intelligent transportation system for smart cities. ACM Transactions on Internet Technology (TOIT), 22(4):106:1– 106:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3507906.

Courcoubetis:2014:SIP

[CGL^{+}14]  Costas  Courcoubetis,

Roch Guérin, Patrick Loiseau, David Parkes, Jean Walrand, and Adam Wierman. Special issue on pricing and incentives in networks and systems: Guest Editors’ introduction. ACM Transactions on Internet Technology (TOIT), 14(2–3): 8:1–8:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Courcoubetis:2016:NPP

[CGL^{+}16]  Costas  Courcoubetis,

Laszlo Gyarmati, Nikolaos Laoutaris, Pablo Rodriguez, and Kostas Sdrolias. Negotiating premium peering prices: a quantitative model with applications. ACM Transactions on Internet Technology (TOIT), 16(2):14:1–14:??, April 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Cho:2003:EFC

[CGM03] Junghoo Cho and Hector Garcia-Molina. Estimating frequency of change. ACM Transactions on Internet Technology (TOIT), 3(3):256– 290, August 2003. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Cho:2006:SWC

[CGMH^{+}06] Junghoo Cho, Hector Garcia-Molina, Taher Haveliwala, Wang Lam, Andreas Paepcke, Sriram Raghavan, and Gary Wesley. Stanford Web-Base components and applications. ACM Transactions on Internet Technology (TOIT), 6(2):153– 186, May 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chen:2023:DSI

[CGS23] Keke Chen, Yuechun Gu, and Sagar Sharma. DisguisedNets: Secure image outsourcing for confidential model training in clouds. ACM Transactions on Internet Technology (TOIT), 23(3): 47:1–47:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3609506.

Choo:2021:ISI

[CGT^{+}21] Kkwang Raymond Choo, Uttam Ghosh, Deepak Tosh, Reza M. Parizi, and Ali Dehghantanha. Introduction to the special issue on Decentralized Blockchain Applications and Infrastructures for Next Generation Cyber-Physical Systems. ACM Transactions on Internet Technology (TOIT), 21(2): 38e:1–38e:3, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3464768.

Chen:2005:FCM

[CH05] Xuan Chen and John Heidemann. Flash crowd mitigation via adaptive admission control based on application-level observations. ACM Transactions on Internet Technology (TOIT), 5(3):532– 569, August 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chen:2021:SSQ

[CHC^{+}21] Yan-Chun Chen, Ren-Hung Hwang, Mu-Yen Chen, Chih-Chin Wen, and Chih-Ping Hsu. Screw slot quality inspection system based on tactile network. ACM Transactions on Internet Technology (TOIT), 21 (4):90:1–90:17, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3423556.

Chaudhry:2021:RBP

[CIY^{+}21] Shehzad Ashraf Chaudhry, Azeem Irshad, Khalid Yahya, Neeraj Kumar, Mamoun Alazab, and Yousaf Bin Zikria. Rotating behind privacy: an improved lightweight authentication scheme for cloud-based IoT environment. ACM Transactions on Internet Technology (TOIT), 21(3): 78:1–78:19, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3425707.

Chen:2023:DPP

[CJW^{+}23] Jing Chen, Wenjun Jiang, Jie Wu, Kenli Li, and Keqin Li. Dynamic personalized POI sequence recommendation with fine-grained contexts. ACM Transactions on Internet Technology (TOIT), 23(2): 32:1–32:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3583687.

Caragiannis:2014:RGG

[CKKK14] Ioannis Caragiannis, Christos Kaklamanis, Panagiotis Kanellopoulos, and Maria Kyropoulou. Revenue guarantees in the generalized second price auction. ACM Transactions on Internet Technology (TOIT), 14(2– 3):17:1–17:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chen:2024:NCD

[CL24] Yi-Cheng Chen and Wang-Chien Lee. A novel cross-domain recommendation with evolution learning. ACM Transactions on Internet Technology (TOIT), 24(1):

6:1–6:??, February 2024. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3639567.

Chen:2019:DSM

[CLF^{+}19] Min Chen, Wei Li, Giancarlo Fortino, Yixue Hao, Long Hu, and Iztok Humar. A dynamic service migration mechanism in edge cognitive computing. ACM Transactions on Internet Technology (TOIT), 19(2):30:1–30:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3239565.

Chen:2024:EDW

[CLFX24] Yanming Chen, Tong Luo, Weiwei Fang, and Neal. N. Xiong. EdgeCI: Distributed workload assignment and model partitioning for CNN inference on edge clusters. ACM Transactions on Internet Technology (TOIT), 24(2): 10:1–10:??, May 2024. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3656041.

Cheng:2021:AGS

[CLJ^{+}21] Lichen Cheng, Jiqiang Liu, Yi Jin, Yidong Li, and Wei Wang. Account guarantee scheme: Making anonymous accounts supervised in blockchain. ACM Transactions on Internet Technology (TOIT), 21(1): 11:1–11:19, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3406092.

Chen:2023:UDL

[CLL23] Mu-Yen Chen, Yi-Wei Lai, and Jiunn-Woei Lian. Using deep learning models to detect fake news about COVID-19. ACM Transactions on Internet Technology (TOIT), 23(2): 25:1–25:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3533431.

Crainiceanu:2011:LBR

[CLM^{+}11] Adina Crainiceanu, Prakash Linga, Ashwin Machanavajjhala, Johannes Gehrke, and Jayavel Shanmugasundaram. Load balancing and range queries in P2P systems using P-Ring. ACM Transactions on Internet Technology (TOIT), 10(4): 16:1–16:??, March 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Concone:2019:FBA

[CLM19] Federico Concone, Giuseppe Lo Re, and Marco Morana. A fog-based application for human activity recognition using personal smart devices. ACM Transactions on Internet Technology (TOIT), 19(2): 20:1–20:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
ft_gateway.cfm?id=3266142.

Chan:2005:MPC

[CLN05] Addison Chan, Rynson W. H. Lau, and Beatrice Ng. Motion prediction for caching and prefetching in mouse-driven DVE navigation. ACM Transactions on Internet Technology (TOIT), 5(1):70– 91, February 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chen:2022:MBJ

[CLS^{+}22] Guihong Chen, Xi Liu, Mohammad Shorfuzzaman, Ali Karime, Yonghua Wang, and Yuanhang Qi. MEC-based jammingaided anti-eavesdropping with deep reinforcement learning for WBANs. ACM Transactions on Internet Technology (TOIT), 22(3):60:1–60:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3453186.

Chen:2022:DDI

[CLW^{+}22] Chen Chen, Lei Liu, Shaohua Wan, Xiaozhe Hui, and Qingqi Pei. Data dissemination for Industry 4.0 applications in Internet of Vehicles based on short-term traffic prediction. ACM Transactions on Internet Technology (TOIT), 22(1):3:1–3:18, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3430505.

Chen:2020:UEG

[CLZ^{+}20] Ting Chen, Zihao Li, Yuxiao Zhu, Jiachi Chen, Xiapu Luo, John Chi-Shing Lui, Xiaodong Lin, and Xiaosong Zhang. Understanding Ethereum via graph analysis. ACM Transactions on Internet Technology (TOIT), 20(2):18:1–18:32, May 2020. CODEN ????

ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/doi/abs/10.1145/
3381036.

Choudhury:2022:DND

[CMML22] Nikumani Choudhury, Rakesh Matam, Mithun Mukherjee, and Jaime Lloret. DADC: a novel duty-cycling scheme for IEEE 802.15.4 clustertree-based IoT applications. ACM Transactions on Internet Technology (TOIT), 22(2): 30:1–30:26, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3409487.

Copil:2016:RFS

[CMTD16] Georgiana Copil, Daniel Moldovan, Hong-Linh Truong, and Schahram Dustdar. rSYBL: a framework for specifying and controlling cloud services elasticity. ACM Transactions on Internet Technology (TOIT), 16(3):18:1–18:??, August 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Charmet:2024:VVP

[CMTT24] Fabien Charmet, Tomohiro Morikawa, Akira Tanaka, and Takeshi Takahashi. VORTEX:

Visual phishing detectiOns aRe Through EXplanations. ACM Transactions on Internet Technology (TOIT), 24(2): 9:1–9:??, May 2024. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3654665.

Costa:2016:MBC

[CO16] Gianni Costa and Riccardo Ortale. Modelbased collaborative personalized recommendation on signed social rating networks. ACM Transactions on Internet Technology (TOIT), 16(3):20:1–20:??, August 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Cooley:2003:UWS

[Coo03]  Robert Cooley.  The

use of web structure and content to identify subjectively interesting web usage patterns. ACM Transactions on Internet Technology (TOIT), 3(2): 93–116, May 2003. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chen:2021:PSD

[CPL^{+}21] Liang Chen, Jiaying Peng, Yang Liu, Jintang Li, Fenfang Xie, and Zibin Zheng. Phishing scams detection in Ethereum transaction network. ACM Transactions on Internet Technology (TOIT), 21(1):10:1–10:16, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3398071.

Claessens:2003:HCM

[CPV03] Joris Claessens, Bart Preneel, and Joos Vandewalle. (how) can mobile agents do secure electronic transactions on untrusted hosts? A survey of the security issues and the current solutions. ACM Transactions on Internet Technology (TOIT), 3(1):28– 48, February 2003. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Cao:2016:MMR

[CPV^{+}16] Tien-Dung Cao, Tran-Vu Pham, Quang-Hieu Vu, Hong-Linh Truong, Duc-Hung Le, and Schahram Dustdar. MARSA: a marketplace for realtime human sensing data. ACM Transactions on Internet Technology (TOIT), 16(3): 16:1–16:??, August 2016.

CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Codetta-Raiteri:2017:DNS

[CRP17] Daniele Codetta-Raiteri and Luigi Portinale. Decision networks for security risk assessment of critical infrastructures. ACM Transactions on Internet Technology (TOIT), 18(3): 29:1–29:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Coyle:2007:SIW

[CS07] Maurice Coyle and Barry Smyth. Supporting intelligent Web search. ACM Transactions on Internet Technology (TOIT), 7(4):20:1–20:??, October 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Colazzo:2009:DCS

[CS09] Dario Colazzo and Carlo Sartiani. Detection of corrupted schema mappings in XML data integration systems. ACM Transactions on Internet Technology (TOIT), 9(4):14:1–14:??, September 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ciesielczyk:2017:PRI

[CSMM17] Michal Ciesielczyk, Andrzej Szwabe, Mikolaj Morzy, and Pawel Misiorek. Progressive random indexing: Dimensionality reduction preserving local network dependencies. ACM Transactions on Internet Technology (TOIT), 17(2): 20:1–20:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chopra:2017:ISI

[CSS17] Amit K. Chopra, Erez Shmueli, and Vivek K. Singh. Introduction to the special issue on advances in social computing. ACM Transactions on Internet Technology (TOIT), 17(2): 11:1–11:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Cheng:2020:CCH

[CSS20] Xia Cheng, Junyang Shi, and Mo Sha. Cracking channel hopping sequences and graph routes in industrial TSCH networks. ACM Transactions on Internet Technology (TOIT), 20(3): 23:1–23:28, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3372881.

Chen:2022:NIM

[CSW^{+}22] Min Chen, Ke Shen, Rui Wang, Yiming Miao, Yingying Jiang, Kai Hwang, Yixue Hao, Guangming Tao, Long Hu, and Zhongchun Liu. Negative information measurement at AI edge: a new perspective for mental health monitoring. ACM Transactions on Internet Technology (TOIT), 22(3): 62:1–62:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3471902.

Carstens:2017:UAI

[CT17] Lucas Carstens and Francesca Toni. Using argumentation to improve classification in natural language problems. ACM Transactions on Internet Technology (TOIT), 17 (3):30:1–30:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chhetri:2024:THA

[CTS^{+}24] Mohan Baruwal Chhetri, Shahroz Tariq, Ronal Singh, Fatemeh Jalalvand, Cecile Paris, and Surya Nepal. Towards human-AI teaming to mitigate alert fatigue in security operations centres. ACM Transactions on Internet Technology (TOIT), 24(3): 12:1–12:??, 2024. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3670009.

Chien:2006:SCQ

[CTZZ06] Shu-Yao Chien, Vassilis J. Tsotras, Carlo Zaniolo, and Donghui Zhang. Supporting complex queries on multiversion XML documents. ACM Transactions on Internet Technology (TOIT), 6(1):53– 84, February 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chau:2014:EVP

[CWC14] Chi-Kin Chau, Qian Wang, and Dah-Ming Chiu. Economic viability of Paris Metro Pricing for digital services. ACM Transactions on Internet Technology (TOIT), 14 (2–3):12:1–12:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Chen:2019:IMC

[CWLZ19] Yanjiao Chen, Xu Wang, Baochun Li, and Qian Zhang. An incentive mechanism for crowdsourcing systems with network effects. ACM Transactions on Internet Technology (TOIT), 19(4):49:1–49:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

                                        URL https://dl.acm.
                                        org/ft_gateway.cfm?
                                        id=3347514.

Cao:2021:UDK

[CWW^{+}21] Bin Cao, Jiawei Wu, Sichao Wang, Honghao Gao, Jing Fan, Shuiguang Deng, Jianwei Yin, and Xuan Liu. Unsupervised derivation of keyword summary for short texts. ACM Transactions on Internet Technology (TOIT), 21(2): 45:1–45:23, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3397162.

Chiu:2021:RGB

[CXG21] David K. Y. Chiu, Tao Xu, and Iker Gondra. Random graph-based multiple instance learning for structured IoT smart city applications. ACM Transactions on Internet Technology (TOIT), 21(3):70:1–70:17, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3448611.

Chen:2021:CWR

[CXH^{+}21] Min Chen, Wenjing Xiao, Long Hu, Yujun Ma, Yin Zhang, and Guangming Tao. Cognitive wearable robotics for autism perception enhancement. ACM Transactions on Internet Technology (TOIT), 21(4): 97:1–97:16, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3450630.

Cui:2021:IVS

[CXW^{+}21] Laizhong Cui, Zhe Xiao, Jiahao Wang, Fei Chen, Yi Pan, Hua Dai, and Jing Qin. Improving vaccine safety using blockchain. ACM Transactions on Internet Technology (TOIT), 21(2): 38:1–38:24, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3388446.

Chen:2020:ORP

[CYD^{+}20] Jiaoyan Chen, Laurence T. Yang, Xianjun Deng, Xianggong Hong, and Lingzhi Yi. Optimal receiver placement for K-barrier coverage in passive bistatic radar sensor networks. ACM Transactions on Internet Technology (TOIT), 20(3):24:1–24:23, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3377402.

Chen:2021:TEE

[CYG^{+}21] Wu Chen, Yong Yu, Keke Gai, Jiamou Liu, and Kim-Kwang Raymond Choo. Time-efficient ensemble learning with sample exchange for edge computing. ACM Transactions on Internet Technology (TOIT), 21(3): 76:1–76:17, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3409265.

Chen:2022:EEE

[CYWW22] Long Chen, Mianyang Yao, Yalan Wu, and Jigang Wu. EECDN: Energy-efficient cooperative DNN edge inference in wireless sensor networks. ACM Transactions on Internet Technology (TOIT), 22 (4):109:1–109:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3544969.

Ciortea:2022:ISI

[CZPS22] Andrei Ciortea, Xiaomin Zhu, Calton Pu, and Munindar P. Singh. Introduction to the special issue on multiagent systems and services in the Internet of Things. ACM Transactions on Internet Technology (TOIT), 22(4):99:1–99:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3584744.

Doncel:2014:PAR

[DABP14] Josu Doncel, Urtzi Ayesta, Olivier Brun, and Balakrishna Prabhu. Is the price of anarchy the right measure for loadbalancing games? ACM Transactions on Internet Technology (TOIT), 14 (2–3):18:1–18:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Dalal:2011:UPQ

[Dal11] Amy Csizmar Dalal.

User-perceived quality assessment of streaming media using reduced feature sets. ACM Transactions on Internet Technology (TOIT), 11(2): 8:1–8:??, December 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Dacosta:2012:OTC

[DCAT12] Italo Dacosta, Saurabh Chakradeo, Mustaque Ahamad, and Patrick Traynor. One-time cookies: Preventing session hijacking attacks with stateless authentication tokens. ACM Transactions on Internet Technology (TOIT), 12(1): 1:1–1:??, June 2012. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Deng:2021:STL

[DCD^{+}21] Song Deng, Fulin Chen, Xia Dong, Guangwei Gao, and Xindong Wu. Short-term load forecasting by using improved GEP and abnormal load recognition. ACM Transactions on Internet Technology (TOIT), 21(4): 95:1–95:28, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3447513.

Dailey:2022:ADN

[DCL^{+}22] Ryan Dailey, Aniesh Chawla, Andrew Liu, Sripath Mishra, Ling Zhang, Josh Majors, Yung-Hsiang Lu, and George K. Thiruvathukal. Automated discovery of network cameras in heterogeneous Web pages. ACM Transactions on Internet Technology (TOIT), 22(1):15:1–15:25, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3450629.

Deng:2021:IDC

[DCZ^{+}21] Shuiguang Deng, Guanjie Cheng, Hailiang Zhao, Honghao Gao, and Jianwei Yin. Incentive-driven computation offloading in blockchain-enabled ecommerce. ACM Transactions on Internet Technology (TOIT), 21(1): 9:1–9:19, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397160.

Dossani:2007:IRO

[DD07] Rafiq Dossani and Nathan Denny. The Internet’s role in offshored services: a case study of India. ACM Transactions on Internet Technology (TOIT), 7(3):15:1–15:??, August 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

DeMeo:2019:TPM

[De 19] Pasquale De Meo. Trust prediction via matrix factorisation. ACM Transactions on Internet Technology (TOIT), 19(4): 44:1–44:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

                                        URL https://dl.acm.
                                        org/ft_gateway.cfm?
                                        id=3323163.

Salve:2023:ATM

[DFL^{+}23] Andrea De Salve, Luca Franceschi, Andrea Lisi, Paolo Mori, and Laura Ricci. L2DART: a trust management system integrating blockchain and off-chain computation. ACM Transactions on Internet Technology (TOIT), 23(1):14:1–14:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3561386.

Dopmann:2022:OBP

[DFLT22] Christoph Döpmann, Felix Fiedler, Sergio Lucia, and Florian Tschorsch.

Optimization-based predictive congestion control for the Tor network: Opportunities and challenges. ACM Transactions on Internet Technology (TOIT), 22(4):97:1–97:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3520440.

Dandekar:2015:SFC

[DGWW15] Pranav Dandekar, Ashish Goel, Michael P. Wellman, and Bryce Wiedenbeck. Strategic formation of credit networks. ACM Transactions on Internet Technology (TOIT), 15(1):3:1–3:??, February 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Dai:2015:EDC

[DJ15] Wei Dai and Scott Jordan. The effect of data caps upon ISP service tier design and users. ACM Transactions on Internet Technology (TOIT), 15 (2):8:1–8:??, June 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Deshpande:2004:SMM

[DK04] Mukund Deshpande and George Karypis. Selective Markov models for predicting Web page accesses. ACM Transactions on Internet Technology (TOIT), 4(2):163– 184, May 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Dill:2002:SSW

[DKM^{+}02] Stephen Dill, Ravi Kumar, Kevin S. Mccurley, Sridhar Rajagopalan, D. Sivakumar, and Andrew Tomkins. Selfsimilarity in the web. ACM Transactions on Internet Technology (TOIT), 2(3):205–223, August 2002. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Dikaiakos:2012:MSR

[DKP12] Marios D. Dikaiakos, Asterios Katsifodimos, and George Pallis. Minersoft: Software retrieval in grid and cloud computing infrastructures. ACM Transactions on Internet Technology (TOIT), 12 (1):2:1–2:??, June 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Damiani:2017:EOS

[DKP17] Ernesto Damiani, Ryszard Kowalczyk, and Gerard Parr. Extending the outreach: From smart cities to connected communities. ACM Transactions on Internet Technology (TOIT), 18(1):1:1–1:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Donato:2007:WGH

[DLLM07] Debora Donato, Luigi Laura, Stefano Leonardi, and Stefano Millozzi. The Web as a graph: How far we are. ACM Transactions on Internet Technology (TOIT), 7(1):4:1– 4:??, February 2007. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Duan:2016:SDC

[DLZ^{+}16] Li Duan, Dongxi Liu, Yang Zhang, Shiping Chen, Ren Ping Liu, Bo Cheng, and Junliang Chen. Secure data-centric access control for smart grid services based on publish/ subscribe systems. ACM Transactions on Internet Technology (TOIT), 16 (4):23:1–23:??, December 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Meo:2017:UCM

[DMGR^{+}17] Pasquale De Meo, Katarzyna Musial-Gabrys, Domenico Rosaci, Giuseppe M. L.

Sarnè, and Lora Aroyo. Using centrality measures to predict helpfulnessbased reputation in trust networks. ACM Transactions on Internet Technology (TOIT), 17(1): 8:1–8:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ding:2007:ESD

[DMT07] Xuhua Ding, Daniele Mazzocchi, and Gene Tsudik. Equipping smart devices with public key signatures. ACM Transactions on Internet Technology (TOIT), 7(1):3:1– 3:??, February 2007. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Dustdar:2019:ISS

[DNJ19] Schahram Dustdar, Surya Nepal, and James Joshi. Introduction to the special section on advances in Internet-based collaborative technologies. ACM Transactions on Internet Technology (TOIT), 19 (3):37:1–37:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Daaji:2022:MCW

[DOG^{+}22] Marwa Daaji, Ali Ouni, Mohamed Mohsen Gammoudi, Salah Bouktif, and Mohamed Wiem Mkaouer. Multi-criteria Web services selection: Balancing the quality of design and quality of service. ACM Transactions on Internet Technology (TOIT), 22(1):12:1– 12:31, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3446388.

Dastani:2017:OCM

[DP17] Mehdi Dastani and Alexander Pankov. Othercondemning moral emotions: Anger, contempt and disgust. ACM Transactions on Internet Technology (TOIT), 17(1): 4:1–4:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Das:2016:CWM

[DPCM16] Aveek K. Das, Parth H.

Pathak, Chen-Nee Chuah, and Prasant Mohapatra. Characterization of wireless multidevice users. ACM Transactions on Internet Technology (TOIT), 16(4): 29:1–29:??, December 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Dan:2022:IGT

[DPD22] Ovidiu Dan, Vaibhav Parikh, and Brian D. Davison. IP geolocation through reverse DNS. ACM Transactions on Internet Technology (TOIT), 22(1):17:1– 17:29, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3457611.

Diaz:2005:PSR

[DR05] Oscar Diaz and Juan J. Rodriguez. Portlet syndication: Raising variability concerns. ACM Transactions on Internet Technology (TOIT), 5(4): 627–659, November 2005. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Dong:2023:PIS

[DRC^{+}23] Yucheng Dong, Qin Ran, Xiangrui Chao, Congcong Li, and Shui Yu. Personalized individual semantics learning to support a large-scale linguistic consensus process. ACM Transactions on Internet Technology (TOIT), 23(2): 26:1–26:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3533432.

David:2007:ODE

[DRJ^{+}07] Esther David, Alex Rogers, Nicholas R. Jennings, Jeremy Schiff, Sarit Kraus, and Michael H. Rothkopf. Optimal design of English auctions with discrete bid levels. ACM Transactions on Internet Technology (TOIT), 7(2):12:1–12:??, May 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Dolog:2008:PAL

[DSNK08] Peter Dolog, Bernd Simon, Wolfgang Nejdl, and Tomaž Klobučar. Personalizing access to learning networks. ACM Transactions on Internet Technology (TOIT), 8(2): 3:1–3:??, February 2008. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Das:2019:PAH

[DSVA19] Saptarshi Das, Shamik Sural, Jaideep Vaidya, and Vijayalakshmi Atluri. Policy adaptation in hierarchical attributebased access control systems. ACM Transactions on Internet Technology (TOIT), 19(3): 40:1–40:??, November 2019. CODEN ????

ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3323233.

Drosatos:2017:PET

[DTE17] George Drosatos, Aimilia Tasidou, and Pavlos S. Efraimidis. Privacyenhanced television audience measurements. ACM Transactions on Internet Technology (TOIT), 17(1):10:1–10:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Doan:2022:EVC

[DvRDHB22] Trinh Viet Doan, Roland van Rijswijk-Deij, Oliver Hohlfeld, and Vaibhav Bajpai. An empirical view on consolidation of the Web. ACM Transactions on Internet Technology (TOIT), 22(3): 70:1–70:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3503158.

Dong:2024:DMS

[DWF24] Chao Dong, Fang Wang, and Dan Feng. Dx-Hash: a memory-saving consistent hashing algorithm. ACM Transactions on Internet Technology (TOIT), 24(1): 3:1–3:??, February 2024.

CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3631708.

Dai:2023:EAC

[DWGC23] Yuanjun Dai, An Wang, Yang Guo, and Songqing Chen. Elastically augmenting the control-path throughput in SDN to deal with Internet DDoS attacks. ACM Transactions on Internet Technology (TOIT), 23(1): 9:1–9:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3559759.

Deng:2023:USS

[DXP^{+}23] Lizhen Deng, Guoxia Xu, Jiaqi Pi, Hu Zhu, and Xiaokang Zhou. Unpaired self-supervised learning for industrial cybermanufacturing spectrum blind deconvolution. ACM Transactions on Internet Technology (TOIT), 23(4):52:1–52:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3590963.

Do:2016:CMD

[DZHV16] Ngoc Do, Ye Zhao, Cheng-Hsin Hsu, and Nalini Venkatasubramanian. Crowdsourced mobile data transfer with delay bound. ACM Transactions on Internet Technology (TOIT), 16 (4):28:1–28:??, December 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Enguehard:2020:ELC

[EDC20] Marcel Enguehard, Yoann Desmouceaux, and Giovanna Carofiglio. Efficient latency control in fog deployments via hardware-accelerated popularity estimation. ACM Transactions on Internet Technology (TOIT), 20(3):21:1–21:23, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3366020.

Eshuis:2019:RAP

[EHY19] Rik Eshuis, Richard Hull, and Mengfei Yi. Reasoning about property preservation in adaptive case management. ACM Transactions on Internet Technology (TOIT), 19(1):12:1–12:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ehrenkranz:2009:SIS

[EL09] Toby Ehrenkranz and Jun Li. On the state of IP spoofing defense. ACM Transactions on Internet Technology (TOIT), 9(2): 6:1–6:??, May 2009. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Estuka:2019:PVA

[EM19] Fadwa Estuka and James Miller. A pure visual approach for automatically extracting and aligning structured Web data. ACM Transactions on Internet Technology (TOIT), 19(4): 51:1–51:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3365376.

Estrada-Torres:2019:MPK

[ETRDRO^{+}19] Bedilia Estrada-Torres, Pedro Henrique Piccoli Richetti, Adela Del-Río-Ortega, Fernanda Araujo Baião, Manuel Resinas, Flávia Maria Santoro, and Antonio Ruiz-Cortés. Measuring performance in knowledge-intensive processes. ACM Transactions on Internet Technology (TOIT), 19(1): 15:1–15:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Eirinaki:2003:WMW

[EV03] Magdalini Eirinaki and Michalis Vazirgiannis. Web mining for web personalization. ACM Transactions on Internet Technology (TOIT), 3 (1):1–27, February 2003. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Eirinaki:2007:WSP

[EV07] Magdalini Eirinaki and Michalis Vazirgiannis. Web site personalization based on link analysis and navigational patterns. ACM Transactions on Internet Technology (TOIT), 7(4):21:1–21:??, October 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Figueiredo:2014:DSM

[FAGB14] Flavio Figueiredo, Jussara M. Almeida, Marcos André Gonçalves, and Fabrício Benevenuto. On the dynamics of social media popularity: a YouTube case study. ACM Transactions on Internet Technology (TOIT), 14(4): 24:1–24:??, December 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ferry:2018:CMD

[FCS^{+}18] Nicolas Ferry, Franck Chauvel, Hui Song, Alessandro Rossini, Maksym Lushpenko, and Arnor Solberg. CloudMF: Model-driven management of multi-cloud applications. ACM Transactions on Internet Technology (TOIT), 18(2): 16:1–16:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Flake:2004:GEMa

[FFGM04a] Gary William Flake, Paolo Frasconi, C. Lee Giles, and Marco Maggini. Guest editorial: Machine learning for the Internet. ACM Transactions on Internet Technology (TOIT), 4(2):125– 128, May 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Flake:2004:GEMb

[FFGM04b] Gary William Flake, Paolo Frasconi, C. Lee Giles, and Marco Maggini. Guest editorial: Machine learning for the Internet. ACM Transactions on Internet Technology (TOIT), 4(4):341– 343, November 2004. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Felemban:2019:TMD

[FFKG19] Muhamad Felemban, Emad Felemban, Jason Kobes, and Arif Ghafoor. Threat management in datacentric IoT-based collaborative systems. ACM Transactions on Internet Technology (TOIT), 19 (3):42:1–42:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Fazzinga:2009:RXD

[FFP09] Bettina Fazzinga, Sergio Flesca, and Andrea Pugliese. Retrieving XML data from heterogeneous sources through vague querying. ACM Transactions on Internet Technology (TOIT), 9(2): 7:1–7:??, May 2009. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Faraci:2020:FCU

[FGS20] Giuseppe Faraci, Christian Grasso, and Giovanni Schembra. Fog in the clouds: UAVs to provide edge computing to IoT devices. ACM Transactions on Internet Technology (TOIT), 20(3):26:1–26:26, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3382756.

Feng:2012:VCC

[FLD12] Qinyuan Feng, Ling Liu, and Yafei Dai. Vulnerabilities and countermeasures in context-aware social rating services. ACM Transactions on Internet Technology (TOIT), 11 (3):11:1–11:??, January 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Fenner:2006:SME

[FLL06] Trevor Fenner, Mark Levene, and George Loizou. A stochastic model for the evolution of the Web allowing link deletion. ACM Transactions on Internet Technology (TOIT), 6(2):117– 130, May 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Fischer:2022:ISS

[FLLM22] Mathias Fischer, Winfried Lamersdorf, Jörg Liebeherr, and Max Mühlhäuser. Introduction to the special section on recent advances in networks and distributed systems. ACM Transactions on Internet Technology (TOIT), 22(4):93:1–93:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3584743.

Fouquet:2023:BBQ

[FLR23] Romain Fouquet, Pierre Laperdrix, and Romain Rouvoy. Breaking bad: Quantifying the addiction of Web elements to JavaScript. ACM Transactions on Internet Technology (TOIT), 23(1):22:1–22:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579846.

Ferretti:2019:FBS

[FMC19] Luca Ferretti, Mirco Marchetti, and Michele Colajanni. Fog-based secure communications for low-power IoT devices. ACM Transactions on Internet Technology (TOIT), 19(2): 27:1–27:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
ft_gateway.cfm?id=3284554.

Fu:2023:TAC

[FPA^{+}23] Xiuwen Fu, Pasquale Pace, Gianluca Aloi, Antonio Guerrieri, Wenfeng Li, and Giancarlo Fortino. Tolerance analysis of cybermanufacturing systems to cascading failures. ACM Transactions on Internet Technology (TOIT), 23(4):50:1–50:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579847.

Farias:2016:IDT

[FPR16] Delia Irazú Hernández Farías, Viviana Patti, and Paolo Rosso. Irony detection in Twitter: The role of affective content. ACM Transactions on Internet Technology (TOIT), 16(3): 19:1–19:??, August 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Fang:2004:LWM

[FS04] Xiao Fang and Olivia R. Liu Sheng. LinkSelector: A Web mining approach to hyperlink selection for Web portals. ACM Transactions on Internet Technology (TOIT), 4(2):209– 237, May 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Falcone:2015:RCT

[FSC15] Rino Falcone, Alessandro Sapienza, and Cristiano Castelfranchi. The relevance of categories for trusting information sources. ACM Transactions on Internet Technology (TOIT), 15(4): 13:1–13:??, December 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Fielding:2002:PDM

[FT02] Roy T. Fielding and Richard N. Taylor. Principled design of the modern Web architecture. ACM Transactions on Internet Technology (TOIT), 2(2):115– 150, May 2002. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Fang:2023:JAD

[FXYX23] Weiwei Fang, Wenyuan Xu, Chongchong Yu, and Neal. N. Xiong. Joint architecture design and workload partitioning for DNN inference on industrial IoT clusters. ACM Transactions on Internet Technology (TOIT), 23(1):7:1–7:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3551638.

Fujikawa:2017:SVN

[FYT17] Hiroshi Fujikawa, Hirofumi Yamaki, and Setsuo Tsuruta. Seamless virtual network for international business continuity in presence of intentional blocks. ACM Transactions on Internet Technology (TOIT), 18 (1):10:1–10:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Fan:2022:SEE

[FYW^{+}22] Zhenyu Fan, Wang Yang, Fan Wu, Jing Cao, and Weisong Shi. Serving at the edge: an edge computing service architecture based on ICN. ACM Transactions on Internet Technology (TOIT), 22(1):22:1–22:27, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3464428.

Feng:2019:PPP

[FYZ19] Jun Feng, Laurence T.

Yang, and Ronghao Zhang. Practical privacypreserving high-order bi-Lanczos in integrated edge-fog-cloud architecture for cyber-physical-social systems. ACM Transactions on Internet Technology (TOIT), 19(2):26:1–26:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3230641.

Feng:2021:BET

[FYZ^{+}21] Jun Feng, Laurence T. Yang, Yuxiang Zhu, Nicholaus J. Gati, and Yijun Mo. Blockchainenabled tensor-based conditional deep convolutional GAN for cyberphysical-social systems. ACM Transactions on Internet Technology (TOIT), 21(2):41:1–41:17, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3404890.

Graves:2018:SCC

[GAC18] James T. Graves, Alessandro Acquisti, and Nicolas Christin. Should credit card issuers reissue cards in response to a data breach?: Uncertainty and transparency in metrics for data security policymaking. ACM Transactions on Internet Technology (TOIT), 18 (4):54:1–54:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gavanelli:2018:APA

[GAL18] Marco Gavanelli, Marco Alberti, and Evelina Lamma. Accountable protocols in abductive logic programming. ACM Transactions on Internet Technology (TOIT), 18 (4):46:1–46:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gu:2022:AET

[GAL^{+}22] Bo Gu, Mamoun Alazab, Ziqi Lin, Xu Zhang, and Jun Huang. AIenabled task offloading for improving quality of computational experience in ultra dense networks. ACM Transactions on Internet Technology (TOIT), 22(3): 68:1–68:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3491217.

Garriga:2021:DCP

[GAT^{+}21] Martin Garriga, Koen Aarns, Christos Tsigkanos, Damian A. Tamburri, and Wjan Van Den Heuvel. DataOps for cyber-physical systems governance: The airport passenger flow case. ACM Transactions on Internet Technology (TOIT), 21(2):36:1–36:25, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432247.

Gupta:2008:SAI

[GBAR08] Manish Gupta, Shamik Banerjee, Manish Agrawal, and H. Raghav Rao. Security analysis of Internet technology components enabling globally distributed workplaces — a framework. ACM Transactions on Internet Technology (TOIT), 8(4):17:1–17:??, September 2008. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ge:2022:PDI

[GCK^{+}22] Mengmeng Ge, Jin-Hee Cho, Dongseong Kim, Gaurav Dixit, and Ing-Ray Chen. Proactive defense for Internet-of-things: Moving target defense with cyberdeception. ACM Transactions on Internet Technology (TOIT), 22(1):24:1– 24:31, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3467021.

Ge:2020:SBP

[GCP^{+}20] Xiaoyu Ge, Panos K. Chrysanthis, Konstantinos Pelechrinis, Demetrios Zeinalipour-Yazti, and Mohamed A. Sharaf. Serendipity-based pointsof-interest navigation. ACM Transactions on Internet Technology (TOIT), 20(4):33:1–33:32, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3391197.

Garcia-Dorado:2017:BMW

[GD17] José Luis García-Dorado.

Bandwidth measurements within the cloud: Characterizing regular behaviors and correlating downtimes. ACM Transactions on Internet Technology (TOIT), 17(4): 39:1–39:??, September 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

GarcIa-DIaz:2022:ISS

[GDLM22]  Vicente  GarcÍa-DÍaz,

Jerry Chun-Wei Lin, and Juan Antonio Morente Molinera. Introduction to the special section on edge computing AI-IoT integrated energy efficient intelligent transportation system for smart cities. ACM Transactions on Internet Technology (TOIT), 22 (4):105:1–105:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3584745.

Guerin:2014:ABS

[GdOW14] Roch Guérin, Jaudelice C. de Oliveira, and Steven Weber. Adoption of bundled services with network externalities and correlated affinities. ACM Transactions on Internet Technology (TOIT), 14 (2–3):13:1–13:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Glavic:2014:ESP

[GEFT14] Boris Glavic, Kyumars Sheykh Esmaili, Peter M. Fischer, and Nesime Tatbul. Efficient stream provenance via operator instrumentation. ACM Transactions on Internet Technology (TOIT), 14 (1):7:1–7:??, July 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gelenbe:2009:ASN

[Gel09] Erol Gelenbe. Analysis of single and networked auctions. ACM Transactions on Internet Technology (TOIT), 9(2):8:1– 8:??, May 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Golbeck:2006:IBT

[GH06] Jennifer Golbeck and James Hendler. Inferring binary trust relationships in Web-based social networks. ACM Transactions on Internet Technology (TOIT), 6(4):497– 529, November 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gao:2021:CEB

[GHD21] Honghao Gao, Wanqiu Huang, and Yucong Duan. The cloud-edge-based dynamic reconfiguration to service workflow for mobile ecommerce environments: a QoS prediction perspective. ACM Transactions on Internet Technology (TOIT), 21(1):6:1–6:23, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3391198.

Goncalves:2017:ESK

[GHK17] Jorge Goncalves, Simo Hosio, and Vassilis Kostakos. Eliciting structured knowledge from situated crowd markets. ACM Transactions on Internet Technology (TOIT), 17(2):

14:1–14:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gomathy:2021:ISC

[GJAT^{+}21] V. Gomathy, K. Janarthanan, Fadi Al-Turjman, R. Sitharthan, M. Rajesh, K. Vengatesan, and T. Priya Reshma. Investigating the spread of coronavirus disease via edge-AI and air pollution correlation. ACM Transactions on Internet Technology (TOIT), 21(4): 105:1–105:10, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3424222.

Gangwar:2023:CDS

[GK23] Arvind Kumar Gangwar and Sandeep Kumar. Concept drift in software defect prediction: a method for detecting and handling the drift. ACM Transactions on Internet Technology (TOIT), 23 (2):31:1–31:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3589342.

Geneves:2014:EIX

[GL14] Pierre Genevès and Nabil Layaïda. Equipping IDEs with XML-Path reasoning capabilities. ACM Transactions on Internet Technology (TOIT), 13 (4):13:1–13:??, July 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gordon:2002:LBD

[GLF02] Michael Gordon, Robert K.

Lindsay, and Weiguo Fan. Literature-based discovery on the World Wide Web. ACM Transactions on Internet Technology (TOIT), 2(4):261– 275, November 2002. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gan:2021:BFU

[GLFV^{+}21] Wensheng Gan, Jerry Chun-Wei Lin, Philippe Fournier-Viger, Han-Chieh Chao, and Philip S. Yu. Beyond frequency: Utility mining with varied item-specific minimum utility. ACM Transactions on Internet Technology (TOIT), 21(1): 3:1–3:32, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3425498.

Guo:2012:TNA

[GLJ^{+}12] Deke Guo, Yunhao Liu, Hai Jin, Zhong Liu, Weiming Zhang, and Hui Liu. Theory and network applications of balanced Kautz tree structures. ACM Transactions on Internet Technology (TOIT), 12(1):3:1–3:??, June 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Geneves:2011:IXS

[GLQ11] Pierre Genevès, Nabil Layaïda, and Vincent Quint. Impact of XML schema evolution. ACM Transactions on Internet Technology (TOIT), 11 (1):4:1–4:??, July 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gurevych:2017:ASM

[GLT17] Iryna Gurevych, Marco Lippi, and Paolo Torroni. Argumentation in social media. ACM Transactions on Internet Technology (TOIT), 17(3): 23:1–23:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gluhovsky:2010:FCT

[Glu10] Ilya Gluhovsky. Forecasting click-through rates based on sponsored search advertiser bids and intermediate variable regression. ACM Transactions on Internet Technology (TOIT), 10(3):11:1– 11:??, October 2010. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Guo:2017:IAC

[GLWH17] Yonghong Guo, Lu Liu, Yan Wu, and James Hardy. Interest-aware content discovery in peer-to-peer social networks. ACM Transactions on Internet Technology (TOIT), 18(3): 39:1–39:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gburzynski:2004:FSW

[GM04] Pawel Gburzynski and Jacek Maitan. Fighting the spam wars: a remailer approach with restrictive aliasing. ACM Transactions on Internet Technology (TOIT), 4 (1):1–30, February 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Groth:2009:MPD

[GMM09] Paul Groth, Simon Miles, and Luc Moreau. A model of process documentation to determine provenance in mashups. ACM Transactions on Internet Technology (TOIT), 9(1):3:1– 3:??, February 2009. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Goebel:2011:CIE

[GNK11] Christoph Goebel, Dirk Neumann, and Ramayya Krishnan. Comparing ingress and egress detection to secure interdomain routing: An experimental analysis. ACM Transactions on Internet Technology (TOIT), 11(2):5:1–5:??, December 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ghose:2019:GEI

[GNR19] Aditya Ghose, Hamid R. Motahari Nezhad, and Manfred Reichert. Guest Editors’ introduction to the special issue on knowledge-driven business process management. ACM Transactions on Internet Technology (TOIT), 19(1): 11:1–11:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gao:2020:RLS

[GNW^{+}20] Weichao Gao, James Nguyen, Yalong Wu, William G. Hatcher, and Wei Yu. Routing in largescale dynamic networks: a Bloom filter-based dual-layer scheme. ACM Transactions on Internet Technology (TOIT), 20(4):38:1–38:24, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3407192.

Gkatziaki:2018:EEU

[GPM^{+}18] Vasiliki Gkatziaki, Symeon Papadopoulos, Richard Mills, Sotiris Diplaris, Ioannis Tsampoulatidis, and Ioannis Kompatsiaris. easIE: Easyto-use information extraction for constructing CSR databases from the Web. ACM Transactions on Internet Technology (TOIT), 18(4): 45:1–45:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Goasdoue:2004:AQU

[GR04] François Goasdoué and Marie-Christine Rousset. Answering queries using views: A KRDB perspective for the semantic Web. ACM Transactions on Internet Technology (TOIT), 4(3):255– 288, August 2004. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ghanem:2020:EAF

[GRR20] Bilal Ghanem, Paolo Rosso, and Francisco Rangel. An emotional analysis of false information in social media and news articles. ACM Transactions on Internet Technology (TOIT), 20(2):19:1–19:18, May 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

                                        URL https://dl.acm.
                                        org/doi/abs/10.1145/
                                        3381750.

Gomes:2005:CNC

[GS05] Daniel Gomes and Mário J. Silva. Characterizing a national community Web. ACM Transactions on Internet Technology (TOIT), 5(3):508– 531, August 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gupta:2007:HKF

[GS07a] Amar Gupta and Satwik Seshasai. 24-hour knowledge factory: Using Internet technology to leverage spatial and temporal separations. ACM Transactions on Internet Technology (TOIT), 7(3): 14:1–14:??, August 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gupta:2007:GEI

[GS07b] Amar Gupta and Satwiksai Seshasai. Guest editorial: The Internet and outsourcing. ACM Transactions on Internet Technology (TOIT), 7(3): 13:1–13:??, August 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Guo:2017:PGE

[GS17] Tian Guo and Prashant Shenoy. Providing geo-elasticity in geographically distributed clouds. ACM Transactions on Internet Technology (TOIT), 18(3): 38:1–38:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Guo:2014:CAC

[GSS^{+}14] Tian Guo, Upendra Sharma, Prashant Shenoy, Timothy Wood, and Sambit Sahu. Cost-aware cloud bursting for enterprise applications. ACM Transactions on Internet Technology (TOIT), 13 (3):10:1–10:??, May 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gai:2023:BBA

[GSZ^{+}23] Keke Gai, Yufeng She, Liehuang Zhu, Kim-Kwang Raymond Choo, and Zhiguo Wan. A blockchain-based access control scheme for zero trust cross-organizational data sharing. ACM Transactions on Internet Technology (TOIT), 23(3):38:1–38:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3511899.

Guo:2002:OSS

[Guo02] Xin Guo. An optimal strategy for sellers in an online auction. ACM Transactions on Internet Technology (TOIT), 2 (1):1–13, February 2002. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gioacchini:2023:DIE

[GVM^{+}23] Luca Gioacchini, Luca Vassio, Marco Mellia, Idilio Drago, Zied Ben Houidi, and Dario Rossi. i-DarkVec: Incremental embeddings for Darknet traffic analysis. ACM Transactions on Internet Technology (TOIT), 23(3):45:1–45:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3595378.

Guan:2021:ASS

[GWF^{+}21] Zhitao Guan, Naiyu Wang, Xunfeng Fan, Xueyan Liu, Longfei Wu, and Shaohua Wan. Achieving secure search over encrypted data for ecommerce: a blockchain approach. ACM Transactions on Internet Technology (TOIT), 21(1): 12:1–12:17, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3408309.

Gao:2021:RFI

[GZL^{+}21] Guangwei Gao, Dong Zhu, Huimin Lu, Yi Yu, Heyou Chang, and Dong Yue. Robust facial image super-resolution by kernel locality-constrained coupled-layer regression. ACM Transactions on Internet Technology (TOIT), 21(3):67:1–67:15, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3418462.

Hsu:2022:ISS

[HAD22] Ching-Hsien Hsu, Amir H.

Alavi, and Mianxiong Dong. Introduction to the special section on cyber security in Internet of Vehicles. ACM Transactions on Internet Technology (TOIT), 22(4):81:1–81:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3584746.

Hoseiny:2021:JQA

[HAST21] Farooq Hoseiny, Sadoon Azizi, Mohammad Shojafar, and Rahim Tafazolli. Joint QoS-aware and cost-efficient task scheduling for fog-cloud resources in a volunteer computing system. ACM Transactions on Internet Technology (TOIT), 21 (4):86:1–86:21, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3418501.

Huck:2002:SCS

[HBGF02] Paul Huck, Michael Butler, Amar Gupta, and Michael Feng. A selfconfiguring and selfadministering name system with dynamic address assignment. ACM Transactions on Internet Technology (TOIT), 2 (1):14–46, February 2002. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Hasan:2014:ASM

[HC14] Souleiman Hasan and Edward Curry. Approximate semantic matching of events for the Internet of Things. ACM Transactions on Internet Technology (TOIT), 14 (1):2:1–2:??, July 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Hernandez-Castro:2023:BCU

[HCBRM23] Carlos Hernández-Castro, David F. Barrero, and Maria Dolores R-Moreno. Breaking CaptchaStar using the BASECASS methodology. ACM Transactions on Internet Technology (TOIT), 23(1):5:1–5:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546867.

Hao:2021:AMT

[HCW^{+}21] Shijie Hao, Tao Chen, Yang Wang, Yanrong Guo, Meng Wang, and For the Alzheimer’s Disease Neuroimaging Initiative. Adaptive multi-task dual-structured learning with its application on Alzheimer’s disease study. ACM Transactions on Internet Technology (TOIT), 21(2): 47:1–47:16, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3398728.

Huang:2007:DDA

[HGW07] Yun Huang, Xianjun Geng, and Andrew B. Whinston. Defeating DDoS attacks by fixing the incentive chain. ACM Transactions on Internet Technology (TOIT), 7(1): 5:1–5:??, February 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Hourany:2021:PNS

[HHF^{+}21] Edy Hourany, Bachir Habib, Camille Fountaine, Abdallah Makhoul, Benoit Piranda, and Julien Bourgeois. PRO-LISEAN: a new security protocol for programmable matter. ACM Transactions on Internet Technology (TOIT), 21(1):22:1–22:29, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3432250.

Hossain:2022:EFS

[HHS^{+}22] Md Arafat Hossain, Jun Han, Jean-Guy Schneider, Jiaojiao Jiang, Muhammad Ashad Kabir, and Steve Versteeg. Extracting formats of service messages with varying payloads. ACM Transactions on Internet Technology (TOIT), 22(3):71:1–71:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3503159.

He:2003:SAA

[HJ03] Minghua He and Nicholas R.

Jennings. Southampton-TAC: an adaptive autonomous trading agent. ACM Transactions on Internet Technology (TOIT), 3(3):218–235, August 2003. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Herzberg:2008:SII

[HJ08] Amir Herzberg and Ahmad Jbara. Security and identification indicators for browsers against spoofing and phishing attacks. ACM Transactions on Internet Technology (TOIT), 8(4):16:1–16:??, September 2008. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

He:2006:HBS

[HJPB06] Minghua He, Nicholas R.

Jennings, and Adam Prügel-Bennett. A heuristic bidding strategy for buying multiple goods in multiple English auctions. ACM Transactions on Internet Technology (TOIT), 6(4):465– 496, November 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Huang:2020:PRR

[HJWW20] Chunli Huang, Wenjun Jiang, Jie Wu, and Guojun Wang. Personalized review recommendation based on users’ aspect sentiment. ACM Transactions on Internet Technology (TOIT), 20(4):42:1–42:26, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3414841.

Hoefer:2014:AAS

[HKV14] Martin Hoefer, Thomas Kesselheim, and Berthold Vöcking. Approximation algorithms for secondary spectrum auctions. ACM Transactions on Internet Technology (TOIT), 14(2–3):16:1–16:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Huang:2021:DAM

[HLG^{+}21] Feiran Huang, Chaozhuo Li, Boyu Gao, Yun Liu, Sattam Alotaibi, and Hao Chen. Deep attentive multimodal network representation learning for social media images. ACM Transactions on Internet Technology (TOIT), 21(3): 69:1–69:17, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3417295.

Hu:2021:DDR

[HLLS21] Kaixi Hu, Lin Li, Jianquan Liu, and Daniel Sun. DuroNet: a dual-robust enhanced spatial-temporal learning network for urban crime prediction. ACM Transactions on Internet Technology (TOIT), 21(1):24:1–24:24, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3432249.

Hu:2021:MBT

[HML^{+}21] He-Xuan Hu, Wen-Jie Mao, Zhen-Zhou Lin, Qiang Hu, and Ye Zhang. Multimodal brain tumor segmentation based on an intelligent UNET–LSTM algorithm in smart hospitals. ACM Transactions on Internet Technology (TOIT), 21(3): 74:1–74:14, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3450519.

Huang:2021:DSD

[HMLH21] Liwei Huang, Yutao Ma, Yanbo Liu, and Keqing He. DAN-SNR: a deep attentive network for social-aware next pointof-interest recommendation. ACM Transactions on Internet Technology (TOIT), 21(1):2:1–2:27, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3430504.

Harumoto:2005:EWB

[HNF^{+}05] Kaname Harumoto, Tadashi Nakano, Shinya Fukumura, Shinji Shimojo, and Shojiro Nishio. Effective Web browsing through content delivery adaptation. ACM Transactions on Internet Technology (TOIT), 5(4):571– 600, November 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Hirsch:2023:TBE

[HNGN23] Sharon Hirsch, Slava Novgorodov, Ido Guy, and Alexander Nus. The tip of the buyer: Extracting product tips from reviews. ACM Transactions on Internet Technology (TOIT), 23(1): 4:1–4:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3547140.

Hochheiser:2002:PPP

[Hoc02] Harry Hochheiser. The platform for privacy preference as a social protocol: an examination within the U.S. policy context. ACM Transactions on Internet Technology (TOIT), 2(4):276– 306, November 2002. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Hosoya:2003:XST

[HP03] Haruo Hosoya and Benjamin C. Pierce. XDuce: a statically typed XML processing language. ACM Transactions on Internet Technology (TOIT), 3(2): 117–148, May 2003. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Hafizoglu:2019:UIP

[HS19] Feyza Merve Hafizoglu and Sandip Sen. Understanding the influences of past experience on trust in humanagent teamwork. ACM Transactions on Internet Technology (TOIT), 19(4):45:1–45:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3324300.

Hu:2017:MDS

[HSLH17] Yan Hu, Weisong Shi, Hong Li, and Xiaohui Hu. Mitigating data sparsity using similarity reinforcementenhanced collaborative filtering. ACM Transactions on Internet Technology (TOIT), 17(3): 31:1–31:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Hao:2023:RSM

[HSRK23] Jianwei Hao, Piyush Subedi, Lakshmish Ramaswamy, and In Kee Kim. Reaching for the sky: Maximizing deep learning inference throughput on edge devices with AI multitenancy. ACM Transactions on Internet Technology (TOIT), 23(1): 2:1–2:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546192.

Hui:2006:OID

[HTG06] Kai-Lung Hui, Bernard C. Y. Tan, and Chyan-Yee Goh. Online information disclosure: Motivators and measurements. ACM Transactions on Internet Technology (TOIT), 6(4):415– 441, November 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Hossain:2022:SSE

[HXB^{+}22] M. Shamim Hossain, Changsheng Xu, Josu Bilbao, Md. Abdur Rahman, Abdulmotaleb El Saddik, and Mohamed Bin Zayed. Special section on edge-AI for connected living. ACM Transactions on Internet Technology (TOIT), 22(3):55:1–55:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3514196.

Hu:2020:SSS

[HXZ^{+}20] Zhengdi Hu, Guangquan Xu, Xi Zheng, Jiang Liu, Zhangbing Li, Quan Z. Sheng, Wenjuan Lian, and Hequn Xian. SSL-SVD: Semi-supervised learning-based sparse trust recommendation. ACM Transactions on Internet Technology (TOIT), 20(1):4:1–4:20, March 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/doi/abs/10.1145/
3369390.

Hurley:2011:NDT

[HZ11] Neil Hurley and Mi Zhang. Novelty and diversity in top-N recommendation — analysis and evaluation. ACM Transactions on Internet Technology (TOIT), 10(4): 14:1–14:??, March 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Huang:2019:IEF

[HZB19] Teng-Chieh Huang, Razieh Nokhbeh Zaeem, and K. Suzanne Barber. It is an equal failing to trust everybody and to trust nobody: Stock price prediction using trust filters and enhanced user sentiment on Twitter. ACM Transactions on Internet Technology (TOIT), 19(4):48:1–48:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3338855.

Huang:2010:PNA

[HZCS10] Tzu-Chi Huang, Sherali Zeadally, Naveen Chilamkurti, and Ce-Kuen Shieh. A programmable network address translator: Design, implementation, and performance. ACM Transactions on Internet Technology (TOIT), 10(1): 3:1–3:??, February 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

He:2012:SWS

[HZHC12] Jing He, Yanchun Zhang, Guangyan Huang, and Jinli Cao. A smart Web service based on the context of things. ACM Transactions on Internet Technology (TOIT), 11 (3):13:1–13:??, January 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ignat:2019:ITS

[IDS19] Claudia-Lavinia Ignat, Quang-Vinh Dang, and Valerie L. Shalin. The influence of trust score on cooperative behavior. ACM Transactions on Internet Technology (TOIT), 19(4): 46:1–46:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3329250.

Iwendi:2021:SSI

[IRJ^{+}21] Celestine Iwendi, Saif Ur Rehman, Abdul Rehman Javed, Suleman Khan, and Gautam Srivastava. Sustainable security for the Internet of Things using artificial intelligence architectures. ACM Transactions on Internet Technology (TOIT), 21(3): 73:1–73:22, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3448614.

Ivkic:2022:SCM

[ISG^{+}22] Igor Ivkić, Patrizia Sailer, Antonios Gouglidis, Andreas Mauthe, and Markus Tauber. A security cost modelling framework for cyber-physical systems. ACM Transactions on Internet Technology (TOIT), 22(2): 53:1–53:31, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3450752.

Jonsson:2006:POS

[JAT^{+}06] Björn Thør Jónsson, María Arinbjarnar, Bjarnsteinn Thórsson, Michael J.

Franklin, and Divesh Srivastava. Performance and overhead of semantic cache management. ACM Transactions on Internet Technology (TOIT), 6(3): 302–331, August 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jiang:2018:FCC

[JCH^{+}18] He Jiang, Xin Chen, Tieke He, Zhenyu Chen, and Xiaochen Li. Fuzzy clustering of crowdsourced test reports for apps. ACM Transactions on Internet Technology (TOIT), 18(2): 18:1–18:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jin:2021:FSL

[JDZ^{+}21] Xin Jin, Yuwei Duan, Ying Zhang, Yating Huang, Mengdong Li, Ming Mao, Amit Kumar Singh, and Yujie Li. Fast search of lightweight block cipher primitives via swarm-like metaheuristics for cyber security. ACM Transactions on Internet Technology (TOIT), 21(4): 93:1–93:15, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3417296.

Jordan:2010:FCT

[JG10] Scott Jordan and Arijit Ghosh. A framework for classification of traffic management practices as reasonable or unreasonable. ACM Transactions on Internet Technology (TOIT), 10 (3):12:1–12:??, October 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jiang:2022:FTC

[JGH^{+}22] Yizhang Jiang, Xiaoqing Gu, Lei Hua, Kang Li, Yuwen Tao, and Bo Li. Forecasting trend of coronavirus disease 2019 using multi-task weighted TSK fuzzy system. ACM Transactions on Internet Technology (TOIT), 22(3): 64:1–64:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3475870.

Jiang:2022:SDM

[JHC^{+}22] Nan Jiang, Debin Huang, Jing Chen, Jie Wen, Heng Zhang, and Honglong Chen. Semi-direct monocular visual-inertial odometry using point and line features for IoV. ACM Transactions on Internet Technology (TOIT), 22(1):5:1–5:23, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3432248.

Ji:2002:ABM

[Ji02]  Minwen Ji.  Affinity-

based management of main memory database clusters. ACM Transactions on Internet Technology (TOIT), 2(4):307– 339, November 2002. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jeong:2021:MPP

[JKI^{+}21] Junho Jeong, Donghyo Kim, Sun-Young Ihm, Yangsun Lee, and Yunsik Son. Multilateral personal portfolio authentication system based on hyperledger fabric. ACM Transactions on Internet Technology (TOIT), 21(1):14:1–14:17, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3423554.

Jacob:2007:ICF

[JKR07] Varghese S. Jacob, Ramayya Krishnan, and Young U. Ryu. Internet content filtering using isotonic separation on content category ratings. ACM Transactions on Internet Technology (TOIT), 7(1):1:1– 1:??, February 2007. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jha:2010:SIL

[JKS^{+}10] Somesh Jha, Stefan Katzenbeisser, Christian Schallhart, Helmut Veith, and Stephen Chenney. Semantic integrity in large-scale online simulations. ACM Transactions on Internet Technology (TOIT), 10(1):2:1–2:??, February 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jiang:2020:EAS

[JLC20] Xingbin Jiang, Michele Lora, and Sudipta Chattopadhyay. An experimental analysis of security vulnerabilities in industrial IoT devices. ACM Transactions on Internet Technology (TOIT), 20(2): 16:1–16:24, May 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1145/3379542.

Jansen:2006:AGW

[JMSP06] Bernard J. Jansen, Tracy Mullen, Amanda Spink, and Jan Pedersen. Automated gathering of Web information: an in-depth examination of agents interacting with search engines. ACM Transactions on Internet Technology (TOIT), 6(4):442– 464, November 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jin:2008:QAS

[JN08] Jingwen Jin and Klara Nahrstedt. QoS-aware service management for component-based distributed applications. ACM Transactions on Internet Technology (TOIT), 8(3):14:1–14:??, May 2008. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jordan:2009:IIA

[Jor09] Scott Jordan. Implications of Internet architecture on net neutrality. ACM Transactions on Internet Technology (TOIT), 9(2):5:1– 5:??, May 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jang:2022:CCA

[JPCL22] Si Young Jang, Sung Kyu Park, Jin Hee Cho, and Dongman Lee. CARES: Context-aware trust estimation for realtime crowdsensing services in vehicular edge networks. ACM Transactions on Internet Technology (TOIT), 22(4): 92:1–92:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3514243.

Jajodia:2017:SSH

[JPSS17] Sushil Jajodia, Noseong Park, Edoardo Serra, and V. S. Subrahmanian. SHARE: a Stackelberg honey-based adversarial reasoning engine. ACM Transactions on Internet Technology (TOIT), 18 (3):30:1–30:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jordan:2013:UIR

[JS13] Scott Jordan and Gwen Shaffer. User and ISP rights of device attachment and device management. ACM Transactions on Internet Technology (TOIT), 13(2):6:1–6:??, December 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Junaid:2022:ASV

[JSAA22] Muhammad Junaid, Adnan Sohail, Fadi Al Turjman, and Rashid Ali. Agile support vector machine for energy-efficient resource allocation in IoT-oriented cloud using PSO. ACM Transactions on Internet Technology (TOIT), 22(1):6:1–6:35, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3433541.

Jiang:2015:SRT

[JWW15] Wenjun Jiang, Jie Wu, and Guojun Wang. On selecting recommenders for trust evaluation in online social networks. ACM Transactions on Internet Technology (TOIT), 15(4):14:1–14:??, December 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Jiang:2019:CNB

[JYW^{+}19] Zexun Jiang, Hao Yin, Yulei Wu, Yongqiang Lyu, Geyong Min, and Xu Zhang. Constructing novel block layouts for webpage analysis. ACM Transactions on Internet Technology (TOIT), 19(3):35:1–35:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3326457.

Kolomvatsos:2020:IEC

[KA20] Kostas Kolomvatsos and Christos Anagnostopoulos. An intelligent edgecentric queries allocation scheme based on ensemble models. ACM Transactions on Internet Technology (TOIT), 20(4):45:1–45:25, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3417297.

Kavurmacioglu:2014:DIP

[KAS14] Emir Kavurmacioglu, Murat Alanyali, and David Starobinski. Demandinvariant price relationships and market outcomes in competitive private commons. ACM Transactions on Internet Technology (TOIT), 14(2–3):15:1–15:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kourtellis:2015:SIF

[KBBI15] Nicolas Kourtellis, Jeremy Blackburn, Cristian Borcea, and Adriana Iamnitchi.

Special issue on foundations of social computing: Enabling social applications via decentralized social data management. ACM Transactions on Internet Technology (TOIT), 15(1):1:1–1:??, February 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kayal:2018:ARN

[KBNV18] Alex Kayal, Willem-Paul Brinkman, Mark A. Neerincx, and M. Birna Van Riemsdijk. Automatic resolution of normative conflicts in supportive technology based on user values. ACM Transactions on Internet Technology (TOIT), 18 (4):41:1–41:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kafali:2017:GEI

[KCR^{+}17] Özgür Kafali, Natalia Criado, Martin Rehak, Jose M. Such, and Pinar Yolum. Guest Editors’ introduction. ACM Transactions on Internet Technology (TOIT), 18(3): 26:1–26:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Karame:2014:MMW

[KFB^{+}14] Ghassan O. Karame, Aurélien Francillon, Victor Budilivschi, Srdjan Capkun, and Vedran Capkun. Microcomputations as micropayments in Web-based services. ACM Transactions on Internet Technology (TOIT), 13(3):8:1–8:??, May 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kuter:2010:UPC

[KG10] Ugur Kuter and Jennifer Golbeck. Using probabilistic confidence models for trust inference in Web-based social networks. ACM Transactions on Internet Technology (TOIT), 10(2): 8:1–8:??, May 2010. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kumar:2022:IIF

[KGAR22] Rahul Kumar, Ankur Gupta, Harkirat Singh Arora, and Balasubramanian Raman. IBRDM: an intelligent framework for brain tumor classification using radiomicsand DWT-based fusion of MRI sequences. ACM Transactions on Internet Technology (TOIT), 22(1):9:1–9:30, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3434775.

Kaur:2021:ESD

[KGKK21] Kuljeet Kaur, Sahil Garg, Georges Kaddoum, and Neeraj Kumar. Energy and SLA-driven MapReduce job scheduling framework for cloudbased cyber-physical systems. ACM Transactions on Internet Technology (TOIT), 21(2): 31:1–31:24, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3409772.

Konstantinidis:2019:IDP

[KIG^{+}19] Andreas Konstantinidis, Panagiotis Irakleous, Zacharias Georgiou, Demetrios Zeinalipour-Yazti, and Panos K. Chrysanthis. IoT data prefetching in indoor navigation SOAs. ACM Transactions on Internet Technology (TOIT), 19(1):10:1–10:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kim:2021:IMB

[KK21] Junho Kim and Mucheol Kim. Intelligent mediatorbased enhanced smart contract for privacy protection. ACM Transactions on Internet Technology (TOIT), 21(1): 8:1–8:16, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3404892.

Kim:2021:DIB

[KKK21] Tae-Yeun Kim, Sung-Hwan Kim, and Hoon Ko. Design and implementation of BCI-based intelligent upper limb rehabilitation robot system. ACM Transactions on Internet Technology (TOIT), 21(3): 60:1–60:17, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392115.

Ko:2016:STU

[KKMK16] In-Young Ko, Han-Gyu Ko, Angel Jimenez Molina, and Jung-Hyun Kwon. SoIoT: Toward a usercentric IoT-based service framework. ACM Transactions on Internet Technology (TOIT), 16(2): 8:1–8:??, April 2016. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kekulluoglu:2018:PPS

[KKY18] Dilara Kekulluoglu, Nadin Kokciyan, and Pinar Yolum. Preserving privacy as social responsibility in online social networks. ACM Transactions on Internet Technology (TOIT), 18(4): 42:1–42:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Knutsson:2003:APS

[KLMH03] Björn Knutsson, Honghui Lu, Jeffrey Mogul, and Bryan Hopkins. Architecture and performance of server-directed transcoding. ACM Transactions on Internet Technology (TOIT), 3(4):392– 424, November 2003. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kim:2017:MDS

[KLS^{+}17] Taehun Kim, Junsung Lim, Heesuk Son, Byoungheon Shin, Dongman Lee, and Soon J. Hyun. A multi-dimensional smart community discovery scheme for IoTenriched smart homes. ACM Transactions on Internet Technology (TOIT), 18(1):3:1–3:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kampik:2022:GAA

[KMB^{+}22] Timotheus Kampik, Adnane Mansour, Olivier Boissier, Sabrina Kirrane, Julian Padget, Terry R. Payne, Munindar P. Singh, Valentina Tamma, and Antoine Zimmermann. Governance of autonomous agents on the Web: Challenges and opportunities. ACM Transactions on Internet Technology (TOIT), 22(4):104:1– 104:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3507910.

Kenny:2009:CES

[KMW09] Alan Kenny, Séamus Mcloone, and Tomás Ward. Controlling entity state updates to maintain remote consistency within a distributed interactive application. ACM Transactions on Internet Technology (TOIT), 9(4): 15:1–15:??, September 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kristol:2001:HCS

[Kri01] David M. Kristol. HTTP Cookies: Standards, privacy, and politics. ACM Transactions on Internet Technology (TOIT), 1(2): 151–198, November 2001. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kwon:2009:FXD

[KRML09] Joonho Kwon, Praveen Rao, Bongki Moon, and Sukho Lee. Fast XML document filtering by sequencing twig patterns. ACM Transactions on Internet Technology (TOIT), 9(4):13:1–13:??, September 2009. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kumar:2006:CAC

[KRRT06] Ravi Kumar, Prabhakar Raghavan, Sridhar Rajagopalan, and Andrew Tomkins. Core algorithms in the CLEVER system. ACM Transactions on Internet Technology (TOIT), 6(2):131– 152, May 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kobsa:2003:PTP

[KS03] Alfred Kobsa and Jörg Schreck. Privacy through pseudonymity in useradaptive systems. ACM Transactions on Internet Technology (TOIT), 3(2): 149–183, May 2003. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Keromytis:2007:RSA

[KS07] Angelos D. Keromytis and Jonathan M. Smith. Requirements for scalable access control and security management architectures. ACM Transactions on Internet Technology (TOIT), 7(2):8:1– 8:??, May 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kumaraguru:2010:TJF

[KSA^{+}10] Ponnurangam Kumaraguru, Steve Sheng, Alessandro Acquisti, Lorrie Faith Cranor, and Jason Hong. Teaching Johnny not to fall for phish. ACM Transactions on Internet Technology (TOIT), 10 (2):7:1–7:??, May 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Krol:2021:PPU

[KSAB^{+}21] Michał Król, Alberto Sonnino, Mustafa Al-Bassam, Argyrios G. Tasiopoulos, Etienne Rivière, and Ioannis Psaras. Proof-of-prestige: a useful work reward system for unverifiable tasks. ACM Transactions on Internet Technology (TOIT), 21(2):

44:1–44:27, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3419483.

Kumar:2021:NCA

[KSL^{+}21] Abhinav Kumar, Sanjay Kumar Singh, K. Lakshmanan, Sonal Saxena, and Sameer Shrivastava. A novel cloud-assisted secure deep feature classification framework for cancer histopathology images. ACM Transactions on Internet Technology (TOIT), 21(2): 52:1–52:22, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3424221.

Kokciyan:2017:AAR

[KYY17] Nadin Kökciyan, Nefise Yaglikci, and Pinar Yolum. An argumentation approach for resolving privacy disputes in online social networks. ACM Transactions on Internet Technology (TOIT), 17(3): 27:1–27:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Kuang:2021:ITS

[KZLG21] Li Kuang, Jianbo Zheng, Kemu Li, and Honghao Gao. Intelligent traffic signal control based on reinforcement learning with state reduction for smart cities. ACM Transactions on Internet Technology (TOIT), 21(4): 102:1–102:24, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3418682.

Lyback:2004:ATS

[LB04] David Lybäck and Magnus Boman. Agent trade servers in financial exchange systems. ACM Transactions on Internet Technology (TOIT), 4(3): 329–339, August 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Li:2024:MMC

[LBC^{+}24] Wen Li, Lingfeng Bao, Jiachi Chen, John Grundy, Xin Xia, and Xiaohu Yang. Market manipulation of cryptocurrencies: Evidence from social media and transaction data. ACM Transactions on Internet Technology (TOIT), 24(2): 8:1–8:??, May 2024. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3643812.

Liu:2012:EMM

[LC12] Ziyang Liu and Yi Chen.

Exploiting and maintaining materialized views for XML keyword queries. ACM Transactions on Internet Technology (TOIT), 12(2):6:1–6:??, December 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Leitner:2016:PCS

[LC16] Philipp Leitner and Jürgen Cito. Patterns in the chaos — a study of performance variation and predictability in public IaaS clouds. ACM Transactions on Internet Technology (TOIT), 16(3):15:1–15:??, August 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Li:2005:CSM

[LCKN05] Mingzhe Li, Mark Claypool, Robert Kinicki, and James Nichols. Characteristics of streaming media stored on the Web. ACM Transactions on Internet Technology (TOIT), 5(4):601– 626, November 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lv:2021:BDP

[LCS21] Zhihan Lv, Dongliang Chen, and Amit Kumar Singh. Big data processing on volunteer computing. ACM Transactions on Internet Technology (TOIT), 21(4): 83:1–83:20, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3409801.

Li:2023:HCV

[LDG^{+}23] Zhenyu Li, Yong Ding, Honghao Gao, Bo Qu, Yujue Wang, and Jun Li. A highly compatible verification framework with minimal upgrades to secure an existing edge network. ACM Transactions on Internet Technology (TOIT), 23(3): 41:1–41:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3511901.

Lopez:2017:BMC

[LFL17] Claudia López, Rosta Farzan, and Yu-Ru Lin. Behind the myths of citizen participation: Identifying sustainability factors of hyper-local information systems. ACM Transactions on Internet Technology (TOIT), 18 (1):11:1–11:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Loukides:2020:OAI

[LGC20] Grigorios Loukides, Robert Gwadera, and Shing-Wan Chang. Overexposureaware influence maximization. ACM Transactions on Internet Technology (TOIT), 20(4): 39:1–39:31, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3408315.

Loukil:2021:DPB

[LGGB^{+}21] Faiza Loukil, Chirine Ghedira-Guegan, Khouloud Boukadi, Aïcha-Nabila Benharkat, and Elhadj Benkhelifa. Data privacy based on IoT device behavior control using blockchain. ACM Transactions on Internet Technology (TOIT), 21(1):23:1–23:20, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3434776.

Luckner:2020:IAU

[LGKL20] Marcin Luckner, Maciej Grzenda, Robert Kunicki, and Jaroslaw Legierski. IoT architecture for urban data-centric services and applications. ACM Transactions on Internet Technology (TOIT), 20(3):29:1–29:30, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3396850.

Le:2022:SIR

[LHAT22] Tu Le, Danny Yuxing Huang, Noah Apthorpe, and Yuan Tian. Skill-Bot: Identifying risky content for children in Alexa skills. ACM Transactions on Internet Technology (TOIT), 22(3): 79:1–79:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3539609.

Lin:2022:EEC

[LHL^{+}22] Weiwei Lin, Tiansheng Huang, Xin Li, Fang Shi, Xiumin Wang, and Ching-Hsien Hsu. Energyefficient computation offloading for UAV-assisted MEC: a two-stage optimization scheme. ACM Transactions on Internet Technology (TOIT), 22(1):4:1–4:23, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3430503.

Lin:2006:ISP

[LHTL06] Jeng-Wei Lin, Jan-Ming Ho, Li-Ming Tseng, and Feipei Lai. IDN server proxy architecture for Internationalized Domain Name resolution and experiences with providing Web services. ACM Transactions on Internet Technology (TOIT), 6 (1):1–19, February 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lin:2021:BBD

[LHZ^{+}21] Chao Lin, Debiao He, Sherali Zeadally, Xinyi Huang, and Zhe Liu. Blockchain-based data sharing system for sensingas-a-service in smart cities. ACM Transactions on Internet Technology (TOIT), 21(2): 40:1–40:21, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397202.

Liu:2012:FPC

[Liu12] Alex X. Liu. Firewall policy change-impact analysis. ACM Transactions on Internet Technology (TOIT), 11(4): 15:1–15:??, March 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Liu:2020:ITO

[Liu20] Ling Liu. Internet technology outlook: From communication to storage and cognitive computing. ACM Transactions on Internet Technology (TOIT), 20(1): 1:1–1:4, March 2020. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/abs/
                                        10.1145/3378661.

Laszka:2018:ASR

[LJG18] Aron Laszka, Benjamin Johnson, and Jens Grossklags. On the assessment of systematic risk in networked systems. ACM Transactions on Internet Technology (TOIT), 18 (4):48:1–48:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Leung:2016:CMS

[LJLN16] Kenneth Wai-Ting Leung, Di Jiang, Dik Lun Lee, and Wilfred Ng. Constructing maintainable semantic relation network from ambiguous concepts in Web content. ACM Transactions on Internet Technology (TOIT), 16(1):6:1–6:??, February 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Laszka:2014:STC

[LJS^{+}14] Aron Laszka, Benjamin Johnson, Pascal Schöttle, Jens Grossklags, and Rainer Böhme. Secure team composition to thwart insider threats and cyber-espionage. ACM Transactions on Internet Technology (TOIT), 14 (2–3):19:1–19:??, October 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lin:2023:VMI

[LLC^{+}23] Yi-Bing Lin, Yuan-Fu Liao, Sin-Horng Chen, Shaw-Hwa Hwang, and Yih-Ru Wang. VoiceTalk: Multimedia-IoT applications for mixing Mandarin, Taiwanese, and English. ACM Transactions on Internet Technology (TOIT), 23(2): 28:1–28:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3543854.

Liang:2022:PPP

[LLG22] Yangfan Liang, Yining Liu, and Brij B. Gupta. PPRP: Preserving-privacy route planning scheme in VANETs. ACM Transactions on Internet Technology (TOIT), 22(4):85:1–85:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3430507.

Lv:2022:ECS

[LLL22] Zhihan Lv, Ranran Lou, and Haibin Lv. Edge computing to solve security issues for infectious disease intelligence prevention. ACM Transactions on Internet Technology (TOIT), 22(3): 63:1–63:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3475869.

Leung:2012:FPW

[LLNF12] Kenneth Wai-Ting Leung, Dik Lun Lee, Wilfred Ng, and Hing Yuet Fung. A framework for personalizing web search with concept-based user profiles. ACM Transactions on Internet Technology (TOIT), 11(4): 17:1–17:??, March 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Li:2008:TSD

[LLSL08] Qing Li, Rynson W. H. Lau, Timothy K. Shih, and Frederick W. B. Li. Technology supports for distributed and collaborative learning over the Internet. ACM Transactions on Internet Technology (TOIT), 8(2):5:1– 5:??, February 2008. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Li:2008:ISI

[LLSM08] Qing Li, Rynson W. H.

Lau, Timothy Shih, and Dennis McLeod. Introduction to special issue Internet technologies for distance education. ACM Transactions on Internet Technology (TOIT), 8(2): 1:1–1:??, February 2008. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lv:2022:TLP

[LLSW22] Zhihan Lv, Ranran Lou, Amit Kumar Singh, and Qingjun Wang. Transfer learning-powered resource optimization for green computing in 5Gaided industrial Internet of Things. ACM Transactions on Internet Technology (TOIT), 22(2): 38:1–38:16, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3434774.

Lempel:2004:ORP

[LM04] Ronny Lempel and Shlomo Moran. Optimizing result prefetching in Web search engines with segmented indices. ACM Transactions on Internet Technology (TOIT), 4(1):31– 59, February 2004. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Li:2021:DSA

[LMS^{+}21] Qianmu Li, Shunmei Meng, Xiaonan Sang, Hanrui Zhang, Shoujin Wang, Ali Kashif Bashir, Keping Yu, and Usman Tariq. Dynamic scheduling algorithm in cyber mimic defense architecture of volunteer computing. ACM Transactions on Internet Technology (TOIT), 21(3): 75:1–75:33, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3408291.

LaMorgia:2023:DWS

[LMSS23] Massimo La Morgia, Alessandro Mei, Francesco Sassi, and Julinda Stefa. The doge of Wall Street: Analysis and detection of pump and dump cryptocurrency manipulations. ACM Transactions on Internet Technology (TOIT), 23(1):11:1– 11:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3561300.

Lukasiewicz:2014:OBQ

[LMSTM14]  Thomas  Lukasiewicz,

Maria Vanina Martinez, Gerardo I. Simari, and Oana Tifrea-Marciuska. Ontology-based query answering with group preferences. ACM Transactions on Internet Technology (TOIT), 14(4): 25:1–25:??, December 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Li:2013:CBA

[LMZ13] Xitong Li, Stuart E.

Madnick, and Hongwei Zhu. A contextbased approach to reconciling data interpretation conflicts in Web services composition. ACM Transactions on Internet Technology (TOIT), 13(1):1:1–1:??, November 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Li:2023:BBZ

[LNTL23] Shancang Li, Surya Nepal, Theo Tryfonas, and Hongwei Li. Blockchainbased zero trust cybersecurity in the Internet of Things. ACM Transactions on Internet Technology (TOIT), 23(3): 36:1–36:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3594535.

Li:2019:DRS

[LOD19] He Li, Kaoru Ota, and Mianxiong Dong. Deep reinforcement scheduling for mobile crowdsensing in fog computing. ACM Transactions on Internet Technology (TOIT), 19(2):21:1–21:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3234463.

Lv:2021:SMD

[LP21] Zhihan Lv and Francesco Piccialli. The security of medical data on Internet based on differential privacy technology. ACM Transactions on Internet Technology (TOIT), 21(3): 55:1–55:18, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3382769.

Lawrence:2017:UAS

[LPB^{+}17] John Lawrence, Joonsuk Park, Katarzyna Budzynska, Claire Cardie, Barbara Konat, and Chris Reed. Using argumentative structure to interpret debates in online deliberative democracy and eRulemaking. ACM Transactions on Internet Technology (TOIT), 17 (3):25:1–25:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Longo:2019:GEI

[LPR19] Francesco Longo, Antonio Puliafito, and Omer Rana. Guest Editors’ introduction to the special issue on fog, edge, and cloud integration for smart environments. ACM Transactions on Internet Technology (TOIT), 19(2):17:1–17:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3319404.

Li:2021:HFS

[LPX^{+}21] Haolun Li, Chi-Man Pun, Feng Xu, Longsheng Pan, Rui Zong, Hao Gao, and Huimin Lu. A hybrid feature selection algorithm based on a discrete artificial bee colony for Parkinson’s diagnosis. ACM Transactions on Internet Technology (TOIT), 21(3):

63:1–63:22, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3397161.

Lv:2021:AEIa

[LQSW21] Zhihan Lv, Liang Qiao, Amit Kumar Singh, and Qingjun Wang. AIempowered IoT security for smart cities. ACM Transactions on Internet Technology (TOIT), 21 (4):99:1–99:21, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3406115.

Lv:2021:AEIb

[LQVK21] Zhihan Lv, Liang Qiao, Sahil Verma, and Kavita. AI-enabled IoT-edge data analytics for connected living. ACM Transactions on Internet Technology (TOIT), 21(4): 104:1–104:20, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3421510.

Lv:2021:CRN

[LQW21] Zhihan Lv, Liang Qiao, and Qingjun Wang. Cognitive robotics on 5G networks. ACM Transactions on Internet Technology (TOIT), 21(4):

92:1–92:18, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3414842.

Lv:2021:BDA

[LS21] Zhihan Lv and Amit Kumar Singh. Big data analysis of Internet of Things system. ACM Transactions on Internet Technology (TOIT), 21(2): 28:1–28:15, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3389250.

Li:2005:OMC

[LSCZ05] Keqiu Li, Hong Shen, Francis Y. L. Chin, and Si Qing Zheng. Optimal methods for coordinated enroute Web caching for tree networks. ACM Transactions on Internet Technology (TOIT), 5(3): 480–507, August 2005. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lawrence:2017:DTD

[LSK^{+}17a] John Lawrence, Mark Snaith, Barbara Konat, Katarzyna Budzynska, and Chris Reed. Debating technology for dialogical argument: Sensemaking, engagement, and analytics. ACM Transactions on Internet Technology (TOIT), 17(3): 24:1–24:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Liu:2017:SLD

[LSK^{+}17b] Xumin Liu, Weishi Shi, Arpeet Kale, Chen Ding, and Qi Yu. Statistical learning of domainspecific quality-of-service features from user reviews. ACM Transactions on Internet Technology (TOIT), 17(2): 22:1–22:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lima:2019:IML

[LSLY19] Eduardo Lima, Weishi Shi, Xumin Liu, and Qi Yu. Integrating multilevel tag recommendation with external knowledge bases for automatic question answering. ACM Transactions on Internet Technology (TOIT), 19(3):34:1–34:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

                                        URL https://dl.acm.
                                        org/ft_gateway.cfm?
                                        id=3319528.

Luo:2021:NMH

[LSZ^{+}21] Ye Luo, Zehai Su, Wei Zheng, Zhaobin Chen, Fuqin Wang, Zhemin Zhang, and Jinjun Chen. A novel memory-hard password hashing scheme for blockchain-based cyberphysical systems. ACM Transactions on Internet Technology (TOIT), 21(2):42:1–42:21, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3408310.

Lippi:2016:AMS

[LT16] Marco Lippi and Paolo Torroni. Argumentation mining: State of the art and emerging trends. ACM Transactions on Internet Technology (TOIT), 16(2): 10:1–10:??, April 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lu:2021:ISS

[LWFD21] Huimin Lu, Liao Wu, Giancarlo Fortino, and Schahram Dustdar. Introduction to the special section on cognitive robotics on 5G/ 6G networks. ACM Transactions on Internet Technology (TOIT), 21 (4):91e:1–91e:3, November 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3476466.

Lan:2021:CED

[LWH^{+}21] Rushi Lan, Jing Wang, Wenming Huang, Zhenrong Deng, Xiyan Sun, Zhuo Chen, and Xiaonan Luo. Chinese emotional dialogue response generation via reinforcement learning. ACM Transactions on Internet Technology (TOIT), 21 (4):94:1–94:17, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3446390.

Liu:2021:OSR

[LWM^{+}21] Xuanzhe Liu, Shangguang Wang, Yun Ma, Ying Zhang, Qiaozhu Mei, Yunxin Liu, and Gang Huang. Operating systems for resourceadaptive intelligent software: Challenges and opportunities. ACM Transactions on Internet Technology (TOIT), 21(2): 27:1–27:19, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3425866.

Liang:2013:SSO

[LXC^{+}13] Yu-Li Liang, Xinyu Xing, Hanqiang Cheng, Jianxun Dang, Sui Huang, Richard Han, Xue Liu, Qin Lv, and Shivakant Mishra. SafeVchat: a system for obscene content detection in online video chat services. ACM Transactions on Internet Technology (TOIT), 12 (4):13:1–13:??, July 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Li:2012:TTO

[LXW^{+}12] Zhisheng Li, Xiangye Xiao, Meng Wang, Chong Wang, Xufa Wang, and Xing Xie. Towards the taxonomy-oriented categorization of yellow pages queries. ACM Transactions on Internet Technology (TOIT), 11(4): 16:1–16:??, March 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Liang:2022:MSA

[LXZ^{+}22] Wei Liang, Songyou Xie, Dafang Zhang, Xiong Li, and Kuan ching Li. A mutual security authentication method for RFID-PUF circuit based on deep learning. ACM Transactions on Internet Technology (TOIT), 22(2): 34:1–34:20, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3426968.

Li:2009:OBR

[LYF^{+}09] Xin Li, Jun Yan, Weiguo Fan, Ning Liu, Shuicheng Yan, and Zheng Chen. An online blog reading system by topic clustering and personalized ranking. ACM Transactions on Internet Technology (TOIT), 9(3):9:1– 9:??, July 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Liu:2018:JIO

[LYM^{+}18] Xuanzhe Liu, Meihua Yu, Yun Ma, Gang Huang, Hong Mei, and Yunxin Liu. i-Jacob: an internetware-oriented approach to optimizing computation-intensive mobile Web browsing. ACM Transactions on Internet Technology (TOIT), 18(2):14:1–14:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lu:2005:WMD

[LYW^{+}05] Hongjun Lu, Jeffrey Xu Yu, Guoren Wang, Shihui Zheng, Haifeng Jiang, Ge Yu, and Aoying Zhou. What makes the differences: benchmarking XML database implementations. ACM Transactions on Internet Technology (TOIT), 5(1):154– 194, February 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lu:2021:SSM

[LYW^{+}21] Wenpeng Lu, Rui Yu, Shoujin Wang, Can Wang, Ping Jian, and Heyan Huang. Sentence semantic matching based on 3D CNN for humanrobot language interaction. ACM Transactions on Internet Technology (TOIT), 21(4): 98:1–98:24, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3450520.

Li:2023:TTS

[LYW23] Ying Li, Yaxin Yu, and Xingwei Wang. Threetier storage framework based on TBchain and IPFS for protecting IoT security and privacy. ACM Transactions on Internet Technology (TOIT), 23(3): 37:1–37:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3549910.

Longo:2017:CSD

[LZBN17] Antonella Longo, Marco Zappatore, Mario Bochicchio, and Shamkant B. Navathe. Crowd-sourced data collection for urban monitoring via mobile sensors. ACM Transactions on Internet Technology (TOIT), 18(1): 5:1–5:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Lin:2021:MGM

[LZJ^{+}21] Zhiyang Lin, Jihua Zhu, Zutao Jiang, Yujie Li, Yaochen Li, and Zhongyu Li. Merging grid maps in diverse resolutions by the context-based descriptor. ACM Transactions on Internet Technology (TOIT), 21(4): 91:1–91:21, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3403948.

Liu:2022:TCE

[LZK^{+}22] Yi Liu, Ruihui Zhao, Jiawen Kang, Abdulsalam Yassine, Dusit Niyato, and Jialiang Peng. Towards communicationefficient and attackresistant federated edge learning for industrial Internet of Things. ACM Transactions on Internet Technology (TOIT), 22(3):59:1–59:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3453169.

Lu:2022:GSG

[LZW^{+}22] Jianfeng Lu, Zhao Zhang, Jiangtao Wang, Ruixuan Li, and Shaohua Wan. A green Stackelberggame incentive mechanism for multi-service exchange in mobile crowdsensing. ACM Transactions on Internet Technology (TOIT), 22(2): 31:1–31:29, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3421506.

Mostafaei:2023:SER

[MA23] Habib Mostafaei and Shafi Afridi. SDNenabled resource provisioning framework for geo-distributed streaming analytics. ACM Transactions on Internet Technology (TOIT), 23(1):18:1–18:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3571158.

Mazouzi:2019:DEE

[MAB19] Houssemeddine Mazouzi, Nadjib Achir, and Khaled Boussetta. DM2-ECOP: an efficient computation offloading policy for multi-user multi-cloudlet mobile edge computing environment. ACM Transactions on Internet Technology (TOIT), 19(2):24:1–24:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

                                        URL https://dl.acm.
                                        org/ft_gateway.cfm?
                                        id=3241666.

Moqurrab:2022:DCI

[MAK^{+}22] Syed Atif Moqurrab, Adeel Anjum, Abid Khan, Mansoor Ahmed, Awais Ahmad, and Gwanggil Jeon. Deep-confidentiality: an IoT-enabled privacypreserving framework for unstructured big biomedical data. ACM Transactions on Internet Technology (TOIT), 22(2): 42:1–42:21, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3421509.

Merialdo:2003:DDD

[MAM03] Paolo Merialdo, Paolo Atzeni, and Giansalvatore Mecca. Design and development of dataintensive web sites: The Araneus approach. ACM Transactions on Internet Technology (TOIT), 3(1): 49–92, February 2003. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Medjahed:2007:ISI

[MBB07] Brahim Medjahed, Athman Bouguettaya, and Boualem Benatallah. Introduction to special issue on semantic Web services. ACM Transactions on Internet Technology (TOIT), 8(1):1:1–1:??, November 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Mobasher:2007:TTR

[MBBW07] Bamshad Mobasher, Robin Burke, Runa Bhaumik, and Chad Williams. Toward trustworthy recommender systems: an analysis of attack models and algorithm robustness. ACM Transactions on Internet Technology (TOIT), 7(4):23:1–23:??, October 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Manolescu:2005:MDD

[MBC^{+}05] Ioana Manolescu, Marco Brambilla, Stefano Ceri, Sara Comai, and Piero Fraternali. Model-driven design and deployment of service-enabled Web applications. ACM Transactions on Internet Technology (TOIT), 5(3):439– 479, August 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Meier:2022:USM

[MBE22] Florian Meier, Alexander Bazo, and David Elsweiler. Using social media data to analyse issue engagement during the 2017 German Federal election. ACM Transactions on Internet Technology (TOIT), 22(1):25:1–25:25, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3467020.

Mancini:2017:IEL

[MBP^{+}17] Maurizio Mancini, Beatrice Biancardi, Florian Pecune, Giovanna Varni, Yu Ding, Catherine Pelachaud, Gualtiero Volpe, and Antonio Camurri. Implementing and evaluating a laughing virtual character. ACM Transactions on Internet Technology (TOIT), 17 (1):3:1–3:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Mrabet:2019:CTC

[MBS19] Manel Mrabet, Yosra Ben Saied, and Leila Azouz Saidane. CAN-TM: Chain augmented naïve Bayes-based trust model for reliable cloud service selection. ACM Transactions on Internet Technology (TOIT), 19(4):47:1–47:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3341732.

Moore:2018:RRB

[MCS18] Tyler Moore, Nicolas Christin, and Janos Szurdi. Revisiting the risks of Bitcoin currency exchange closure. ACM Transactions on Internet Technology (TOIT), 18 (4):50:1–50:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Murturi:2022:DDC

[MD22] Ilir Murturi and Schahram Dustdar. DECENT: a decentralized configurator for controlling elasticity in dynamic edge networks. ACM Transactions on Internet Technology (TOIT), 22(3):

78:1–78:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530692.

Merlino:2019:EWE

[MDDB19] Giovanni Merlino, Rustem Dautov, Salvatore Distefano, and Dario Bruneo. Enabling workload engineering in edge, fog, and cloud computing through OpenStackbased middleware. ACM Transactions on Internet Technology (TOIT), 19(2):28:1–28:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3309705.

Marques:2021:FBM

[MEAK^{+}21] Rafael Salema Marques, Gregory Epiphaniou, Haider Al-Khateeb, Carsten Maple, Mohammad Hammoudeh, Paulo André Lima De Castro, Ali Dehghantanha, and Kim Kwang Raymond Choo. A flow-based multi-agent data exfiltration detection architecture for ultra-low latency networks. ACM Transactions on Internet Technology (TOIT), 21(4): 103:1–103:30, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3419103.

Mezghani:2019:ACP

[MED19] Emna Mezghani, Ernesto Exposito, and Khalil Drira. An autonomic cognitive pattern for smart IoT-based system manageability: Application to comorbidity management. ACM Transactions on Internet Technology (TOIT), 19(1):8:1–8:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Mohammed:2021:BES

[MFR^{+}21] Sabah Mohammed, Jinan Fiaidhi, Carlos Ramos, Tai-Hoon Kim, Wai Chi Fang, and Tarek Abdelzaher. Blockchain in eCommerce: a special issue of the ACM Transactions on Internet of Things. ACM Transactions on Internet Technology (TOIT), 21(1): 4:11–4:55, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3445788.

Mrissa:2007:CBM

[MGB^{+}07] Michael Mrissa, Chirine Ghedira, Djamal Benslimane, Zakaria Maamar, Florian Rosenberg, and Schahram Dustdar. A context-based mediation approach to compose semantic Web services. ACM Transactions on Internet Technology (TOIT), 8(1):4:1–4:??, November 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Mehta:2021:PCT

[MGB^{+}21] Vikram Mehta, Daniel Gooch, Arosha Bandara, Blaine Price, and Bashar Nuseibeh. Privacy Care: a tangible interaction framework for privacy management. ACM Transactions on Internet Technology (TOIT), 21(1):25:1–25:32, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3430506.

Makhoul:2016:UEA

[MGHB16] Abdallah Makhoul, Christophe Guyeux, Mourad Hakem, and Jacques M. Bahi. Using an epidemiological approach to maximize data survival in the Internet of Things. ACM Transactions on Internet Technology (TOIT), 16(1):5:1–5:??, February 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Masud:2021:PTC

[MHA^{+}21] Mehedi Masud, M. Shamim Hossain, Hesham Alhumyani, Sultan S. Alshamrani, Omar Cheikhrouhou, Saleh Ibrahim, Ghulam Muhammad, Amr E. Eldin Rashed, and B. B. Gupta. Pre-trained convolutional neural networks for breast cancer detection using ultrasound images. ACM Transactions on Internet Technology (TOIT), 21 (4):85:1–85:17, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3418355.

Major:2022:AWS

[MHCF22] David Major, Danny Yuxing Huang, Marshini Chetty, and Nick Feamster. Alexa, who am I speaking to?: Understanding users’ ability to identify third-party apps on Amazon Alexa. ACM Transactions on Internet Technology (TOIT), 22(1):11:1–11:22, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3446389.

Mehrabi:2022:ECH

[MJ22] Mohamad Ali Mehrabi and Alireza Jolfaei. Efficient cryptographic hardware for safety message verification in Internet of Connected Vehicles. ACM Transactions on Internet Technology (TOIT), 22(4):86:1– 86:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3431499.

Maiti:2021:NII

[MKJB21] Somanka Maiti, Ashish Kumar, Smriti Jain, and Gaurav Bhatnagar. A novel image inpainting framework using regression. ACM Transactions on Internet Technology (TOIT), 21(3): 62:1–62:16, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3402177.

Mahmoud:2008:GES

[ML08] Qusay H. Mahmoud and Peter Langendoerfer. Guest editorial: Service-oriented computing. ACM Transactions on Internet Technology (TOIT), 8(3):11:1–11:??, May 2008. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Murata:2005:TXS

[MLMK05] Makoto Murata, Dongwon Lee, Murali Mani, and Kohsuke Kawaguchi. Taxonomy of XML schema languages using formal language theory. ACM Transactions on Internet Technology (TOIT), 5(4): 660–704, November 2005. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Mishra:2023:RTP

[MMI23] Pankaj Mishra, Ahmed Moustafa, and Takayuki Ito. Real-time pricingbased resource allocation in open market environments. ACM Transactions on Internet Technology (TOIT), 23(1): 1:1–1:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3465237.

Mehrabi:2021:PSC

[MMJ21] Mohamad Ali Mehrabi, Naila Mukhtar, and Alireza Jolfaei. Power side-channel analysis of RNS GLV ECC using machine and deep learning algorithms. ACM Transactions on Internet Technology (TOIT), 21(3):65:1–65:20, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3423555.

Ma:2022:PPD

[MMK^{+}22] Xindi Ma, Jianfeng Ma, Saru Kumari, Fushan Wei, Mohammad Shojafar, and Mamoun Alazab. Privacy-preserving distributed multi-task learning against inference attack in cloud computing. ACM Transactions on Internet Technology (TOIT), 22(2): 45:1–45:24, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3426969.

Molinaro:2014:PPA

[MMP^{+}14] Cristian Molinaro, Vincenzo Moscato, Antonio Picariello, Andrea Pugliese, Antonino Rullo, and V. S. Subrahmanian. PADUA: Parallel Architecture to Detect Unexplained Activities. ACM Transactions on Internet Technology (TOIT), 14 (1):3:1–3:??, July 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Malik:2016:SRE

[MMR16] Zaki Malik, Brahim Medjahed, and Abdelmounaam Rezgui. sCARE: Reputation estimation for uncertain Web services. ACM Transactions on Internet Technology (TOIT), 16(1):7:1–7:??, February 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Meiss:2011:PEI

[MMV11] Mark Meiss, Filippo Menczer, and Alessandro Vespignani. Properties and evolution of Internet traffic networks from anonymized flow data. ACM Transactions on Internet Technology (TOIT), 10(4): 15:1–15:??, March 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Moreau:2017:CFP

[Mor17] Luc Moreau. A canonical form for PROV documents and its application to equality, signature, and validation. ACM Transactions on Internet Technology (TOIT), 17 (4):35:1–35:??, September 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Mutschler:2014:ASP

[MP14] Christopher Mutschler and Michael Philippsen. Adaptive speculative processing of out-of-order event streams. ACM Transactions on Internet Technology (TOIT), 14 (1):4:1–4:??, July 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Min:2006:CEA

[MPC06] Jun-Ki Min, Myung-Jae Park, and Chin-Wan Chung. A compressor for effective archiving, retrieval, and updating of XML documents. ACM Transactions on Internet Technology (TOIT), 6(3): 223–258, August 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Muscariello:2023:SSR

[MPR^{+}23] Luca Muscariello, Michele Papalini, Olivier Roques, Mauro Sardara, and Arthur Tran Van. Securing scalable realtime multiparty communications with hybrid information-centric networking. ACM Transactions on Internet Technology (TOIT), 23(2): 33:1–33:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3593585.

Menczer:2004:TWC

[MPS04] Filippo Menczer, Gautam Pant, and Padmini Srinivasan. Topical web crawlers: Evaluating adaptive algorithms. ACM Transactions on Internet Technology (TOIT), 4(4):378– 419, November 2004. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Mistry:2022:LBC

[MQB22] Sajib Mistry, Lie Qu, and Athman Bouguettaya. Layer-based composite reputation bootstrapping. ACM Transactions on Internet Technology (TOIT), 22(1):13:1– 13:28, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3448610.

Manogaran:2022:GEI

[MQUXK22] Gunasekaran Manogaran, Hassan Qudrat-Ullah, Qin Xin, and Latifur Khan. Guest editorial introduction for the special section on deep learning algorithms and systems for enhancing security in cloud services. ACM Transactions on Internet Technology (TOIT), 22(2): 39e:1–39e:5, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3516806.

Mahmud:2019:LAA

[MRB19] Redowan Mahmud, Kotagiri Ramamohanarao, and Rajkumar Buyya. Latency-aware application module management for fog computing environments. ACM Transactions on Internet Technology (TOIT), 19(1): 9:1–9:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Manogaran:2022:TBA

[MRS^{+}22a] Gunasekaran Manogaran, Bharat S. Rawal, Vijayalakshmi Saravanan, Priyan M. K., Qin Xin, and P. Shakeel. Tokenbased authorization and authentication for secure Internet of Vehicles communication. ACM Transactions on Internet Technology (TOIT), 22(4):90:1–90:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3491202.

Manogaran:2022:OEC

[MRS^{+}22b] Gunasekaran Manogaran, Bharat S. Rawal, Houbing Song, Huihui Wang, Chinghsien Hsu, Vijayalakshmi Saravanan, Seifedine Nimer Kadry, and P. Mohamed Shakeel. Optimal energy-centric resource allocation and offloading scheme for green Internet of Things using machine learning. ACM Transactions on Internet Technology (TOIT), 22(2): 36:1–36:19, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3431500.

Ma:2023:VGG

[MRY^{+}23] Fuchen Ma, Meng Ren, Fu Ying, Wanting Sun, Houbing Song, Heyuan Shi, Yu Jiang, and Huizhong Li. V-Gas: Generating high gas consumption inputs to avoid out-of-gas vulnerability. ACM Transactions on Internet Technology (TOIT), 23(3): 40:1–40:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3511900.

Mok:2005:LAS

[MS05] Wilson Wai Ho Mok and R. P. Sundarraj. Learning algorithms for single-instance electronic negotiations using the time-dependent behavioral tactic. ACM Transactions on Internet Technology (TOIT), 5(1):195– 230, February 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Meo:2017:PAS

[MS17] Rosa Meo and Emilio Sulis. Processing affect in social media: a comparison of methods to distinguish emotions in tweets. ACM Transactions on Internet Technology (TOIT), 17(1): 7:1–7:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Masud:2021:CCS

[MSG^{+}21] Mehedi Masud, Parminder Singh, Gurjot Singh Gaba, Avinash Kaur, Roobaea Alrobaea Alghamdi, Mubarak Alrashoud, and Salman Ali Alqahtani. CROWD: Crow search and deep learning based feature extractor for classification of Parkinson’s disease. ACM Transactions on Internet Technology (TOIT), 21(3): 77:1–77:18, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3418500.

Mohammad:2017:SST

[MSK17] Saif M. Mohammad, Parinaz Sobhani, and Svetlana Kiritchenko. Stance and sentiment in tweets. ACM Transactions on Internet Technology (TOIT), 17(3): 26:1–26:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ma:2016:SAD

[MSW^{+}16] Jiangang Ma, Le Sun, Hua Wang, Yanchun Zhang, and Uwe Aickelin. Supervised anomaly detection in uncertain pseudoperiodic data streams. ACM Transactions on Internet Technology (TOIT), 16(1):4:1–4:??, February 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Maekawa:2012:CAW

[MYS^{+}12] Takuya Maekawa, Yutaka Yanagisawa, Yasushi Sakurai, Yasue Kishino, Koji Kamei, and Takeshi Okadome. Context-aware Web search in ubiquitous sensor environments. ACM Transactions on Internet Technology (TOIT), 11(3): 12:1–12:??, January 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Norman:2015:ITS

[NBFZ15] Timothy J. Norman, Suzanne Barber, Rino Falcone, and Jie Zhang. Introduction to theme section on trust in social networks and systems. ACM Transactions on Internet Technology (TOIT), 15(4): 12:1–12:??, December 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Neiat:2019:IBC

[NBM19] Azadeh Ghari Neiat, Athman Bouguettaya, and Sajib Mistry. Incentivebased crowdsourcing of hotspot services. ACM Transactions on Internet Technology (TOIT), 19 (1):5:1–5:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Nentwich:2002:XCC

[NCEF02] Christian Nentwich, Licia Capra, Wolfgang Emmerich, and Anthony Finkelstein. xlinkit: a consistency checking and smart link generation service. ACM Transactions on Internet Technology (TOIT), 2(2):151– 185, May 2002. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ng:2007:MUP

[NDL07] Wilfred Ng, Lin Deng, and Dik Lun Lee. Mining User preference using Spy voting for search engine personalization. ACM Transactions on Internet Technology (TOIT), 7(4):19:1–19:??, October 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Neves:2017:MPI

[NDO^{+}17] Vitor C. Neves, Daniel De Oliveira, Kary A. C. S. Ocaña, Vanessa Braganholo, and Leonardo Murta. Managing provenance of implicit data flows in scientific experiments. ACM Transactions on Internet Technology (TOIT), 17(4): 36:1–36:??, September 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Novgorodov:2020:DCC

[NGER20] Slava Novgorodov, Ido Guy, Guy Elad, and Kira Radinsky. Descriptions from the customers: Comparative analysis of review-based product description generation methods. ACM Transactions on Internet Technology (TOIT), 20(4):44:1–44:31, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3418202.

Ni:2021:HSN

[NLLC21] Pin Ni, Yuming Li, Gangmin Li, and Victor Chang. A hybrid Siamese neural network for natural language inference in cyber-physical systems. ACM Transactions on Internet Technology (TOIT), 21(2): 33:1–33:25, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3418208.

Novo:2019:MCT

[Nov19] Oscar Novo. Making constrained things reachable: a secure IP-agnostic NAT traversal approach for IoT. ACM Transactions on Internet Technology (TOIT), 19(1):3:1–3:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Nepal:2015:IBR

[NPP^{+}15] Surya Nepal, Cecile Paris, Payam Aghaei Pour, Jill Freyne, and Sanat Kumar Bista. Interaction-based recommendations for online communities. ACM Transactions on Internet Technology (TOIT), 15 (2):6:1–6:??, June 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Nguyen:2021:BBI

[NT21] Truc D. T. Nguyen and My T. Thai. A blockchain-based iterative double auction protocol using multiparty state channels. ACM Transactions on Internet Technology (TOIT), 21(2):39:1–39:22, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3389249.

Ning:2019:SAC

[NYB^{+}19] Xiaodong Ning, Lina Yao, Boualem Benatallah, Yihong Zhang, Quan Z. Sheng, and Salil S. Kanhere. Sourceaware crisis-relevant tweet identification and key information summarization. ACM Transactions on Internet Technology (TOIT), 19(3):

37:1–37:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Nguyen:2022:MCS

[NZ22] Tu N. Nguyen and Sherali Zeadally. Mobile crowd-sensing applications: Data redundancies, challenges, and solutions. ACM Transactions on Internet Technology (TOIT), 22(2): 48:1–48:15, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3431502.

Ni:2022:LPP

[NZQX22] Tongguang Ni, Jiaqun Zhu, Jia Qu, and Jing Xue. Labeling privacy protection SVM using privileged information for COVID-19 diagnosis. ACM Transactions on Internet Technology (TOIT), 22(3): 65:1–65:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3475868.

Otterbacher:2017:SMY

[OALA17] Jahna Otterbacher, Chee Siang Ang, Marina Litvak, and David Atkins. Show me you care: Trait empathy, linguistic style, and mimicry on Facebook. ACM Transactions on Internet Technology (TOIT), 17(1):6:1–6:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ouyang:2018:ASE

[OGP^{+}18] Xue Ouyang, Peter Garraghan, Bernhard Primas, David Mckee, Paul Townend, and Jie Xu. Adaptive speculation for efficient Internetware application execution in clouds. ACM Transactions on Internet Technology (TOIT), 18(2): 15:1–15:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

OMahony:2004:CRR

[OHKS04] Michael O’Mahony, Neil Hurley, Nicholas Kushmerick, and Guénolé Silvestre. Collaborative recommendation: a robustness analysis. ACM Transactions on Internet Technology (TOIT), 4(4): 344–377, November 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Otoum:2021:CSA

[OKM21] Safa Otoum, Burak Kantarci, and Hussein Mouftah. A comparative study of AI-based intrusion detection techniques in critical infrastructures. ACM Transactions on Internet Technology (TOIT), 21(4): 81:1–81:22, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3406093.

Ottenwalder:2014:MMA

[OKR^{+}14] Beate Ottenwälder, Boris Koldehofe, Kurt Rothermel, Kirak Hong, David Lillethun, and Umakishore Ramachandran. MCEP: a mobility-aware complex event processing system. ACM Transactions on Internet Technology (TOIT), 14(1): 6:1–6:??, July 2014. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Oberle:2005:SAD

[OSSV05] Daniel Oberle, Steffen Staab, Rudi Studer, and Raphael Volz. Supporting application development in the Semantic Web. ACM Transactions on Internet Technology (TOIT), 5(2):328– 358, May 2005. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ouni:2019:HAI

[OWK^{+}19] Ali Ouni, Hanzhang Wang, Marouane Kessentini, Salah Bouktif, and Katsuro Inoue. A hybrid approach for improving the design quality of Web service interfaces. ACM Transactions on Internet Technology (TOIT), 19 (1):4:1–4:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Panagidi:2020:TTC

[PACH20] K. Panagidi, C. Anagnostopoulos, A. Chalvatzaras, and S. Hadjiefthymiades. To transmit or not to transmit: Controlling communications in the mobile IoT domain. ACM Transactions on Internet Technology (TOIT), 20(3):22:1–22:23, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3369389.

Pranata:2013:MDR

[PAS13] Ilung Pranata, Rukshan Athauda, and Geoff Skinner. Modeling decentralized reputation-based trust for initial transactions in digital environments. ACM Transactions on Internet Technology (TOIT), 12(3): 8:1–8:??, May 2013. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Piccialli:2021:ISS

[PBJP21] Francesco Piccialli, Nik Bessis, Gwanggil Jeon, and Calton Pu. Introduction to the special section on Data Science for Cyber-Physical Systems. ACM Transactions on Internet Technology (TOIT), 21(2): 28e:1–28e:7, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3464766.

Silva:2022:FCP

[PBL^{+}22] Thiago Pereira Da Silva, Thais Batista, Frederico Lopes, Aluizio Rocha Neto, Flávia C. Delicato, Paulo F. Pires, and Atslands R. Da Rocha. Fog computing platforms for smart city applications: a survey. ACM Transactions on Internet Technology (TOIT), 22(4):96:1–96:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488585.

Pavlopoulou:2022:PEC

[PC22] Niki Pavlopoulou and Edward Curry. PoS-SUM: an entity-centric publish/subscribe system for diverse summarization in Internet of Things. ACM Transactions on Internet Technology (TOIT), 22(3): 73:1–73:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3507911.

Pore:2019:CEE

[PCBG19] Madhurima Pore, Vinaya Chakati, Ayan Banerjee, and Sandeep K. S. Gupta. ContextAiDe: End-to-end architecture for mobile crowd-sensing applications. ACM Transactions on Internet Technology (TOIT), 19(2):19:1–19:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3301444.

Pachilakis:2020:DIC

[PCP^{+}20] Michalis Pachilakis, Antonios A. Chariton, Panagiotis Papadopoulos, Panagiotis Ilia, Eirini Degkleri, and Evangelos P. Markatos. Design and implementation of a compressed certificate status protocol. ACM Transactions on Internet Technology (TOIT), 20(4):34:1–34:25, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392096.

Peng:2021:EDD

[PCV^{+}21] Cong Peng, Jianhua Chen, Pandi Vijayakumar, Neeraj Kumar, and Debiao He. Efficient distributed decryption scheme for IoT gatewaybased applications. ACM Transactions on Internet Technology (TOIT), 21(1):19:1–19:23, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3414475.

Plaza-Del-Arco:2020:DMX

[PDAMGULMV20] Flor-Miriam Plaza-Del-Arco, M. Dolores Molina-González, L. Alfonso Ureña-López, and M. Teresa Martín-Valdivia. Detecting misogyny and xenophobia in Spanish tweets using language technologies. ACM Transactions on Internet Technology (TOIT), 20(2): 12:1–12:19, May 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1145/3369869.

Pennekamp:2023:OTW

[PDF^{+}23] Jan Pennekamp, Markus Dahlmanns, Frederik Fuhrmann, Timo Heutmann, Alexander Kreppein, Dennis Grunert, Christoph Lange, Robert H. Schmitt, and Klaus Wehrle. Offering two-way privacy for evolved purchase inquiries. ACM Transactions on Internet Technology (TOIT), 23(4): 53:1–53:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3599968.

Pascoal:2020:MPA

[PDS20] Rui Pascoal, Ana De Almeida, and Rute C. Sofia. Mobile Pervasive Augmented Reality Systems — MPARS: The role of user preferences in the perceived quality of experience in outdoor applications. ACM Transactions on Internet Technology (TOIT), 20 (1):7:1–7:17, March 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1145/3375458.

Piccialli:2021:PAS

[PGP^{+}21] Francesco Piccialli, Fabio Giampaolo, Edoardo Prezioso, Danilo Crisci, and Salvatore Cuomo. Predictive analytics for smart parking: a deep learning approach in forecasting of IoT data. ACM Transactions on Internet Technology (TOIT), 21(3): 68:1–68:21, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3412842.

Peng:2018:COM

[PGT^{+}18] Xin Peng, Jingxiao Gu, Tian Huat Tan, Jun Sun, Yijun Yu, Bashar Nuseibeh, and Wenyun Zhao. CrowdService: Optimizing mobile crowdsourcing and service composition. ACM Transactions on Internet Technology (TOIT), 18(2): 19:1–19:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Peng:2021:EEP

[PHC^{+}21] Cong Peng, Debiao He, Jianhua Chen, Neeraj Kumar, and Muhammad Khurram Khan. EPRT: an efficient privacypreserving medical service recommendation and trust discovery scheme for eHealth system. ACM Transactions on Internet Technology (TOIT), 21(3):61:1–61:24, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3397678.

Pradhan:2021:GTS

[PHR^{+}21]  Buddhadeb  Pradhan,

Nirmal Baran Hui, Diptendu Sinha Roy, Gautam Srivastava, and Jerry Chun-Wei Lin. Gametheoretic strategic coordination and navigation of multiple wheeled robots. ACM Transactions on Internet Technology (TOIT), 21(4): 96:1–96:15, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3450521.

Pahl:2018:APC

[PJZ18] Claus Pahl, Pooyan Jamshidi, and Olaf Zimmermann. Architectural principles for cloud software. ACM Transactions on Internet Technology (TOIT), 18(2): 17:1–17:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Pal:2020:EPS

[PK20] Amitangshu Pal and Krishna Kant. Exploiting proxy sensing for efficient monitoring of large-scale sensor networks. ACM Transactions on Internet Technology (TOIT), 20(2): 14:1–14:31, May 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1145/3376919.

Palanisamy:2018:PPP

[PLZW18] Balaji Palanisamy, Ling Liu, Yang Zhou, and Qingyang Wang. Privacypreserving publishing of multilevel utilitycontrolled graph datasets. ACM Transactions on Internet Technology (TOIT), 18(2):24:1–24:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Prandi:2017:NTS

[PMFS17] Catia Prandi, Silvia Mirri, Stefano Ferretti, and Paola Salomoni. On the need of trustworthy sensing and crowdsourcing for urban accessibility in Smart City. ACM Transactions on Internet Technology (TOIT), 18(1):4:1–4:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Puliafito:2019:FCI

[PML^{+}19] Carlo Puliafito, Enzo Mingozzi, Francesco Longo, Antonio Puliafito, and Omer Rana. Fog computing for the Internet of Things: a survey. ACM Transactions on Internet Technology (TOIT), 19(2):18:1–18:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3301443.

Paun:2023:WBP

[PMN23] Iulia Paun, Yashar Moshfeghi, and Nikos Ntarmos. White box: On the prediction of collaborative filtering recommendation systems’ performance. ACM Transactions on Internet Technology (TOIT), 23(1): 8:1–8:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3554979.

Pussewalage:2019:ADA

[PO19] Harsha S. Gardiyawasam Pussewalage and Vladimir A. Oleshchuk. An anonymous delegatable attributebased credential scheme for a collaborative ehealth environment. ACM Transactions on Internet Technology (TOIT), 19 (3):41:1–41:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Park:2011:ACC

[PP11] Ki-Woong Park and Kyu Ho Park. AC-CENT: Cognitive cryptography plugged compression for SSL/TLSbased cloud computing services. ACM Transactions on Internet Technology (TOIT), 11(2): 7:1–7:??, December 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Pourmirza:2019:BRN

[PPDG19] Shaya Pourmirza, Sander Peters, Remco Dijkman, and Paul Grefen. BPMS-RA: a novel reference architecture for business process management systems. ACM Transactions on Internet Technology (TOIT), 19(1):13:1–13:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

                                        URL https://dl.acm.
                                        org/ft_gateway.cfm?
                                        id=3232677.

Petropoulos:2005:GQI

[PPV05] Michalis Petropoulos, Yannis Papakonstantinou, and Vasilis Vassalos.

Graphical query interfaces for semistructured data: the QURSED system. ACM Transactions on Internet Technology (TOIT), 5(2):390– 438, May 2005. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). See address correction [Vas05].

Platzer:2009:WSC

[PRD09] Christian Platzer, Florian Rosenberg, and Schahram Dustdar. Web service clustering using multidimensional angles as proximity measures. ACM Transactions on Internet Technology (TOIT), 9(3):11:1–11:??, July 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Pal:2020:SBN

[PRKD20] Amitangshu Pal, Mayank Raj, Krishna Kant, and Sajal K. Das. A smartphone-based network architecture for post-disaster operations using WiFi tethering. ACM Transactions on Internet Technology (TOIT), 20(1):6:1–6:27, March 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/doi/abs/10.1145/
3372145.

Paschalides:2020:MBD

[PSA^{+}20] Demetris Paschalides, Dimosthenis Stephanidis, Andreas Andreou, Kalia Orphanou, George Pallis, Marios D. Dikaiakos, and Evangelos Markatos. MANDOLA: a big-data processing and visualization platform for monitoring and detecting online hate speech. ACM Transactions on Internet Technology (TOIT), 20(2):11:1–11:21, March 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

                                        URL https://dl.acm.
                                        org/doi/abs/10.1145/
                                        3371276.

Pfitzner:2021:FLM

[PSA21] Bjarne Pfitzner, Nico Steckhan, and Bert Arnrich. Federated learning in a medical context: a systematic literature review. ACM Transactions on Internet Technology (TOIT), 21(2): 50:1–50:31, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3412357.

Pang:2010:PPS

[PSK10] Hweehwa Pang, Jialie Shen, and Ramayya Krishnan. Privacypreserving similaritybased text retrieval. ACM Transactions on Internet Technology (TOIT), 10(1):4:1–4:??, February 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Pu:2020:BAR

[PSL^{+}20] Calton Pu, Abhijit Suprem, Rodrigo Alves Lima, Aibek Musaev, De Wang, Danesh Irani, Steve Webb, and Joao Eduardo Ferreira. Beyond artificial reality: Finding and monitoring live events from social sensors. ACM Transactions on Internet Technology (TOIT), 20 (1):2:1–2:21, March 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1145/3374214.

Polachan:2022:DDS

[PSP22] Kurian Polachan, Chandramani Singh, and T. V. Prabhakar. Decentralized dynamic scheduling of TCPS flows and a simulator for time-sensitive networking. ACM Transactions on Internet Technology (TOIT), 22(4): 94:1–94:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3498729.

Pan:2024:ETI

[PSZ24] Bofeng Pan, Natalia Stakhanova, and Zhongwen Zhu. EtherShield: Time-interval analysis for detection of malicious behavior on Ethereum. ACM Transactions on Internet Technology (TOIT), 24(1):2:1–2:??, February 2024. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3633514.

Pitoura:2009:DFI

[PT09] Theoni Pitoura and Peter Triantafillou. Distribution fairness in Internetscale networks. ACM Transactions on Internet Technology (TOIT), 9(4):16:1–16:??, September 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Padget:2017:FGA

[PV17] Julian A. Padget and Wamberto W. Vasconcelos. Fine-grained access control via policycarrying data. ACM Transactions on Internet Technology (TOIT), 18 (3):31:1–31:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Perentis:2017:AUF

[PVL^{+}17] Christos Perentis, Michele Vescovi, Chiara Leonardi, Corrado Moiso, Mirco Musolesi, Fabio Pianesi, and Bruno Lepri. Anonymous or not? Understanding the factors affecting personal mobile data disclosure. ACM Transactions on Internet Technology (TOIT), 17 (2):13:1–13:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Pei:2022:NNP

[PWGQ22] Songwen Pei, Yusheng Wu, Jin Guo, and Meikang Qiu. Neural network pruning by recurrent weights for finance market. ACM Transactions on Internet Technology (TOIT), 22(3): 56:1–56:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3433547.

Pagani:2022:NNA

[PWSG22] Alessio Pagani, Zhuangkun Wei, Ricardo Silva, and Weisi Guo. Neural network approximation of graph Fourier transform for sparse sampling of networked dynamics. ACM Transactions on Internet Technology (TOIT), 22(1):21:1– 21:18, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3461838.

Qian:2019:SNA

[QLJ^{+}19] Jianwei Qian, Xiang-Yang Li, Taeho Jung, Yang Fan, Yu Wang, and Shaojie Tang. Social network de-anonymization: More adversarial knowledge, more users reidentified? ACM Transactions on Internet Technology (TOIT), 19(3): 33:1–33:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3310363.

Qiao:2022:MCB

[QZDG22] Yanchen Qiao, Weizhe Zhang, Xiaojiang Du, and Mohsen Guizani. Malware classification based on multilayer perception and Word2Vec for IoT security. ACM Transactions on Internet Technology (TOIT), 22(1):10:1–10:22, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3436751.

Ren:2022:TSM

[RAR22] Haoyu Ren, Darko Anicic, and Thomas A. Runkler. Towards semantic management of ondevice applications in industrial IoT. ACM Transactions on Internet Technology (TOIT), 22 (4):102:1–102:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3510820.

Ricci:2022:WDT

[RCM^{+}22] Alessandro Ricci, Angelo Croatti, Stefano Mariani, Sara Montagna, and Marco Picone. Web of digital twins. ACM Transactions on Internet Technology (TOIT), 22 (4):101:1–101:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3507909.

Ros:2015:COC

[RCP^{+}15] Santiago Pina Ros, Ángel Pina Canelles, Manuel Gil Pérez, Félix Gómez Mármol, and Gregorio Martínez Pérez. Chasing offensive conduct in social networks: a reputation-based practical approach for Frisber. ACM Transactions on Internet Technology (TOIT), 15(4): 15:1–15:??, December 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Rodriguez:2016:MQA

[RDC16] Carlos Rodríguez, Florian Daniel, and Fabio Casati. Mining and quality assessment of mashup model patterns with the crowd: a feasibility study. ACM Transactions on Internet Technology (TOIT), 16(3): 17:1–17:??, August 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Rohrer:2020:ERV

[RHT20] Elias Rohrer, Steffen Heidel, and Florian Tschorsch. Enabling reference verifiability for the World Wide Web with Webchain. ACM Transactions on Internet Technology (TOIT), 20(4):35:1–35:23, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392097.

Rezvani:2018:PAM

[RIB18] Mohsen Rezvani, Aleksandar Ignjatovic, and Elisa Bertino. A provenanceaware multi-dimensional reputation system for online rating systems. ACM Transactions on Internet Technology (TOIT), 18 (4):55:1–55:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Rinaldi:2009:ODA

[Rin09] Antonio M. Rinaldi. An ontology-driven approach for semantic information retrieval on the Web. ACM Transactions on Internet Technology (TOIT), 9(3):10:1–10:??, July 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Rahman:2022:LDH

[RKY^{+}22] Mohammad Saidur Rahman, Ibrahim Khalil, Xun Yi, Mohammed Atiquzzaman, and Elisa Bertino. A lossless datahiding based IoT data authenticity model in Edge-AI for connected living. ACM Transactions on Internet Technology (TOIT), 22(3): 57:1–57:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3453171.

Rosenthal:2017:DIM

[RM17] Sara Rosenthal and Kathleen Mckeown. Detecting influencers in multiple online genres. ACM Transactions on Internet Technology (TOIT), 17 (2):12:1–12:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Robu:2012:UPO

[RML12] Valentin Robu, Lonneke Mous, and Han La Poutré. Using priced options to solve the exposure problem in sequential auctions. ACM Transactions on Internet Technology (TOIT), 12(2):5:1–5:??, December 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Rawal:2022:MTS

[RMMH22] Bharat S. Rawal, Poongodi M., Gunasekaran Manogaran, and Mounir Hamdi. Multi-tier stack of block chain with proxy re-encryption method scheme on the Internet of Things platform. ACM Transactions on Internet Technology (TOIT), 22(2): 41:1–41:20, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3421508.

Rajab:2010:PTC

[RMP10] Moheeb Abu Rajab, Fabian Monrose, and Niels Provos. Peeking through the cloud: Client density estimation via DNS cache probing. ACM Transactions on Internet Technology (TOIT), 10(3):9:1–9:??, October 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Rathore:2017:HBI

[RPA^{+}17] M. Mazhar Rathore, Anand Paul, Awais Ahmad, Marco Anisetti, and Gwanggil Jeon. Hadoop-Based Intelligent Care System (HICS): Analytical approach for big data in IoT. ACM Transactions on Internet Technology (TOIT), 18(1): 8:1–8:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Rust:2022:RDC

[RPR22] Pierre Rust, Gauthier Picard, and Fano Ramparany. Resilient distributed constraint reasoning to autonomously configure and adapt IoT environments. ACM Transactions on Internet Technology (TOIT), 22 (4):100:1–100:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3507907.

Ren:2021:IVM

[RQL^{+}21] Yongjun Ren, Jian Qi, Yepeng Liu, Jin Wang, and Gwang-Jun Kim. Integrity verification mechanism of sensor data based on bilinear map accumulator. ACM Transactions on Internet Technology (TOIT), 21(1): 5:1–5:19, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3380749.

Ruffo:2009:PPR

[RS09] Giancarlo Ruffo and Rossano Schifanella. A peer-to-peer recommender system based on spontaneous affinities. ACM Transactions on Internet Technology (TOIT), 9(1): 4:1–4:??, February 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Roy:2017:FRD

[RSS17] Atanu Roy, Ayush Singhal, and Jaideep Srivastava. Formation and reciprocation of dyadic trust. ACM Transactions on Internet Technology (TOIT), 17(2): 15:1–15:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Rodriguez:2019:DDT

[RTcR19] Ricardo J. Rodríguez, Rafael Tolosana-calasanz, and Omer F. Rana. A dynamic data-throttling approach to minimize workflow imbalance. ACM Transactions on Internet Technology (TOIT), 19(3):32:1–32:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3278720.

Ravi:2022:DIU

[RTR^{+}22] Chandrasekar Ravi, Anmol Tigga, G. Thippa Reddy, Saqib Hakak, and Mamoun Alazab. Driver identification using optimized deep learning model in smart transportation. ACM Transactions on Internet Technology (TOIT), 22(4): 84:1–84:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3412353.

Ren:2020:SRE

[RWXC20] Hongshuai Ren, Yang Wang, Chengzhong Xu, and Xi Chen. SMig-RL: an evolutionary migration framework for cloud services based on deep reinforcement learning. ACM Transactions on Internet Technology (TOIT), 20(4):43:1– 43:18, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3414840.

Ruan:2017:MTB

[RZAD17] Yefeng Ruan, Ping Zhang, Lina Alfantoukh, and Arjan Durresi. Measurement theory-based trust management framework for online social communities. ACM Transactions on Internet Technology (TOIT), 17(2): 16:1–16:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ruan:2020:URP

[RZJ20] Na Ruan, Dongli Zhou, and Weijia Jia. Ursa: Robust performance for Nakamoto consensus with self-adaptive throughput. ACM Transactions on Internet Technology (TOIT), 20(4):41:1– 41:26, November 2020.

CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3412341.

Rodic:2022:MLS

[RZP^{+}22] Lea Dujić Rodić, Tomislav Zupanović, Toni Perković, Petar Solić, and Joel J. P. C. Rodrigues. Machine learning and soil humidity sensing: Signal strength approach. ACM Transactions on Internet Technology (TOIT), 22(2): 39:1–39:21, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3418207.

Saez:2018:UBD

[SAB^{+}18] Santiago Gómez Sáez, Vasilios Andrikopoulos, Marina Bitsaki, Frank Leymann, and André van Hoorn. Utility-based decision making for migrating cloud-based applications. ACM Transactions on Internet Technology (TOIT), 18(2): 22:1–22:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Stamatogiannakis:2017:PPP

[SABG17] Manolis Stamatogiannakis, Elias Athanasopoulos, Herbert Bos, and Paul Groth. PROV_{2R}: Practical provenance analysis of unstructured processes. ACM Transactions on Internet Technology (TOIT), 17(4): 37:1–37:??, September 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Saryazdi:2024:URL

[SABL24] Sepehr Saryazdi, Balsam Alkouz, Athman Bouguettaya, and Abdallah Lakhdari. Using reinforcement learning and error models for drone precise landing. ACM Transactions on Internet Technology (TOIT), 24 (3):14:1–14:??, 2024. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3670997.

Saeed:2016:IID

[SAJL16] Ahmed Saeed, Ali Ahmadinia, Abbas Javed, and Hadi Larijani. Intelligent intrusion detection in low-power IoTs. ACM Transactions on Internet Technology (TOIT), 16 (4):27:1–27:??, December 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Sekaran:2022:TTM

[SATPR22] Ramesh Sekaran, Fadi Al-Turjman, Rizwan Patan, and Velmani Ramasamy. Tripartite transmitting methodology for intermittently connected mobile network (ICMN). ACM Transactions on Internet Technology (TOIT), 22(4):89:1–89:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3433545.

Seeger:2020:OSH

[SBC20] Jan Seeger, Arne Bröring, and Georg Carle. Optimally self-healing IoT choreographies. ACM Transactions on Internet Technology (TOIT), 20(3):27:1–27:20, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3386361.

Shehab:2007:WSD

[SBG07] Mohamed Shehab, Kamal Bhattacharya, and Arif Ghafoor. Web services discovery in secure collaboration environments. ACM Transactions on Internet Technology (TOIT), 8(1):5:1– 5:??, November 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Sun:2019:MOO

[SCL^{+}19] Daniel Sun, Shiping Chen, Guoqiang Li, Yuanyuan Zhang, and Muhammad Atif. Multiobjective optimisation of online distributed software update for DevOps in clouds. ACM Transactions on Internet Technology (TOIT), 19 (3):43:1–43:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Song:2024:PEB

[SCLB24] Zhiyi Song, Dipankar Chaki, Abdallah Lakhdari, and Athman Bouguettaya. Positional encodingbased resident identification in multi-resident smart homes. ACM Transactions on Internet Technology (TOIT), 24(1):1:1–1:??, February 2024. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3631353.

Soldani:2022:MAR

[SCPB22] Jacopo Soldani, Marco Cameriero, Giulio Paparelli, and Antonio Brogi. Modelling and analysing replica- and fault-aware management of horizontally scalable applications. ACM Transactions on Internet Technology (TOIT), 22(3): 74:1–74:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3511302.

Shen:2017:TES

[SCW17] Haiying Shen, Harrison Chandler, and Haoyu Wang. Toward efficient short-video sharing in the YouTube social network. ACM Transactions on Internet Technology (TOIT), 18(3): 33:1–33:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Song:2021:SCS

[SCZ^{+}21] A. Qun Song, Yuhao Chen, Yan Zhong, Kun Lan, Simon Fong, and B. Rui Tang. A supplychain system framework based on Internet of Things using blockchain technology. ACM Transactions on Internet Technology (TOIT), 21(1): 13:1–13:24, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3409798.

Sheng:2012:ISI

[SD12] Quan Z. Sheng and Schahram Dustdar. Introduction to special issue on context-aware Web services for the future Internet. ACM Transactions on Internet Technology (TOIT), 11(3):9:1–9:??, January 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Sharma:2021:EUC

[SDB21] Tanusree Sharma, Hunter A. Dyer, and Masooda. Bashir. Enabling usercentered privacy controls for mobile applications: COVID-19 perspective. ACM Transactions on Internet Technology (TOIT), 21(1): 26:1–26:24, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3434777.

Silva:2014:RCW

[SdMA^{+}14] Thiago H. Silva, Pedro O. S. Vaz de Melo, Jussara M. Almeida, Juliana Salles, and Antonio A. F. Loureiro. Revealing the city that we cannot see. ACM Transactions on Internet Technology (TOIT), 14 (4):26:1–26:??, December 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Savaglio:2021:SDM

[SF21] Claudio Savaglio and Giancarlo Fortino. A simulation-driven methodology for IoT data mining based on edge computing. ACM Transactions on Internet Technology (TOIT), 21(2): 30:1–30:22, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3402444.

Saxena:2016:API

[SGC16] Neetesh Saxena, Santiago Grijalva, and Narendra S. Chaudhari. Authentication protocol for an IoT-enabled LTE network. ACM Transactions on Internet Technology (TOIT), 16(4): 25:1–25:??, December 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Shih:2019:GPB

[SGOS19] Timothy K. Shih, W. K.

T. M. Gunarathne, Ankhtuya Ochirbat, and Huang-Ming Su. Grouping peers based on complementary degree and social relationship using genetic algorithm. ACM Transactions on Internet Technology (TOIT), 19(1): 2:1–2:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Shorfuzzaman:2022:PAE

[SH22] Mohammad Shorfuzzaman and M. Shamim Hossain. Predictive analytics of energy usage by IoT-based smart home appliances for green urban development. ACM Transactions on Internet Technology (TOIT), 22(2):35:1–35:26, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3426970.

Schroeder:2006:WSU

[SHB06] Bianca Schroeder and Mor Harchol-Balter. Web servers under overload: How scheduling can help. ACM Transactions on Internet Technology (TOIT), 6(1):20– 52, February 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Stolfo:2006:BBM

[SHH^{+}06] Salvatore J. Stolfo, Shlomo Hershkop, Chia-Wei Hu, Wei-Jen Li, Olivier Nimeskern, and Ke Wang. Behaviorbased modeling and its application to Email analysis. ACM Transactions on Internet Technology (TOIT), 6(2):187– 221, May 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Singh:2013:TAU

[Sin13a] Munindar P. Singh.

TOIT administrative updates. ACM Transactions on Internet Technology (TOIT), 12(4): 11:1–11:??, July 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Singh:2013:VT

[Sin13b] Munindar P. Singh. Vision for TOIT. ACM Transactions on Internet Technology (TOIT), 12 (4):10:1–10:??, July 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Singh:2017:TR

[Sin17] Munindar P. Singh.

TOIT reviewers over 2015 and 2016. ACM Transactions on Internet Technology (TOIT), 18 (1):12:1–12:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Singh:2018:TR

[Sin18] Munindar P. Singh.

TOIT reviewers over 2017. ACM Transactions on Internet Technology (TOIT), 18(4):

57:1–57:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Shue:2013:RRC

[SK13] Craig A. Shue and Andrew J. Kalafut. Resolvers revealed: Characterizing DNS resolvers and their clients. ACM Transactions on Internet Technology (TOIT), 12 (4):14:1–14:??, July 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Saenko:2017:GAS

[SK17] Igor Saenko and Igor Kotenko. Genetic algorithms for solving problems of access control design and reconfiguration in computer networks. ACM Transactions on Internet Technology (TOIT), 18(3): 27:1–27:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Sachan:2024:SEQ

[SK24] Anuj Sachan and Neetesh Kumar. SDN-enabled quantized LQR for smart traffic light controller to optimize congestion. ACM Transactions on Internet Technology (TOIT), 24(1):7:1–7:??, February 2024. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3641104.

Sharma:2023:TCN

[SKA^{+}23] Sidharth Sharma, Aniruddha Kushwaha, Mohammad Alizadeh, George Varghese, and Ashwin Gumaste. Tuneman: Customizing networks to guarantee application bandwidth and latency. ACM Transactions on Internet Technology (TOIT), 23(1):20:1– 20:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3575657.

Shin:2022:SEH

[SKH22] Hyungjune Shin, Dongyoung Koo, and Junbeom Hur. Secure and efficient hybrid data deduplication in edge computing. ACM Transactions on Internet Technology (TOIT), 22(3): 80:1–80:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3537675.

Son:2022:EIP

[SL22] Heesuk Son and Dongman Lee. An efficient interaction protocol inference scheme for incompatible updates in IoT environments. ACM Transactions on Internet Technology (TOIT), 22(2):54:1–54:25, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3430501.

Sengupta:2020:PPN

[SLBD20] Binanda Sengupta, Yingjiu Li, Kai Bu, and Robert H. Deng. Privacy-preserving network path validation. ACM Transactions on Internet Technology (TOIT), 20(1):5:1–5:27, March 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:/

/dl.acm.org/doi/abs/
10.1145/3372046.

Singal:2022:QAM

[SLG^{+}22] Gaurav Singal, Vijay Laxmi, Manoj Singh Gaur, D. Vijay Rao, Riti Kushwaha, Deepak Garg, and Neeraj Kumar. QoS-aware meshbased multicast routing protocols in edge ad hoc networks: Concepts and challenges. ACM Transactions on Internet Technology (TOIT), 22(1): 1:1–1:27, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3428150.

Srivastava:2023:SSA

[SLPZ23] Gautam Srivastava, Jerry Chun-Wei Lin, Calton Pu, and Yudong Zhang. Special section on “Advances in Cyber-Manufacturing: Architectures, Challenges, & Future Research Directions”. ACM Transactions on Internet Technology (TOIT), 23(4): 49:1–49:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3627990.

Salomoni:2008:MBS

[SMFR08] Paola Salomoni, Silvia Mirri, Stefano Ferretti, and Marco Roccetti. A multimedia broker to support accessible and mobile learning through learning objects adaptation. ACM Transactions on Internet Technology (TOIT), 8(2):4:1– 4:??, February 2008. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Sherchan:2012:CSU

[SNBC12] Wanita Sherchan, Surya Nepal, Athman Bouguettaya, and Shiping Chen.

Context-sensitive user interfaces for semantic services. ACM Transactions on Internet Technology (TOIT), 11(3):14:1– 14:??, January 2012. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Shao:2017:ECI

[SO17] Jianhua Shao and Hoang Ong. Exploiting contextual information in attacking set-generalized transactions. ACM Transactions on Internet Technology (TOIT), 17 (4):40:1–40:??, September 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Sekaran:2021:NAE

[SPAT21] Ramesh Sekaran, Rizwan Patan, and Fadi Al-Turjman. A novel approach for efficient packet transmission in volunteered computing MANET. ACM Transactions on Internet Technology (TOIT), 21(4): 100:1–100:15, November 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3418203.

Song:2023:CHD

[SPCC23] Pei-Cheng Song, Jeng-Shyang Pan, Han-Chieh Chao, and Shu-Chuan Chu. Collaborative hotspot data collection with drones and 5G edge computing in smart city. ACM Transactions on Internet Technology (TOIT), 23(4):55:1– 55:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3617373.

Shankar:2022:SDL

[SPE^{+}22] K. Shankar, Eswaran Perumal, Mohamed Elhoseny, Fatma Taher, B. B. Gupta, and Ahmed A. Abd El-Latif. Synergic deep learning for smart health diagnosis of COVID-19 for connected living and smart cities. ACM Transactions on Internet Technology (TOIT), 22(3): 61:1–61:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3453168.

Stergiou:2022:IFB

[SPG22] Christos L. Stergiou, Konstantinos E. Psannis, and Brij B. Gupta. In-FeMo: Flexible big data management through a federated cloud system. ACM Transactions on Internet Technology (TOIT), 22(2): 46:1–46:22, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3426972.

Stein:2009:FPW

[SPJ09] Sebastian Stein, Terry R. Payne, and Nicholas R. Jennings. Flexible provisioning of Web service workflows. ACM Transactions on Internet Technology (TOIT), 9(1):2:1– 2:??, February 2009. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Shudrenko:2022:NAE

[SPKTG22] Yevhenii Shudrenko, Daniel Plöger, Koojana Kuladinithi, and Andreas Timm-Giel. A novel approach to enhance the end-to-end quality of service for avionic wireless sensor networks. ACM Transactions on Internet Technology (TOIT), 22(4):95:1–95:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3520441.

Sun:2013:IUP

[SPM^{+}13] San-Tsai Sun, Eric Pospisil, Ildar Muslukhov, Nuray Dindar, Kirstie Hawkey, and Konstantin Beznosov. Investigating users’ perspectives of Web single sign-on: Conceptual gaps and acceptance model. ACM Transactions on Internet Technology (TOIT), 13(1): 2:1–2:??, November 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Sayyadi:2013:GAA

[SR13] Hassan Sayyadi and Louiqa Raschid. A graph analytical approach for topic detection. ACM Transactions on Internet Technology (TOIT), 13(2):4:1–4:??, December 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Sood:2022:ETI

[SRK22] Sandeep Kumar Sood, Keshav Singh Rawat, and Dheeraj Kumar. Emerging trends of ICT in airborne disease prevention. ACM Transactions on Internet Technology (TOIT), 22(4):110:1– 110:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3564783.

Szykman:2006:DIW

[SS06] Simon Szykman and Ram D. Sriram. Design and implementation of the Web-enabled NIST design repository. ACM Transactions on Internet Technology (TOIT), 6(1): 85–116, February 2006. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Shen:2011:ADC

[SS11] Haifeng Shen and Chengzheng Sun. Achieving data consistency by contextualization in Web-Based collaborative applications. ACM Transactions on Internet Technology (TOIT), 10(4): 13:1–13:??, March 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Shi:2020:PSA

[SS20] Junyang Shi and Mo Sha. Parameter self-adaptation for industrial wireless sensor–actuator networks. ACM Transactions on Internet Technology (TOIT), 20(3):28:1–28:28, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3388240.

Shahid:2021:MLB

[SSA^{+}21] Huniya Shahid, Munam Ali Shah, Ahmad Almogren, Hasan Ali Khattak, Ikram Ud Din, Neeraj Kumar, and Carsten Maple. Machine learning-based mist computing enabled Internet of Battlefield Things. ACM Transactions on Internet Technology (TOIT), 21(4): 101:1–101:26, November 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3418204.

Sheng:2023:GEI

[SSC23] Quan Z. Sheng, Arun Kumar Sangaiah, and Ankit Chaudhary. Guest editors’ introduction for special issue on applications of computational linguistics in multimedia IoT services. ACM Transactions on Internet Technology (TOIT), 23(2): 24:1–24:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3591355.

Sofia:2020:ISI

[SSKW20] Rute C. Sofia, Eve M.

Schooler, Dirk Kutscher, and Chris Winkler. Introduction to the special issue on evolution of IoT networking architectures. ACM Transactions on Internet Technology (TOIT), 20(3):20:1–20:2, October 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3406087.

Siboni:2016:AST

[SST^{+}16] Shachar Siboni, Asaf Shabtai, Nils O. Tippenhauer, Jemin Lee, and Yuval Elovici. Advanced security testbed framework for wearable IoT devices. ACM Transactions on Internet Technology (TOIT), 16(4): 26:1–26:??, December 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Samie:2019:OOO

[STB^{+}19] Farzad Samie, Vasileios Tsoutsouras, Lars Bauer, Sotirios Xydis, Dimitrios Soudris, and Jörg Henkel. Oops: Optimizing operation-mode selection for IoT edge devices. ACM Transactions on Internet Technology (TOIT), 19(2): 22:1–22:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
ft_gateway.cfm?id=3230642.

Singh:2021:JEC

[STJ^{+}21] A. K. Singh, S. Thakur, Alireza Jolfaei, Gautam Srivastava, MD. Elhoseny, and A. Mohan. Joint encryption and compression-based watermarking technique for security of digital documents. ACM Transactions on Internet Technology (TOIT), 21(1):18:1–18:20, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3414474.

Stolba:2017:QPL

[STK17] Michal Stolba, Jan Tozicka, and Antonín Komenda. Quantifying privacy leakage in multi-agent planning. ACM Transactions on Internet Technology (TOIT), 18(3): 28:1–28:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Singh:2021:ISS

[SWAHP21] Amit Kumar Singh, Jonathan Wu, Ali Al-Haj, and Calton Pu. Introduction to the special section on security and privacy of medical data for smart healthcare. ACM Transactions on Internet Technology (TOIT), 21(3):53:1–53:4, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3460870.

Sutcliffe:2015:MRT

[SWD15] Alistair G. Sutcliffe, Di Wang, and Robin I. M. Dunbar. Modelling the role of trust in social relationships. ACM Transactions on Internet Technology (TOIT), 15 (4):16:1–16:??, December 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Sun:2021:RRS

[SXZ^{+}21] You Sun, Rui Xue, Rui Zhang, Qianqian Su, and Sheng Gao. RTChain: a reputation system with transaction and consensus incentives for e-commerce blockchain. ACM Transactions on Internet Technology (TOIT), 21(1): 15:1–15:24, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3430502.

Shen:2022:CDA

[SZT22] Chaonan Shen, Kai Zhang, and Jinshan Tang. A COVID-19 detection algorithm using deep features and discrete social learning particle swarm optimization for edge computing devices. ACM Transactions on Internet Technology (TOIT), 22(3): 58:1–58:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3453170.

Tajalizadehkhoob:2018:RAB

[TBG^{+}18] Samaneh Tajalizadehkhoob, Rainer Böhme, Carlos Gañán, Maciej Korczyński, and Michel Van Eeten. Rotten apples or bad harvest? What we are measuring when we are measuring abuse. ACM Transactions on Internet Technology (TOIT), 18 (4):49:1–49:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Taherkordi:2019:CDR

[TEMH19] Amir Taherkordi, Frank Eliassen, Michael Mcdonald, and Geir Horn. Context-driven and realtime provisioning of datacentric IoT services in the cloud. ACM Transactions on Internet Technology (TOIT), 19(1): 7:1–7:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Tsai:2021:EHO

[TF21] Chun-Wei Tsai and Zhi-Yan Fang. An effective hyperparameter optimization algorithm for DNN to predict passengers at a metro station. ACM Transactions on Internet Technology (TOIT), 21(2): 32:1–32:24, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3410156.

Tsigkanos:2020:CDT

[TGBG20] Christos Tsigkanos, Martin Garriga, Luciano Baresi, and Carlo Ghezzi. Cloud deployment tradeoffs for the analysis of spatially distributed Internet of Things systems. ACM Transactions on Internet Technology (TOIT), 20(2): 17:1–17:23, May 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/abs/
                                        10.1145/3381452.

Toch:2007:SAA

[TGRBD07] Eran Toch, Avigdor Gal, Iris Reinhartz-Berger, and Dov Dori. A semantic approach to approximate service retrieval. ACM Transactions on Internet Technology (TOIT), 8(1):2:1– 2:??, November 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Thiemann:2005:EDS

[Thi05] Peter Thiemann. An embedded domain-specific language for type-safe server-side Web scripting. ACM Transactions on Internet Technology (TOIT), 5(1):1– 46, February 2005. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Tsoi:2006:CCP

[THS06] Ah Chung Tsoi, Markus Hagenbuchner, and Franco Scarselli. Computing customized page ranks. ACM Transactions on Internet Technology (TOIT), 6(4):381– 414, November 2006. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Tiwari:2022:SES

[TJGY22] Prayag Tiwari, Amit Kumar Jaiswal, Sahil Garg, and Ilsun You. SANTM: Efficient self-attention-driven network for text matching. ACM Transactions on Internet Technology (TOIT), 22(3): 55:1–55:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3426971.

Tu:2008:NLA

[TJLC08] Xuping Tu, Hai Jin, Xiaofei Liao, and Jiannong Cao. Nearcast: a localityaware P2P live streaming approach for distance education. ACM Transactions on Internet Technology (TOIT), 8(2):2:1– 2:??, February 2008. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Totok:2011:ESU

[TK11] Alexander Totok and Vijay Karamcheti. Exploiting service usage information for optimizing server resource management. ACM Transactions on Internet Technology (TOIT), 11(1):1:1–1:??, July 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Tyson:2012:JMP

[TMK^{+}12] Gareth Tyson, Andreas Mauthe, Sebastian Kaune, Paul Grace, Adel Taweel, and Thomas Plagemann. Juno: a middleware platform for supporting delivery-centric applications. ACM Transactions on Internet Technology (TOIT), 12(2):4:1–4:??, December 2012. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Tedeschi:2022:OTF

[TNJJ22] Enrico Tedeschi, Tor-Arne S. Nordmo, Dag Johansen, and Håvard D. Johansen. On optimizing transaction fees in bitcoin using AI: Investigation on miners inclusion pattern. ACM Transactions on Internet Technology (TOIT), 22(3): 77:1–77:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3528669.

Turner:2010:MBB

[TPK10] David Michael Turner, Vassilis Prevelakis, and Angelos D. Keromytis. A market-based bandwidth charging framework. ACM Transactions on Internet Technology (TOIT), 10(1):1:1–1:??, February 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Tian:2022:IBP

[TPQC22] Hui Tian, Fang Peng, Hanyu Quan, and Chin-Chen Chang. Identitybased public auditing for cloud storage of Internetof-Vehicles data. ACM Transactions on Internet Technology (TOIT), 22(4):88:1–88:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3433543.

Tanveer:2021:LSL

[TSM21] M. Tanveer, S. Sharma, and K. Muhammad. Large-scale least squares twin SVMs. ACM Transactions on Internet Technology (TOIT), 21(2): 29:1–29:19, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3398379.

Trevisan:2023:ADE

[TSM^{+}23] Martino Trevisan, Francesca Soro, Marco Mellia, Idilio Drago, and Ricardo Morla. Attacking DoH and ECH: Does server name encryption protect users’ privacy? ACM Transactions on Internet Technology (TOIT), 23(1):19:1–19:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3570726.

Tata:2019:GEI

[TSS19] Samir Tata, Quan Z.

Sheng, and Eleni Stroulia. Guest Editors’ introduction for special issue on service management for the Internet of Things. ACM Transactions on Internet Technology (TOIT), 19(1): 6:1–6:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Tan:2021:BEA

[TSY^{+}21] Liang Tan, Na Shi, Keping Yu, Moayad Aloqaily, and Yaser Jararweh. A blockchainempowered access control framework for smart devices in green Internet of Things. ACM Transactions on Internet Technology (TOIT), 21(3): 80:1–80:20, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3433542.

Ullah:2022:IBC

[UNBAT22] Farhan Ullah, Muhammad Rashid Naeem, Abdullah S. Bajahzar, and Fadi Al-Turjman. IoTbased cloud service for secured Android markets using PDG-based deep learning classification. ACM Transactions on Internet Technology (TOIT), 22(2): 40:1–40:17, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3418206.

Ungureanu:2005:UCP

[Ung05] Victoria Ungureanu. Using certified policies to regulate E-commerce transactions. ACM Transactions on Internet Technology (TOIT), 5(1): 129–153, February 2005. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Urgaonkar:2009:ROA

[USR09] Bhuvan Urgaonkar, Prashant Shenoy, and Timothy Roscoe. Resource overbooking and application profiling in a shared Internet hosting platform. ACM Transactions on Internet Technology (TOIT), 9(1):1:1– 1:??, February 2009. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ulusoy:2022:PPA

[UY22] Onuralp Ulusoy and Pinar Yolum. PANOLA: a personal assistant for supporting users in preserving privacy. ACM Transactions on Internet Technology (TOIT), 22(1):27:1–27:32, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3471187.

Weth:2017:CPS

[VAK17] Christian Von Der Weth, Ashraf M. Abdul, and Mohan Kankanhalli. Cyberphysical social networks. ACM Transactions on Internet Technology (TOIT), 17(2):17:1–17:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

VonDerWeth:2019:CCD

[VAKK19] Christian Von Der Weth, Ashraf Abdul, Abhinav R. Kashyap, and Mohan S. Kankanhalli. CloseUp — a communitydriven live online search engine. ACM Transactions on Internet Technology (TOIT), 19(3): 39:1–39:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

VanEngelen:2008:FSO

[Van08] Robert A. Van Engelen. A framework for serviceoriented computing with C and C++ Web service components. ACM Transactions on Internet Technology (TOIT), 8(3): 12:1–12:??, May 2008. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Varshney:2003:LMM

[Var03] Upkar Varshney. Location management for mobile commerce applications in wireless Internet environment. ACM Transactions on Internet Technology (TOIT), 3(3): 236–255, August 2003. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Vassalos:2005:C

[Vas05] Vasilis Vassalos. Corrigenda. ACM Transactions on Internet Technology (TOIT), 5(3):570, August 2005. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). Address correction for [PPV05].

Vali:2024:RRC

[VAS24] Ali Akbar Vali, Sadoon Azizi, and Mohammad Shojafar. RESP: a recursive clustering approach for edge server placement in mobile edge computing. ACM Transactions on Internet Technology (TOIT), 24(3): 13:1–13:??, 2024. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3666091.

Vasconcelos:2019:CFM

[VASD19] D. R. Vasconcelos, R. M. C.

Andrade, V. Severino, and J. N. De Souza. Cloud, fog, or mist in IoT? That is the question. ACM Transactions on Internet Technology (TOIT), 19(2): 25:1–25:??, April 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
ft_gateway.cfm?id=3309709.

Verde:2022:DLA

[VBD^{+}22] Laura Verde, Nadia Brancati, Giuseppe De Pietro, Maria Frucci, and Giovanna Sannino. A deep learning approach for voice disorder detection for smart connected living environments. ACM Transactions on Internet Technology (TOIT), 22(1): 8:1–8:16, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3433993.

vanderAalst:2008:CCS

[vdADO^{+}08] Wil M. P. van der Aalst, Marlon Dumas, Chun Ouyang, Anne Rozinat, and Eric Verbeek. Conformance checking of service behavior. ACM Transactions on Internet Technology (TOIT), 8 (3):13:1–13:??, May 2008. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Verdiesen:2018:MMA

[VDV18] Ilse Verdiesen, Virginia Dignum, and Jeroen Van Den Hoven. Measuring moral acceptability in edeliberation: a practical application of ethics by participation. ACM Transactions on Internet Technology (TOIT), 18 (4):43:1–43:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Vosecky:2014:ISA

[VJL^{+}14] Jan Vosecky, Di Jiang, Kenneth Wai-Ting Leung, Kai Xing, and Wilfred Ng. Integrating social and auxiliary semantics for multifaceted topic modeling in Twitter. ACM Transactions on Internet Technology (TOIT), 14 (4):27:1–27:??, December 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Villela:2007:PSA

[VPR07] Daniel Villela, Prashant Pradhan, and Dan Rubenstein. Provisioning servers in the application tier for e-commerce systems. ACM Transactions on Internet Technology (TOIT), 7(1):7:1– 7:??, February 2007. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Vogler:2016:SFP

[VSID16] Michael Vögler, Johannes M. Schleicher, Christian Inzinger, and Schahram Dustdar. A scalable framework for provisioning large-scale IoT deployments. ACM Transactions on Internet Technology (TOIT), 16(2):11:1–11:??, April 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Vargas-Solar:2022:GHE

[VSKEOZM22] Genoveva Vargas-Solar, Maysaa Khalil, Javier A. Espinosa-Oviedo, and José-Luis Zechinelli-Martini. GREENHOME: a household energy consumption and CO_{2} footprint metering environment. ACM Transactions on Internet Technology (TOIT), 22(3): 72:1–72:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3505264.

Wang:2017:RTT

[WARCD17] Di Wang, Ahmad Al-Rubaie, Sandra Stincić Clarke, and John Davies. Real-time traffic event detection from social media. ACM Transactions on Internet Technology (TOIT), 18(1):9:1–9:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wang:2020:ESF

[WCC20] Meng Wang, Bo Cheng, and Junliang Chen. An efficient service function chaining placement algorithm in mobile edge computing. ACM Transactions on Internet Technology (TOIT), 20(4): 32:1–32:21, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3388241.

Wang:2023:BEB

[WCX^{+}23] Jin Wang, Jiahao Chen, Neal Xiong, Osama Alfarraj, Amr Tolba, and Yongjun Ren. S-BDS: an effective blockchainbased data storage scheme in zero-trust IoT. ACM Transactions on Internet Technology (TOIT), 23(3):42:1–42:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3511902.

Wu:2023:DTI

[WCY^{+}23] Yirui Wu, Hao Cao, Guoqiang Yang, Tong Lu, and Shaohua Wan. Digital twin of intelligent small surface defect detection with cybermanufacturing systems. ACM Transactions on Internet Technology (TOIT), 23(4):51:1–51:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3571734.

Wang:2021:MGC

[WCZ^{+}21] Wei Wang, Junyang Chen, Yushu Zhang, Zhiguo Gong, Neeraj Kumar, and Wei Wei. A multi-graph convolutional network framework for tourist flow prediction. ACM Transactions on Internet Technology (TOIT), 21(4):

106:1–106:13, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3424220.

Wu:2024:OSD

[WDK^{+}24] Jiashu Wu, Hao Dai, Kenneth B. Kent, Jerome Yen, Chengzhong Xu, and Yang Wang. Open set dandelion network for IoT intrusion detection. ACM Transactions on Internet Technology (TOIT), 24(1):4:1–4:??, February 2024. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3639822.

Weber:2017:FAI

[Web17] Steven Weber. Facilitating adoption of Internet technologies and services with externalities via cost subsidization. ACM Transactions on Internet Technology (TOIT), 17 (4):38:1–38:??, September 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wilkin:2014:DFT

[WEJ14] Gregory Aaron Wilkin, Patrick Eugster, and K. R. Jayaram. Decentralized fault-tolerant event correlation. ACM Transactions on Internet Technology (TOIT), 14 (1):5:1–5:??, July 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wu:2020:EIM

[WFZ^{+}20] Xudong Wu, Luoyi Fu, Zixin Zhang, Huan Long, Jingfan Meng, Xinbing Wang, and Guihai Chen. Evolving influence maximization in evolving networks. ACM Transactions on Internet Technology (TOIT), 20(4): 40:1–40:31, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3409370.

Wazid:2023:BEN

[WG23] Mohammad Wazid and Prosanta Gope. BACKM-EHA: a novel blockchainenabled security solution for IoMT-based ehealthcare applications. ACM Transactions on Internet Technology (TOIT), 23(3):39:1–39:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3511898.

Wang:2024:EVF

[WGW^{+}24] Zichen Wang, Xiangshan Gao, Cong Wang, Peng Cheng, and Jiming Chen. Efficient vertical federated unlearning via fast retraining. ACM Transactions on Internet Technology (TOIT), 24(2): 11:1–11:??, May 2024. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3657290.

Wu:2022:IDD

[WHM^{+}22] Chao Wu, Shingo Horiuchi, Kenji Murase, Hiroaki Kikushima, and Kenichi Tayama. An intent-driven DaaS management framework to enhance user quality of experience. ACM Transactions on Internet Technology (TOIT), 22(4):98:1–98:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488586.

Williamson:2002:FEW

[Wil02] Carey Williamson. On filter effects in web caching hierarchies. ACM Transactions on Internet Technology (TOIT), 2(1):47– 77, February 2002. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wang:2022:IGL

[WJL^{+}22] Jingjing Wang, Wenjun Jiang, Kenli Li, Guojun Wang, and Keqin Li. Incremental group-level popularity prediction in online social networks. ACM Transactions on Internet Technology (TOIT), 22(1):20:1–20:26, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3461839.

Willnecker:2018:MOO

[WK18] Felix Willnecker and Helmut Krcmar. Multiobjective optimization of deployment topologies for distributed applications. ACM Transactions on Internet Technology (TOIT), 18(2): 21:1–21:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wong:2007:AWI

[WL07] Tak-Lam Wong and Wai Lam. Adapting Web information extraction knowledge via mining site-invariant and sitedependent features. ACM Transactions on Internet Technology (TOIT), 7(1): 6:1–6:??, February 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wang:2023:MAR

[WL23] Yu-Jhen Wang and Anthony J. T. Lee. Movie account recommendation on Instagram. ACM Transactions on Internet Technology (TOIT), 23(1):23:1–23:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3579852.

Wang:2022:NTN

[WLB22] Changda Wang, Xiaowei Li, and Elisa Bertino. Network temperature: a novel statistical index for networks measurement and management. ACM Transactions on Internet Technology (TOIT), 22(3): 66:1–66:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3511093.

Wang:2013:WAM

[WLL^{+}13] Meng Wang, Guangda Li, Zheng Lu, Yue Gao, and Tat-Seng Chua. When Amazon meets Google: Product visualization by exploring multiple Web sources. ACM Transactions on Internet Technology (TOIT), 12(4): 12:1–12:??, July 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wang:2023:PAT

[WLW^{+}23] Fan Wang, Guangshun Li, Yilei Wang, Wajid Rafique, Mohammad R. Khosravi, Guanfeng Liu, Yuwen Liu, and Lianyong Qi. Privacy-aware traffic flow prediction based on multi-party sensor data with zero trust in Smart City. ACM Transactions on Internet Technology (TOIT), 23(3): 44:1–44:??, August 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3511904.

Wang:2021:BBP

[WMG^{+}21] Hao Wang, Shenglan Ma, Chaonian Guo, Yulei Wu, Hong-Ning Dai, and Di Wu. Blockchainbased power energy trading management. ACM Transactions on Internet Technology (TOIT), 21(2):43:1–43:16, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3409771.

Wu:2022:ATT

[WMW^{+}22] Tingmin Wu, Wanlun Ma, Sheng Wen, Xin Xia, Cecile Paris, Surya Nepal, and Yang Xiang. Analysis of trending topics and text-based channels of information delivery in cybersecurity. ACM Transactions on Internet Technology (TOIT), 22(2): 52:1–52:27, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3483332.

Wang:2020:ELE

[WMWM20] Shuo Wang, Aishan Maoliniyazi, Xinle Wu, and Xiaofeng Meng. Emo2Vec: Learning emotional embeddings via multi-emotion category. ACM Transactions on Internet Technology (TOIT), 20(2): 13:1–13:17, May 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/abs/
                                        10.1145/3372152.

Wang:2022:DLB

[WNN^{+}22] Xiaojie Wang, Laisen Nie, Zhaolong Ning, Lei Guo, Guoyin Wang, Xinbo Gao, and Neeraj Kumar. Deep learningbased network traffic prediction for secure backbone networks in Internet of Vehicles. ACM Transactions on Internet Technology (TOIT), 22(4):87:1–87:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3433548.

Wang:2019:BCB

[WQC^{+}19] Kai Wang, Wei Quan, Nan Cheng, Mingyuan Liu, Yu Liu, and H. Anthony Chan. Betweenness centrality based software defined routing: Observation from practical Internet datasets. ACM Transactions on Internet Technology (TOIT), 19(4):50:1–50:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3355605.

Waldman:2001:ARP

[WRC01] Marc Waldman, Aviel D.

Rubin, and Lorrie Faith Cranor. The architecture of robust publishing systems. ACM Transactions on Internet Technology (TOIT), 1(2):199– 230, November 2001. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wei:2021:SSA

[WRWM21] Wei Wei, Ammar Rayes, Wei Wang, and Yiduo Mei. Special section on AI-empowered Internet of Things for smart cities. ACM Transactions on Internet Technology (TOIT), 21(3):64:1–64:3, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3460868.

Wachsmuth:2017:UMD

[WS17] Henning Wachsmuth and Benno Stein. A universal model for discourselevel argumentation analysis. ACM Transactions on Internet Technology (TOIT), 17(3): 28:1–28:??, July 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wu:2021:MTA

[WSLT21] Jimmy Ming-Tai Wu, Gautam Srivastava, Jerry Chun-Wei Lin, and Qian Teng. A multi-threshold ant colony system-based sanitization model in shared medical environments. ACM Transactions on Internet Technology (TOIT), 21(2): 49:1–49:26, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3408296.

Wu:2021:NRT

[WSM21] Di Wu, Wei Shi, and Xiangyu Ma. A novel realtime anti-spam framework. ACM Transactions on Internet Technology (TOIT), 21(4):88:1– 88:27, November 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3423153.

Wu:2021:EMS

[WTS^{+}21] Jimmy Ming-Tai Wu, Qian Teng, Gautam Srivastava, Matin Pirouz, and Jerry Chun-Wei Lin. The efficient mining of skyline patterns from a volunteer computing network. ACM Transactions on Internet Technology (TOIT), 21(4): 89:1–89:20, July 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3423557.

Weiss:2021:PQM

[WVHTK21] Iris Weiss, Birgit Vogel-Heuser, Emanuel Trunzer, and Simon Kruppa. Product quality monitoring in hydraulic presses using a minimal sample of sensor and actuator data. ACM Transactions on Internet Technology (TOIT), 21(2): 37:1–37:23, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3436238.

Wang:2022:NBF

[WWJ^{+}22] Derui Wang, Sheng Wen, Alireza Jolfaei, Mohammad Sayad Haghighi, Surya Nepal, and Yang Xiang. On the neural backdoor of federated generative models in edge computing. ACM Transactions on Internet Technology (TOIT), 22(2): 43:1–43:21, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3425662.

Wang:2023:HSF

[WWZ^{+}23] Hucheng Wang, Zhi Wang, Lei Zhang, Xiaonan Luo, and Xinheng Wang. A highly stable fusion positioning system of smartphone under NLoS acoustic indoor environment. ACM Transactions on Internet Technology (TOIT), 23 (2):30:1–30:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3589765.

Wolf:2001:BLC

[WY01] Joel L. Wolf and Philip S.

Yu. On balancing the load in a clustered web farm. ACM Transactions on Internet Technology (TOIT), 1(2):231– 261, November 2001. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Wu:2023:FSM

[WYC^{+}23] Feijie Wu, Ho Yin Yuen, Henry Chan, Victor C. M. Leung, and Wei Cai. Facilitating serverless matchbased online games with novel blockchain technologies. ACM Transactions on Internet Technology (TOIT), 23(1): 10:1–10:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3565884.

Wang:2021:FPP

[WZB^{+}21] Tao Wang, Zhigao Zheng, Ali Kashif Bashir, Alireza Jolfaei, and Yanyan Xu. FinPrivacy: a privacypreserving mechanism for fingerprint identification. ACM Transactions on Internet Technology (TOIT), 21(3):

56:1–56:15, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3387130.

Wang:2019:MTR

[WZKP19] Qingyang Wang, Shungeng Zhang, Yasuhiko Kanemasa, and Calton Pu. Mitigating tail response time of n-tier applications: The impact of asynchronous invocations. ACM Transactions on Internet Technology (TOIT), 19(3): 36:1–36:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3340462.

Xiong:2007:PDP

[XCL07] Li Xiong, Subramanyam Chitti, and Ling Liu. Preserving data privacy in outsourcing data aggregation services. ACM Transactions on Internet Technology (TOIT), 7(3): 17:1–17:??, August 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Xiao:2022:PTP

[XCRY22] Song Xiao, Kai Chen, Xiaoxiang Ren, and Haitao Yuan. Pedestrian trajectory prediction in heterogeneous traffic using facial keypoints-based convolutional encoderdecoder network. ACM Transactions on Internet Technology (TOIT), 22(4):83:1–83:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3410444.

Xiao:2023:UFT

[XFL^{+}23] Wenhua Xiao, Xudong Fang, Bixin Liu, Ji Wang, and Xiaomin Zhu. UNION: Fault-tolerant cooperative computing in opportunistic mobile edge cloud. ACM Transactions on Internet Technology (TOIT), 23(4): 59:1–59:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3617994.

Xu:2022:NQD

[XIS22] Lanyu Xu, Arun Iyengar, and Weisong Shi. NLU-Broker: a QoE-driven broker system for natural language understanding services. ACM Transactions on Internet Technology (TOIT), 22(3): 69:1–69:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3497807.

Xu:2020:TTT

[XJ20] Runhua Xu and James Joshi. Trustworthy and transparent thirdparty authority. ACM Transactions on Internet Technology (TOIT), 20(4):31:1–31:23, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3386262.

Xie:2020:RLA

[XLL20] Hong Xie, Yongkun Li, and John C. S. Lui. A reinforcement learning approach to optimize discount and reputation tradeoffs in ecommerce systems. ACM Transactions on Internet Technology (TOIT), 20(4):37:1–37:26, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3400024.

Xu:2017:CBD

[XM17] Zhen Xu and James Miller. Cross-browser differences detection based on an empirical metric for Web page visual similarity. ACM Transactions on Internet Technology (TOIT), 18(3): 34:1–34:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Xu:2023:MSC

[XSSD23] Yibin Xu, Jianhua Shao, Tijs Slaats, and Boris Düdder. MWPoW+: a strong consensus protocol for intra-shard consensus in blockchain sharding. ACM Transactions on Internet Technology (TOIT), 23(2): 34:1–34:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3584020.

Xu:2022:EDN

[XSW^{+}22] Minxian Xu, Chenghao Song, Huaming Wu, Sukhpal Singh Gill, Kejiang Ye, and Chengzhong Xu. es-DNN: Deep neural network based multivariate workload prediction in cloud computing environments. ACM Transactions on Internet Technology (TOIT), 22(3): 75:1–75:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3524114.

Xie:2018:ISI

[XvHWW18] Tao Xie, Andre van Hoorn, Huaimin Wang, and Ingo Weber. Introduction to the special issue on emerging software technologies for Internetbased systems: Internetware and DevOps. ACM Transactions on Internet Technology (TOIT), 18(2):13:1–13:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Xie:2019:PYS

[XWML19] Hong Xie, Weijie Wu, Richard T. B. Ma, and John C. S. Lui. Pay as your service needs: an application-driven pricing approach for the Internet economics. ACM Transactions on Internet Technology (TOIT), 19(4):52:1–52:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

                                        URL https://dl.acm.
                                        org/ft_gateway.cfm?
                                        id=3361148.

Xia:2022:CIP

[XZG^{+}22] Zhuoqun Xia, Lingxuan Zeng, Ke Gu, Xiong Li, and Weijia Jia. Conditional identity privacypreserving authentication scheme based on cooperation of multiple fog servers under fog computing-based IoVs. ACM Transactions on Internet Technology (TOIT), 22(4):107:1– 107:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3538381.

Xia:2022:ISS

[XZJO22] Kaijian Xia, Wenbing Zhao, Alireza Jolfaei, and Tamer Ozsu. Introduction to the special section on edge/fog computing for infectious disease intelligence. ACM Transactions on Internet Technology (TOIT), 22(3): 63:1–63:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3494119.

Xu:2021:CPB

[XZY^{+}21] Xiaolong Xu, Dawei Zhu, Xiaoxian Yang, Shuo Wang, Lianyong Qi, and Wanchun Dou. Concurrent practical Byzantine fault tolerance for integration of blockchain and supply chain. ACM Transactions on Internet Technology (TOIT), 21(1):7:1–7:17, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3395331.

Xue:2008:IWS

[XZZ08] Xiao-Bing Xue, Zhi-Hua Zhou, and Zhongfei (Mark) Zhang. Improving Web search using image snippets. ACM Transactions on Internet Technology (TOIT), 8(4):21:1–21:??, September 2008. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yin:2002:EWC

[YADI02] Jian Yin, Lorenzo Alvisi, Mike Dahlin, and Arun Iyengar. Engineering web cache consistency. ACM Transactions on Internet Technology (TOIT), 2(3): 224–259, August 2002. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yoshikawa:2001:XPB

[YASU01] Masatoshi Yoshikawa, Toshiyuki Amagasa, Takeyuki Shimura, and Shunsuke Uemura. XRel: a pathbased approach to storage and retrieval of XML documents using relational databases. ACM Transactions on Internet Technology (TOIT), 1(1): 110–141, August 2001. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yus:2022:SES

[YBMV22] Roberto Yus, Georgios Bouloukakis, Sharad Mehrotra, and Nalini Venkatasubramanian. The SemIoTic ecosystem: a semantic bridge between IoT devices and smart spaces. ACM Transactions on Internet Technology (TOIT), 22(3): 76:1–76:??, August 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3527241.

Yousfi:2019:ABP

[YBW19] Alaaeddine Yousfi, Kimon Batoulis, and Mathias Weske. Achieving business process improvement via ubiquitous decision-aware business processes. ACM Transactions on Internet Technology (TOIT), 19(1): 14:1–14:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ye:2014:EMD

[YBZ14] Zhen Ye, Athman Bouguettaya, and Xiaofang Zhou. Economic model-driven cloud service composition. ACM Transactions on Internet Technology (TOIT), 14(2–3):20:1– 20:??, October 2014. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yang:2018:GTM

[YC18] Zhi Yang and Wei Chen. A game theoretic model for the formation of navigable small-world networks-the tradeoff between distance and reciprocity. ACM Transactions on Internet Technology (TOIT), 18(4): 56:1–56:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yuksel:2017:BBH

[YCC17] Beste F. Yuksel, Penny Collisson, and Mary Czerwinski. Brains or beauty: How to engender trust in user-agent interactions. ACM Transactions on Internet Technology (TOIT), 17(1): 2:1–2:??, March 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yi:2022:ISI

[YCH^{+}22] Haibo Yi, Ruinan Chi, Xin Huang, Xuejun Cai, and Zhe Nie. Improving security of Internet of Vehicles based on postquantum signatures with systolic divisions. ACM Transactions on Internet Technology (TOIT), 22(4):82:1–82:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3410445.

Yuan:2013:PVQ

[YCM^{+}13] Lihua Yuan, Chao-Chih Chen, Prasant Mohapatra, Chen-Nee Chuah, and Krishna Kant. A proxy view of quality of Domain Name Service, poisoning attacks and survival strategies. ACM Transactions on Internet Technology (TOIT), 12 (3):9:1–9:??, May 2013. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yue:2021:PPT

[YDZ^{+}21] Zijie Yue, Shuai Ding, Lei Zhao, Youtao Zhang, Zehong Cao, M. Tanveer, Alireza Jolfaei, and Xi Zheng. Privacypreserving time-series medical images analysis using a hybrid deep learning framework. ACM Transactions on Internet Technology (TOIT), 21(3):57:1–57:21, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3383779.

Yan:2022:CFC

[YJL^{+}22] Hongyang Yan, Nan Jiang, Kang Li, Yilei Wang, and Guoyu Yang. Collusion-free for cloud verification toward the view of game theory. ACM Transactions on Internet Technology (TOIT), 22(2):33:1–33:21, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3423558.

Yuan:2022:AFG

[YLC^{+}22] Yali Yuan, Chencheng Liang, Xu Chen, Thar Baker, and Xiaoming Fu. Adaptive fuzzy gamebased energy-efficient localization in 3D underwater sensor networks. ACM Transactions on Internet Technology (TOIT), 22(2): 29:1–29:20, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3406533.

Yang:2017:DSC

[YLL^{+}17] Zhenguo Yang, Qing Li, Zheng Lu, Yun Ma, Zhiguo Gong, and Wenyin Liu. Dual structure constrained multimodal feature coding for social event detection from Flickr data. ACM Transactions on Internet Technology (TOIT), 17 (2):19:1–19:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yang:2023:III

[YLM^{+}23] Li Yang, Xi Li, Zhuoru Ma, Lu Li, Neal Xiong, and Jianfeng Ma. IRGA: an intelligent implicit real-time gait authentication system in heterogeneous complex scenarios. ACM Transactions on Internet Technology (TOIT), 23(2): 35:1–35:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3594538.

Yuan:2021:DMS

[YLZ^{+}21] Bin Yuan, Chen Lin, Deqing Zou, Laurence Tianruo Yang, and Hai Jin. Detecting malicious switches for a secure software-defined tactile Internet. ACM Transactions on Internet Technology (TOIT), 21(4):84:1–84:23, November 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3415146.

Yang:2023:PCS

[YMY^{+}23] Fanyi Yang, Huifang Ma, Cairui Yan, Zhixin Li, and Liang Chang. Polarized communities search via co-guided random walk in attributed signed networks. ACM Transactions on Internet Technology (TOIT), 23(4):58:1–58:??, November 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3613449.

Yan:2021:SCD

[YPFY21] Zheng Yan, Li Peng, Wei Feng, and Laurence T. Yang. Socialchain: Decentralized trust evaluation based on blockchain in pervasive social networking. ACM Transactions on Internet Technology (TOIT), 21(1):17:1–17:28, February 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3419102.

Yao:2016:TIR

[YSNL16] Lina Yao, Quan Z. Sheng, Anne H. H. Ngu, and Xue Li. Things of interest recommendation by leveraging heterogeneous relations in the Internet of Things. ACM Transactions on Internet Technology (TOIT), 16 (2):9:1–9:??, April 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yao:2017:CLR

[YSW^{+}17] Lina Yao, Quan Z.

Sheng, Xianzhi Wang, Wei Emma Zhang, and Yongrui Qin. Collaborative location recommendation by integrating multi-dimensional contextual information. ACM Transactions on Internet Technology (TOIT), 18(3):32:1–32:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yang:2022:FPP

[YSZ^{+}22] Huijie Yang, Jian Shen, Tianqi Zhou, Sai Ji, and Pandi Vijayakumar. A flexible and privacypreserving collaborative filtering scheme in cloud computing for VANETs. ACM Transactions on Internet Technology (TOIT), 22(2):44:1–44:19, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3425708.

Yadav:2022:LIN

[YV22] Ashima Yadav and Dinesh Kumar Vishwakarma. A language-independent network to analyze the impact of COVID-19 on the world via sentiment analysis. ACM Transactions on Internet Technology (TOIT), 22(1):28:1–28:30, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3475867.

Yue:2010:BTP

[YW10] Chuan Yue and Haining Wang. BogusBiter: a transparent protection against phishing attacks. ACM Transactions on Internet Technology (TOIT), 10(2):6:1–6:??, May 2010. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Ye:2019:PBS

[YWML19] Li Ye, Weijie Wu, Richard T. B. Ma, and John C. S. Lui. On the profitability of bundling sale strategy for online service markets with network effects. ACM Transactions on Internet Technology (TOIT), 19(3):31:1–31:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.

org/ft_gateway.cfm?
id=3277667.

Yin:2021:LDA

[YXL^{+}21] Yuyu Yin, Haoran Xu, Tingting Liang, Manman Chen, Honghao Gao, and Antonella Longo. Leveraging data augmentation for service QoS prediction in cyber-physical systems. ACM Transactions on Internet Technology (TOIT), 21(2): 35:1–35:25, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3425795.

Yang:2018:IVA

[YXP^{+}18] Wenhua Yang, Chang Xu, Minxue Pan, Xiaoxing Ma, and Jian Lu. Improving verification accuracy of CPS by modeling and calibrating interaction uncertainty. ACM Transactions on Internet Technology (TOIT), 18(2):20:1–20:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yu:2019:FGE

[YYM^{+}19] Zhiwen Yu, Fei Yi, Chao Ma, Zhu Wang, Bin Guo, and Liming Chen. Finegrained emotion role detection based on retweet information. ACM Transactions on Internet Technology (TOIT), 19 (1):1:1–1:??, March 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Yang:2024:ATA

[YZL^{+}24] Xuezheng Yang, Zhiwen Zeng, Anfeng Liu, Neal N. Xiong, and Shaobo Zhang. ADTO: a trust active detectingbased task offloading scheme in edge computing for Internet of Things. ACM Transactions on Internet Technology (TOIT), 24(1): 5:1–5:??, February 2024. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3640013.

Yang:2014:SMU

[YZY^{+}14] Dingqi Yang, Daqing Zhang, Zhiyong Yu, Zhiwen Yu, and Djamal Zeghlache. SESAME: Mining user digital footprints for fine-grained preference-aware social media search. ACM Transactions on Internet Technology (TOIT), 14 (4):28:1–28:??, December 2014. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhong:2020:CEC

[ZB20] Zhiheng Zhong and Rajkumar Buyya. A costefficient container orchestration strategy in Kubernetes-based cloud computing infrastructures with heterogeneous resources. ACM Transactions on Internet Technology (TOIT), 20(2): 15:1–15:24, May 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/abs/
10.1145/3378447.

Zhang:2019:ISS

[ZBF^{+}19] Jie Zhang, Jamal Bentahar, Rino Falcone, Timothy J. Norman, and Murat Sensoy. Introduction to the special section on trust and AI. ACM Transactions on Internet Technology (TOIT), 19(4):44:1–44:??, November 2019. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

URL https://dl.acm.
org/ft_gateway.cfm?
id=3365675.

Zhou:2018:OAT

[ZDCB18] Bowen Zhou, Amir Vahid Dastjerdi, Rodrigo N. Calheiros, and Rajkumar Buyya. An online algorithm for task offloading in heterogeneous mobile clouds. ACM Transactions on Internet Technology (TOIT), 18(2): 23:1–23:??, March 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zdun:2008:PBD

[Zdu08]  Uwe Zdun.  Pattern-

based design of a serviceoriented middleware for remote object federations. ACM Transactions on Internet Technology (TOIT), 8(3):15:1–15:??, May 2008. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zaeem:2018:PAS

[ZGB18] Razieh Nokhbeh Zaeem, Rachel L. German, and K. Suzanne Barber. PrivacyCheck: Automatic summarization of privacy policies using data mining. ACM Transactions on Internet Technology (TOIT), 18(4): 53:1–53:??, November 2018. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2023:PRS

[ZGD23] Wenzhao Zhang, Yi Gao, and Wei Dong. Providing realtime support for containerized edge services. ACM Transactions on Internet Technology (TOIT), 23(4):56:1– 56:??, November 2023.

CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3617123.

Zhang:2023:FSN

[ZGF^{+}23] Chong Zhang, Qiang Guo, Luoyi Fu, Jiaxin Ding, Xinde Cao, Fei Long, Xinbing Wang, and Chenghu Zhou. Finding the source in networks: an approach based on structural entropy. ACM Transactions on Internet Technology (TOIT), 23(1):17:1–17:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3568309.

Zhou:2009:UFC

[ZH09] Duanning Zhou and Wayne Wei Huang. Using a fuzzy classification approach to assess e-commerce Web sites: an empirical investigation. ACM Transactions on Internet Technology (TOIT), 9(3):12:1–12:??, July 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhou:2007:SAH

[ZHDD07] Jing Zhou, Wendy Hall, David C. De Roure, and Vijay K. Dialani. Supporting ad-hoc resource sharing on the Web: a peer-to-peer approach to hypermedia link services. ACM Transactions on Internet Technology (TOIT), 7(2):11:1–11:??, May 2007. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhu:2004:PMC

[ZHH04] Jianhan Zhu, Jun Hong, and John G. Hughes. PageCluster: Mining conceptual link hierarchies from Web log files for adaptive Web site navigation. ACM Transactions on Internet Technology (TOIT), 4(2):185– 208, May 2004. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2016:TAD

[ZHL^{+}16] Peng Zhang, Jing He, Guodong Long, Guangyan Huang, and Chengqi Zhang. Towards anomalous diffusion sources detection in a large network. ACM Transactions on Internet Technology (TOIT), 16(1):2:1–2:??, February 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhuang:2015:PBM

[ZJL^{+}15] Yi Zhuang, Nan Jiang, Qing Li, Lei Chen, and Chunhua Ju. Progressive batch medical image retrieval processing in mobile wireless networks. ACM Transactions on Internet Technology (TOIT), 15 (3):9:1–9:??, September 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2021:EDU

[ZJQ^{+}21] Yuanpeng Zhang, Yizhang Jiang, Lianyong Qi, Md Zakirul Alam Bhuiyan, and Pengjiang Qian. Epilepsy diagnosis using multi-view & multimedoid entropy-based clustering with privacy protection. ACM Transactions on Internet Technology (TOIT), 21(2): 48:1–48:21, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3404893.

Zendehdel:2022:ASA

[ZKC^{+}22] Ghazale Amel Zendehdel, Ratinder Kaur, Inderpreet Chopra, Natalia Stakhanova, and Erik Scheme. Automated security assessment framework for wearable BLE-enabled health monitoring devices. ACM Transactions on Internet Technology (TOIT), 22(1):14:1–14:31, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3448649.

Zabolotnyi:2015:JCG

[ZLHD15] Rostyslav Zabolotnyi, Philipp Leitner, Waldemar Hummer, and Schahram Dustdar. JCloudScale: Closing the gap between IaaS and PaaS. ACM Transactions on Internet Technology (TOIT), 15 (3):10:1–10:??, September 2015. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhou:2020:ICF

[ZLL^{+}20] Yang Zhou, Ling Liu, Kisung Lee, Balaji Palanisamy, and Qi Zhang. Improving collaborative filtering with social influence over heterogeneous information networks. ACM Transactions on Internet Technology (TOIT), 20(4):36:1–36:29, November 2020. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3397505.

Zheng:2022:ESP

[ZLS^{+}22] Xiao Zheng, Mingchu Li, Syed Bilal Hussain Shah, Dinh-Thuan Do, Yuanfang Chen, Constandinos X. Mavromoustakis, George Mastorakis, and Evangelos Pallis. Enhancing security-problem-based deep learning in mobile edge computing. ACM Transactions on Internet Technology (TOIT), 22(2):49:1–49:15, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3458931.

Zeng:2023:FRL

[ZLZ^{+}23] Man Zeng, Dandan Li, Pei Zhang, Kun Xie, and Xiaohong Huang. Federated route leak detection in inter-domain routing with privacy guarantee. ACM Transactions on Internet Technology (TOIT), 23(1):12:1– 12:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3561051.

Zhou:2022:PRS

[ZMGW22] Ao Zhou, Xiao Ma, Siyi Gao, and Shangguang Wang. Providing reliable service for parkedvehicle-assisted mobile edge computing. ACM Transactions on Internet Technology (TOIT), 22(4):91:1–91:??, November 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3514242.

Zhang:2023:SLS

[ZMT^{+}23] Yazhou Zhang, Dan Ma, Prayag Tiwari, Chen Zhang, Mehedi Masud, Mohammad Shorfuzzaman, and Dawei Song. Stance-level sarcasm detection with BERT and stance-centered graph attention networks. ACM Transactions on Internet Technology (TOIT), 23 (2):27:1–27:??, May 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3533430.

Zhan:2011:ADD

[ZOC11] Justin Zhan, B. John Oommen, and Johanna Crisostomo. Anomaly detection in dynamic systems using weak estimators. ACM Transactions on Internet Technology (TOIT), 11(1):3:1–3:??, July 2011. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2017:DDP

[ZSL^{+}17] Wei Emma Zhang, Quan Z. Sheng, Jey Han Lau, Ermyas Abebe, and Wenjie Ruan. Duplicate detection in programming question answering communities. ACM Transactions on Internet Technology (TOIT), 18(3): 37:1–37:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2017:LBF

[ZSY^{+}17] Wei Emma Zhang, Quan Z.

Sheng, Lina Yao, Kerry Taylor, Ali Shemshadi, and Yongrui Qin. A learning-based framework for improving querying on Web interfaces of curated knowledge bases. ACM Transactions on Internet Technology (TOIT), 18(3): 35:1–35:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2023:CEC

[ZTH^{+}23] Bolin Zhang, Zhiying Tu, Shaoshi Hang, Dianhui Chu, and Xiaofei Xu. Conco-ERNIE: Complex user intent detect model for smart healthcare cognitive bot. ACM Transactions on Internet Technology (TOIT), 23(1):21:1–21:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3574135.

Zhang:2021:PRF

[ZTL^{+}21] Chen Zhang, Zhuo Tang, Kenli Li, Jianzhong Yang, and Li Yang. A polishing robot force control system based on time series data in industrial Internet of Things. ACM Transactions on Internet Technology (TOIT), 21(2):34:1–34:22, June 2021. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3419469.

Zander:2017:WTY

[ZW17] Sebastian Zander and Xuequn Wang. Are we there yet? IPv6 in Australia and China. ACM Transactions on Internet Technology (TOIT), 18 (3):36:1–36:??, May 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2017:AMR

[ZWC^{+}17] Haibo Zhang, Luming Wan, Yawen Chen, Laurence T. Yang, and Lizhi Peng. Adaptive message routing and replication in mobile opportunistic networks for connected communities. ACM Transactions on Internet Technology (TOIT), 18(1):

2:1–2:??, December 2017. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2022:TOT

[ZWC^{+}22] Rui Zhang, Libing Wu, Shuqin Cao, Xinrong Hu, Shan Xue, Dan Wu, and Qingan Li. Task offloading with task classification and offloading nodes selection for MEC-enabled IoV. ACM Transactions on Internet Technology (TOIT), 22(2):51:1–51:24, May 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL

https://dl.acm.org/
doi/10.1145/3475871.

Zhang:2023:SMT

[ZWW^{+}23] Rongjunchen Zhang, Tingmin Wu, Sheng Wen, Surya Nepal, Cecile Paris, and Yang Xiang. SAM: Multi-turn response selection based on semantic awareness matching. ACM Transactions on Internet Technology (TOIT), 23(1):3:1–3:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3545570.

Zhang:2016:PAG

[ZXH16] Yuexin Zhang, Yang Xiang, and Xinyi Huang.

Password-authenticated group key exchange: a cross-layer design. ACM Transactions on Internet Technology (TOIT), 16 (4):24:1–24:??, December 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2022:VRA

[ZXP^{+}22] Di Zhang, Feng Xu, Chi-Man Pun, Yang Yang, Rushi Lan, Liejun Wang, Yujie Li, and Hao Gao. Virtual reality aided high-quality 3D reconstruction by remote drones. ACM Transactions on Internet Technology (TOIT), 22(1):18:1–18:20, February 2022. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3458930.

Zhuge:2008:RSM

[ZXS08] Hai Zhuge, Yunpeng Xing, and Peng Shi. Resource space model, OWL and database: Mapping and integration. ACM Transactions on Internet Technology (TOIT), 8(4):20:1–20:??, September 2008. CO-DEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2016:DEP

[ZXYL16] Rui Zhang, Rui Xue, Ting Yu, and Ling Liu. Dynamic and efficient private keyword search over inverted index-based encrypted data. ACM Transactions on Internet Technology (TOIT), 16(3):21:1–21:??, August 2016. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Zhang:2023:LCD

[ZZF^{+}23] Wenzhao Zhang, Yuxuan Zhang, Hongchang Fan, Yi Gao, and Wei Dong. A low-code development framework for cloud-native edge systems. ACM Transactions on Internet Technology (TOIT), 23(1):15:1– 15:??, February 2023. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3563215.

Zhang:2022:DDG

[ZZW^{+}22] Xiongtao Zhang, Xiaomin Zhu, Ji Wang, Weidong Bao, and Laurence T. Yang. DANCE: Distributed generative adversarial networks with communication compression. ACM Transactions on Internet Technology (TOIT), 22(2): 50:1–50:32, May 2022.

CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3458929.